### v1

# Asymmetric Perturbation in Solving Bilinear Saddle-Point Optimization

Abstract
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in min-max problems, game theory, and constrained optimization. Perturbing payoffs or values are known to be effective in stabilizing learning dynamics and finding an exact solution or equilibrium. However, it requires careful adjustment of the perturbation magnitude; otherwise, learning dynamics converge to only an equilibrium. We establish an impossibility result that it almost never reaches an exact equilibrium as long as both players’ payoff functions are perturbed. To overcome this, we introduce an asymmetric perturbation approach, where only one player’s payoff function is perturbed. This ensures convergence to an equilibrium without requiring parameter adjustments, provided the perturbation strength parameter is sufficiently low. Furthermore, we empirically demonstrate fast convergence toward equilibria in both normal-form and extensive-form games.
1 Introduction
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in min-max problems, game theory, and constrained optimization. Over the past decade, no-regret learning algorithms have been extensively studied for computing (approximate) solutions or equilibria. When each player minimizes regret, the time-averaged strategies approximate Nash equilibria in two-player zero-sum games, that is, the algorithms guarantee average-iterate convergence. However, the actual sequence of strategies does not necessarily converge: it can cycle or diverge even in simple bilinear cases (Mertikopoulos et al., 2018; Bailey and Piliouras, 2018; Cheung and Piliouras, 2019). This is problematic, especially in large-scale games where neural networks are used for strategy representation, as averaging requires storing a separate model copy at each iteration.
This is one of the main motivations for the study of last-iterate convergence, a stronger notion than average-iterate convergence, in which the strategies themselves converge to an equilibrium. One successful approach is to use optimistic learning algorithms, which essentially incorporate a one-step optimistic prediction that the environment will behave similarly in the next step. This idea has led to a number of effective algorithms, including Extra-Gradient methods (EG) (Liang and Stokes, 2019; Mokhtari et al., 2020), Optimistic Gradient Descent Ascent (OGDA) (Daskalakis and Panageas, 2019; Gidel et al., 2019; Mertikopoulos and Zhou, 2019), and Optimistic Multiplicative Weights Update (OMWU) (Daskalakis and Panageas, 2019; Lei et al., 2021a).
Alternatively, perturbing the payoffs by introducing strongly convex penalties (Facchinei and Pang, 2003) has long been recognized as a technique for achieving last-iterate convergence (Koshal et al., 2010; Tatarenko and Kamgarpour, 2019). This was typically done by decaying the perturbation magnitude over time, leading to asymptotic convergence. This idea has re-emerged through a different way of periodically updating the magnitude of perturbation, thereby stabilizing the learning dynamics (Perolat et al., 2021; Abe et al., 2022, 2023, 2024). Notably, such approaches have been successfully applied in practical settings, including learning in large-scale games (Perolat et al., 2022) and fine-tuning large language models via preference optimization (Munos et al., 2024), often in place of optimistic algorithms. However, these methods still require careful tuning of the perturbation strength; otherwise, the dynamics may converge only to an approximate equilibrium.
To clarify this limitation, we first prove an impossibility result showing that learning dynamics almost never reach an exact equilibrium when both players’ payoff functions are perturbed—a setting considered in most existing perturbation-based methods. In other words, symmetric regularization typically results in solutions that do not exactly match the equilibrium of the original game, regardless of how small the perturbation is.
Motivated by this observation, we develop an asymmetric perturbation approach, where only one player’s payoff function is perturbed, while the other remains unperturbed. This simple modification yields a qualitatively different outcome: the learning dynamics converge to an exact saddle point, as long as the perturbation strength is sufficiently small—within a broad and practical range—without requiring fine-tuned parameter selection. Intuitively, leaving the opponent player $y$ unperturbed preserves the linearity of player $x$’s original objective, so adding a strongly convex perturbation does not significantly alter the solution. Consequently, even if strongly convex penalties are applied to the payoff function, the solution remains consistent between the original game and the perturbed game.
Furthermore, to demonstrate the effectiveness of our proposed approach, we provide two applications in normal-form and extensive-form games. First, we incorporate the technique into a gradient-based learning algorithm, which provably converges to a saddle point with a guaranteed rate. Empirical results on benchmark normal-form games show accelerated convergence. Second, we apply the technique to counterfactual regret (CFR) minimization (Zinkevich et al., 2007), a widely used method in extensive-form games, and demonstrate significant improvements in convergence speed on standard benchmarks.
2 Preliminaries
Bilinear saddle-point optimization problems.
In this study, we focus on the following bilinear saddle-point problem:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay,$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ (resp. $\mathcal{Y}\subseteq\mathbb{R}^{n}$) represents the $m$-dimensional (resp. $n$-dimensional) convex strategy space for player $x$ (resp. player $y$), and $A\in\mathbb{R}^{m\times n}$ is a game matrix. We assume that $\mathcal{X}$ and $\mathcal{Y}$ are polytopes. We refer to the function $x^{\top}Ay$ as the payoff function, and write $z=(x,y)$ as the strategy profile. This formulation includes many well-studied classes of games, such as two-player normal-form games and extensive-form games with perfect recall [^1] Two-player extensive-form games with perfect recall can be expressed as bilinear problems using sequence-form strategies (von Stengel, 1996).
Nash equilibrium.
This study aims to compute a minimax or maximin strategy in the optimization problem (1). Let $\mathcal{X}^{\ast}:=\mathop{\rm arg~{}min}\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}x^{\top}Ay$ denote the set of minimax strategies, and let $\mathcal{Y}^{\ast}:=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}x^{\top}Ay$ denote the set of maximin strategies. It is well-known that any strategy profile $(x^{\ast},y^{\ast})\in\mathcal{X}^{\ast}\times\mathcal{Y}^{\ast}$ is a Nash equilibrium, which satisfies the following condition:
$\forall(x,y)\in\mathcal{X}\times\mathcal{Y},~{}(x^{\ast})^{\top}Ay\leq(x^{\ast})^{\top}Ay^{\ast}\leq x^{\top}Ay^{\ast}.$
Based on the minimax theorem (v. Neumann, 1928), every equilibrium $(x^{\ast},y^{\ast})\in\mathcal{X}^{\ast}\times\mathcal{Y}^{\ast}$ attains the identical value, denoted as $v^{\ast}$, which can be expressed as:
$v^{\ast}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay=\max_{y\in\mathcal{Y}}\min_{x\in\mathcal{X}}x^{\top}Ay.$
We refer to $v^{\ast}$ as the game value. To quantify the proximity to equilibrium for a given strategy profile $(x,y)$, we use NashConv, which is defined as follows:
$\mathrm{NashConv}(x,y)$ $=\max_{\tilde{y}\in\mathcal{Y}}x^{\top}A\tilde{y}-\min_{\tilde{x}\in\mathcal{X}}(\tilde{x})^{\top}Ay.$
3 Asymmetric payoff perturbation
In this section, we explain our novel technique of asymmetric payoff perturbation. We first revisit the commonly used symmetric perturbation, where we show that the resulting solution of the perturbed game generally deviates from the equilibrium in the original game. Then, we demonstrate that a seemingly minor structural change—perturbing only one player’s payoff—can yield a dramatically different outcome: the solution of the perturbed game exactly matches the equilibrium in many cases.
3.1 Equilibrium shift under symmetric perturbation
Payoff perturbation is an extensively studied technique for solving games (Facchinei and Pang, 2003; Liu et al., 2023). In this approach, the payoff functions of all players are perturbed by a strongly convex function $\psi$. For example, in bilinear games, instead of solving the original game in (1), we solve the following perturbed game:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\mu\psi(x)-\mu\psi(y)\right\},$
where $\mu\in(0,\infty)$ is the perturbation strength. Since the perturbation is applied to both players’ payoff functions, we refer to this perturbed game as a symmetrically perturbed game. In this study, we specifically focus on the standard case, where the perturbation payoff function $\psi$ is given by the squared $\ell^{2}$-norm, i.e., $\psi(x)=\frac{1}{2}\|x\|^{2}$:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$ (2)
Let $x^{\mu}$ (resp. $y^{\mu}$) denote the minimax (resp. maximin) strategy in the symmetrically perturbed game (2) 222 The minimax strategy is uniquely determined because symmetrically perturbed games satisfy the strongly convex–strongly concave property., which can be solved at an exponentially fast rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). It has also been shown that the distance between the solution of the symmetrically perturbed game $(x^{\mu},y^{\mu})$ and the solution in the original game $(x^{\ast},y^{\ast})$ is upper bounded by $\mathcal{O}(\mu)$ (Liu et al., 2023; Abe et al., 2024). However, they do not guarantee that the two solutions coincide, even for a small $\mu>0$. In contrast, the following theorem provides the first formal impossibility result, proving that $(x^{\mu},y^{\mu})$ almost never coincides with $(x^{\ast},y^{\ast})$.
Theorem 3.1.
Consider a normal-form game with a unique interior equilibrium. Assume that at this equilibrium, neither player chooses their actions uniformly at random, i.e., $(x^{\ast},y^{\ast})\neq\left(\frac{1}{m}\mathbf{1}_{m},\frac{1}{n}\mathbf{1}_{n}\right)$. Then, for any $\mu>0$ such that $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$, the minimax strategy $x^{\mu}$ in (2) satisfies $x^{\mu}\neq x^{\ast}$. Furthermore, for any $\mu>0$ such that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the maximin strategy $y^{\mu}$ in (2) satisfies $y^{\mu}\neq y^{\ast}$.
The proof is provided in Appendix D. Additionally, we extend our analysis to the case where both players have different perturbation strengths, i.e., $\mu_{x}>0$ and $\mu_{y}>0$, as shown in Appendix B.
Discussion on Theorem 3.1.
The term $\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$ can be interpreted as a measure of the difference between the equilibrium $(x^{\ast},y^{\ast})$ and the uniform random strategy profile $(\frac{1}{m}\mathbf{1}_{m},\frac{1}{n}\mathbf{1}_{n})$. Specifically, the numerator $\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}$ represents the difference in the payoffs, while the denominator $\left\|x^{\ast}\right\|^{2}-\frac{1}{m}$ represents the difference in the squared $\ell^{2}$-norms, respectively. A promising direction for future research is to theoretically demonstrate that, when $\mu=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the corresponding equilibrium coincides exactly with the equilibrium in the original game, i.e., $(x^{\mu},y^{\mu})=(x^{*},y^{*})$. We have experimentally confirmed this, and the results are presented in the Appendix A.2.
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 1: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix $A$ is given by $[[0,1,-3],[-1,0,1],[3,-1,0]]$. The proximity to the minimax strategy is measured by the value of $\max_{y\in\mathcal{Y}}(x^{\mu})^{\top}Ay-v^{\ast}$.
When the game is symmetric.
Next, let us consider the case when $A^{\top}=-A$, as in Rock-Paper-Scissors. In this scenario, the equilibrium strategies $x^{\mu}$ and $y^{\mu}$ are not identical to the minimax or maximin strategies of the original game, regardless of the choice of $\mu>0$.
Corollary 3.2.
Assume that $A^{\top}=-A$. Under the same setup as Theorem 3.1, the equilibrium $(x^{\mu},y^{\mu})$ in (2) always satisfies $x^{\mu}\neq x^{\ast}$ and $y^{\mu}\neq y^{\ast}$ for any $\mu>0$.
This is because it always holds that $v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)=0$ when $A^{\top}=-A$. Figure 1 shows the proximity of $x^{\mu}$ to $x^{\ast}$ with varying perturbation strength $\mu$ in a simple biased Rock-Paper-Scissors game. We observe that as long as $\mu>0$, $x^{\mu}$ remains distant from $x^{\ast}$. This observation supports the theoretical results in Theorem 3.1 and Corollary 3.2.
3.2 Asymmetric payoff perturbation
Instead of incorporating the perturbation into both players’ payoff functions, we consider the case where only player $x$’s payoff function is perturbed:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}\right\}.$ (3)
The procedure we are going to describe in Theorem 3.3 focuses on computing the minimax strategy $x^{\ast}$, rather than the maximin strategy $y^{\ast}$. To compute $y^{\ast}$, we simply solve the corresponding maximin problem for player $y$:
$\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$
The same reasoning applies to this perturbed maximin optimization problem. Thus, hereafter, we primarily focus on the perturbed game (3) from the perspective of player $x$.
Figure 2: The landscape of the objective function for player $x$ in asymmetrically perturbed games. The functions $g(x)$ and $g^{\mu}_{\mathrm{asym}}(x)$ are defined as $g(x):=\max\limits_{y\in\mathcal{Y}}x^{\top}Ay$ and $g^{\mu}_{\mathrm{asym}}(x):=g(x)+\frac{\mu}{2}\|x\|^{2}$, respectively.
Since the function $\max_{y\in\mathcal{Y}}x^{\top}Ay$ is convex with respect to $x$ (Boyd and Vandenberghe, 2004), the perturbed objective $\max_{y\in\mathcal{Y}}x^{\top}Ay+\frac{\mu}{2}\|x\|^{2}$ is $\mu$-strongly convex. Therefore, the minimax strategy for the perturbed game (3) is unique. We denote it by $x^{\mu}$ and denote the set of maximin strategies in (3) by $\mathcal{Y}^{\mu}$. Since both the minimax and maximin strategies constitute a Nash equilibrium of the perturbed game, the pair $(x^{\mu},y^{\mu})$ with $y^{\mu}\in\mathcal{Y}^{\mu}$ satisfies the following conditions: for all $\tilde{y}^{\mu}\in\mathcal{Y}^{\mu}$ and $x\in\mathcal{X}$,
$(x^{\mu})^{\top}A\tilde{y}^{\mu}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\leq x^{\top}A\tilde{y}^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2},$ (4)
and for all $y\in\mathcal{Y}$,
$(x^{\mu})^{\top}Ay^{\mu}\geq(x^{\mu})^{\top}Ay.$ (5)
3.3 Equilibrium invariance under the asymmetric perturbation
In this section, we discuss the properties of the minimax strategies for asymmetrically perturbed games. Surprisingly, we can show that $x^{\mu}$ in (3) does correspond to a minimax strategy in the original game (1) for a sufficiently small $\mu$:
Theorem 3.3.
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x^{\ast}\in\mathcal{X}^{\ast}}\left\|x^{\ast}\right\|})$, where $\alpha>0$ is a constant depending only on $\mathcal{X}^{\ast}$. Then, the minimax strategy $x^{\mu}$ in the corresponding asymmetrically perturbed game (3) satisfies $x^{\mu}\in\mathcal{X}^{\ast}$.
Therefore, for all perturbation strengths $\mu$ less than a certain positive constant, $x^{\mu}$ coincides with the minimax strategy of the original game. Figure 1 visualizes this feature in a simple example. As long as the value of $\mu$ is less than $2.5$, the equilibrium strategy of the asymmetric perturbed game is invariant with respect to the original equilibrium strategy.
The key element in proving Theorem 3.3 is the near-linear behavior of the objective function $g(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay$ for player $x$ in the original game. Specifically, according to Wei et al. (2021), there exists a constant $\alpha>0$ such that:
$\forall x^{\ast}\in\mathcal{X}^{\ast},~{}g(x)-g\left(x^{\ast}\right)\geq\alpha\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|,$
where $\prod_{\mathcal{A}}(a)=\mathop{\rm arg~{}min}\limits_{a^{\prime}\in\mathcal{A}}\left\|a-a^{\prime}\right\|$ represents the projection operator onto a given closed convex set $\mathcal{A}$. This inequality implies that deviating from the minimax strategy set $\mathcal{X}^{\ast}$ results in an increase in the objective function proportionally to the distance $\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|$. In contrast, the variation (i.e., the gradient) of the perturbation payoff function $\frac{\mu}{2}\left\|x\right\|^{2}$ can always be bounded by $\mathcal{O}(\mu)$ over $\mathcal{X}$. Hence, by choosing $\mu$ sufficiently small, we can ensure that the perturbation payoff function does not significantly incentivize player $x$ to deviate from $x^{\ast}$. In Figure 2, we illustrate this fact intuitively. The addition of the strongly convex function $\frac{\mu}{2}\|x\|^{2}$ does not shift the optimum $x^{*}$ of the original $g(x)$ if $\mu$ is sufficiently small, as the kink of the lines through $g(x)$ dominates. The detailed proof of Theorem 3.3 is provided in Appendix E.
In summary, to compute a minimax strategy $x^{\ast}\in\mathcal{X}^{\ast}$ in the original game, it is sufficient to solve the asymmetrically perturbed game (3) with a small perturbation strength $\mu>0$. Note that, as mentioned above, we can also compute a maximin strategy $y^{\ast}\in\mathcal{Y}^{\ast}$ by solving the game $\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}$, where the payoff perturbation is applied only to player $y$.
4 Asymmetrically perturbed gradient descent ascent
This section proposes a gradient-based online learning algorithm, Asymmetrically Perturbed Gradient Descent Ascent (AsymP-GDA), for solving asymmetrically perturbed games (3). At each iteration $t\in[T]$, AsymP-GDA updates each player’s strategy according to the following update rule [^3] AsymP-GDA employs alternating updates rather than simultaneous updates, as recent work has demonstrated the advantages of the former over the latter (Lee et al., 2024).:
[…]
where $\eta>0$ is the learning rate. In AsymP-GDA, player $x$’s strategy $x^{t}$ is updated based on the gradient of the perturbed payoff function $x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}$, while player $y$’s strategy $y^{t}$ is updated based on the gradient of the original payoff function $x^{\top}Ay$.
Since the perturbed payoff function of player $x$ is strongly convex, it is anticipated that AsymP-GDA enjoys a last-iterate convergence guarantee. By combining this observation with Theorem 3.3, when $\mu$ is sufficiently small, the updated strategy $x^{t}$ should converge to a minimax strategy $x^{\ast}$ in the original game. We confirm this empirically by plotting the trajectory of $x^{t}$ updated by AsymP-GDA in a sample normal-form game, as shown in Figure 3. We also provide the trajectories of GDA and SymP-GDA; in the latter, the squared $\ell^{2}$-norm perturbs both players’ gradients. For both AsymP-GDA and SymP-GDA, the perturbation strength is set to $\mu=1$. As expected, AsymP-GDA successfully converges to the minimax strategy (red point) in the original game, whereas SymP-GDA converges to a point far from the minimax strategy, and GDA cycles around the minimax strategy. Further details and additional experiments in normal-form games can be found in Appendix A.3.
Figure 3: Trajectories of strategies for player $x$ using AsymP-GDA, SymP-GDA, and GDA. The game matrix $A$ is set to $A=[[0,1,-3],[-1,0,1],[3,-1,0]]$, and the strategy spaces are set to $\mathcal{X}=\mathcal{Y}=\Delta^{3}$. The red point represents the minimax strategy in the original game. The trajectories originate from different initial strategies, demonstrating the learning dynamics under each method.
4.1 Last-iterate convergence rate
In this section, we provide the convergence result of AsymP-GDA. Let $\|A\|$ denote the largest singular value of a given matrix $A$, and $D:=\max_{z,z^{\prime}\in\mathcal{X}\times\mathcal{Y}}\left\|z-z^{\prime}\right\|$ denote the diameter of $\mathcal{X}\times\mathcal{Y}$. Theorem 4.1 demonstrates that AsymP-GDA converges to the minimax/maximin strategies in the asymmetrically perturbed game at a rate of $\mathcal{O}(1/t)$:
Theorem 4.1.
For an arbitrary perturbation strength $\mu>0$, if the learning rate satisfies $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta}{\mu}\right)}\right)$, then $x^{t}$ and $y^{t}$ satisfy for any $t\geq 1$:
$\left\|x^{\mu}-x^{t}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)t},$
where $\beta>0$ is a positive constant depending only on $\mathcal{Y}^{\mu}$.
Note that the statement of Theorem 4.1 holds for any fixed $\mu>0$. Thus, by combining Theorems 3.3 and 4.1, we can conclude that if $\mu$ is sufficiently small (which does not need to depend on the number of iterations $t$ or $T$), player $x$’s strategy $x^{t}$ updated by AsymP-GDA converges to a minimax strategy in the original game (1):
Corollary 4.2.
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x^{\ast}\in\mathcal{X}^{\ast}}\left\|x^{\ast}\right\|})$, and the learning rate is set such that $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta}{\mu}\right)}\right)$. Then, AsymP-GDA ensures the convergence of $x^{t}$ to an equilibrium $x^{\ast}$ in the original game at a rate of $\mathcal{O}(1/t)$:
$\left\|x^{\ast}-x^{t}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)t}.$
Remark 4.3 (Overcoming the limitation of small fixed $\mu$).
One might still be concerned that the permissible perturbation strength $\mu$ in Corollary 4.2 must remain very small when $\alpha$ is close to 00. While it is true that $\alpha$ can, in theory, become arbitrarily small depending on the game’s payoff structure, such pathological cases are uncommon in practice, as demonstrated by our experiments. To mitigate this potential issue, a simple anchoring mechanism can be employed. This approach replaces the perturbation term $\frac{\mu}{2}\|x\|^{2}$ with $\frac{\mu}{2}\|x-\sigma\|^{2}$, where the anchor point $\sigma$ is periodically updated to the current strategy $x^{t}$ (Perolat et al., 2021; Abe et al., 2024). As discussed in AppendixC, this modification keeps the convergence guarantee and remains effective even when $\mu$ is moderately large.
4.2 Proof sketch of Theorem 4.1
This section outlines the proof sketch for Theorem 4.1. The complete proofs are provided in Appendix F.
(1) Monotonic decrease of the distance function.
Firstly, leveraging the strong convexity of the perturbation payoff function, $\frac{\mu}{2}\left\|x\right\|^{2}$, we can show that the distance between the current strategy profile $z^{t}=(x^{t},y^{t})$ and any equilibrium $z^{\mu}=(x^{\mu},y^{\mu})$ monotonically decreases. Specifically, we have for any $t\geq 1$:
$\left\|z^{\mu}-z^{t+1}\right\|^{2}-\left\|z^{\mu}-z^{t}\right\|^{2}\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{t+1}-z^{t}\right\|^{2}.$ (7)
(2) Lower bound on the path length.
The primary technical challenge is deriving the term related to the distance between $y^{t+1}$ and the maximin strategies set $\mathcal{Y}^{\mu}$, i.e., $\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}$, which leads to the last-iterate convergence rate. To this end, we first derive the following lower bound on the path length $\left\|z^{t+1}-z^{t}\right\|$ by the distance $\|x^{\mu}-x^{t+1}\|^{2}$ and $\|A(y^{\mu}-y^{t+1})\|^{2}$ (as shown in Lemmas F.3 and F.4):
$\left\|z^{t+1}-z^{t}\right\|$ $\geq\Omega\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\right).$ (8)
One might think that the term of $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}$ can be 00 even if $y^{t+1}\notin\mathcal{Y}^{\mu}$. However, by proving that $Ay^{\ast}$ attains a unique vector $b^{\ast}\in\mathbb{R}^{m}$ for all $y^{\ast}\in\mathcal{Y}^{\mu}$, and this vector cannot be achieved by any strategy $y\notin\mathcal{Y}^{\mu}$, we demonstrate that $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\neq 0$ as long as $y\notin\mathcal{Y}^{\mu}$. Using this fact, we demonstrate that the term of $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}$ can be lower bounded by the distance between $y^{t+1}$ and the maximin strategy set:
$\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\geq\Omega\left(\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right).$ (9)
By combining (8) and (9), we obtain the following lower bound on $\left\|z^{t+1}-z^{t}\right\|$ by the distance between the current strategy profile and the equilibrium set:
$\!\left\|z^{t+1}\!-\!z^{t}\right\|\!\geq\!\Omega\left(\!\left\|x^{\mu}\!-\!x^{t+1}\right\|^{2}\!+\!\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})\!-\!y^{t+1}\right\|^{2}\right)\!.$ (10)
(3) Last-iterate convergence rate.
Putting (10) into (7), we have for any $t\geq 1$:
$\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{2}\leq\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}-\Omega\left(\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{4}\right),$
where $\mathcal{Z}^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$. Finally, utilizing an auxiliary lemma on recursive formulas by (Wei et al., 2021), we obtain the following upper bound on the distance $\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}$:
$\left\|\prod_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\leq\mathcal{O}(1/t).$
Remark 4.4 (Technical challenge in proving Theorem 4.1).
The main technical challenge in proving Theorem 4.1 arises from the asymmetric nature of the perturbation, which is applied only to player $x$. Unlike the symmetric case, where strong convexity in both players’ payoff functions directly yields contraction, the asymmetric setting requires a more subtle analysis since player $y$’s payoff function remains linear. Our proof establishes a last-iterate convergence rate by lower bounding the projection error of $y^{t}$ through the gap in payoff vectors $Ay^{t}-Ay^{\mu}$, as shown in (9), rather than relying on strong convexity. This contrasts with techniques for unconstrained linear-quadratic games (Zhang et al., 2022), where the dynamics reduce to a linear system; in our constrained setting, projections alter the dynamics fundamentally. See Appendix F for details.
5 Asymmetrically perturbed CFR+
We introduce the concept of asymmetric perturbation into learning algorithms for extensive-form games, which involve sequential decision-making and imperfect information, making them a more complex and realistic setting for strategic interactions. Here, we provide only empirical simulations, unlike in Section 4, as analyzing extensive-form games is significantly more difficult than analyzing normal-form games.
Extensive-form games are commonly solved using Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) and its variants, such as CFR+ (Tammelin, 2014). In CFR+, the cumulative counterfactual regret $\boldsymbol{R}^{x}$ of player $x$ is updated to ensure it does not fall below zero:
$\boldsymbol{R}^{x}_{t+1}$ $=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})\right]^{+},$
where $\boldsymbol{r}({x}_{t})$ represents the immediate regret incurred at iteration $t$. To investigate the impact of perturbations in extensive-form games, we propose Asymmetric Perturbed CFR+ (AsymP-CFR+), which incorporates an asymmetric perturbation, alongside Symmetric Perturbed CFR+ (SymP-CFR+) as a baseline, which applies a symmetric perturbation. In AsymP-CFR+, the cumulative counterfactual regret $\boldsymbol{R}^{x}$ for player $x$ is updated using the following rule:
$\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu\boldsymbol{x}_{t}\right]^{+}.$
In contrast, SymP-CFR+ applies perturbation symmetrically to both players. The pseudocode for AsymP-CFR+ is provided in Algorithm 1 of Appendix A.4.
We compare the NashConv of the last-iterate strategy for AsymP-CFR+ against SymP-CFR+ and baseline algorithms, including CFR, CFR+, LCFR, and DCFR (both from Brown and Sandholm (2019)). Our experiments focus on six different extensive-form games: Kuhn Poker, Leduc Poker, Goofspiel (with four-card and five-card variants), and Liar’s Dice (with four-sided and six-sided dice), all of which are implemented using OpenSpiel (Lanctot et al., 2019). For both AsymP-CFR+ and SymP-CFR+, we set $\mu=0.01$. For DCFR, we use OpenSpiel’s default parameters.
Figure 4 shows the NashConv values for each game. As indicated by these results, AsymP-CFR+ not only converges faster than any other method in all games but also directly reaches an equilibrium strategy, whereas SymP-CFR+ converges near the equilibrium. These results confirm that the asymmetric perturbation leads to convergence in extensive-form games.
Figure 4: Performance of AsymP-CFR+ in extensive-form games.
6 Related literature
Saddle-point optimization problems have attracted significant attention due to their applications in machine learning, such as training generative adversarial networks (Daskalakis et al., 2018). No-regret learning algorithms have been extensively studied with the aim of achieving either average-iterate or last-iterate convergence. To attain last-iterate convergence, many recent algorithms incorporate optimism (Rakhlin and Sridharan, 2013a, b), including optimistic multiplicative weights update (Daskalakis and Panageas, 2019; Lei et al., 2021b; Wei et al., 2021), optimistic gradient descent ascent (Daskalakis et al., 2018; Mertikopoulos et al., 2019; de Montbrun and Renault, 2022), and extra-gradient methods (Golowich et al., 2020; Mokhtari et al., 2020).
As an alternative approach, payoff perturbation has gained renewed attention. In this approach, players’ payoff functions are regularized with strongly convex terms (Cen et al., 2021, 2023; Pattathil et al., 2023), which stabilizes the dynamics and leads to convergence. Some existing works have shown convergence to an approximate equilibrium under fixed perturbation (Sokota et al., 2023; Tuyls et al., 2006; Coucheney et al., 2015; Leslie and Collins, 2005; Abe et al., 2022; Hussain et al., 2023). To recover equilibria of the original game, later studies have employed a decreasing schedule or iterative regularization (Facchinei and Pang, 2003; Koshal et al., 2013; Yousefian et al., 2017; Bernasconi et al., 2022; Liu et al., 2023; Cai et al., 2023), or have updated the regularization center periodically (Perolat et al., 2021; Abe et al., 2023, 2024). In contrast to these approaches, our algorithms converge without decaying or modifying the perturbation.
Extensive-form games, which model sequential and imperfect-information interactions, have also been studied from both theoretical and empirical perspectives. Strategy representations can be broadly classified into sequence-form (von Stengel, 1996) and behavioral-form. Under the sequence-form, optimistic algorithms have been shown to enjoy last-iterate convergence guarantees (Lee et al., 2021). Under the behavioral-form, perturbation-based approaches have also been applied successfully (Perolat et al., 2021; Sokota et al., 2023; Liu et al., 2023). Our asymmetric perturbation is compatible with both representations. Notably, our theoretical results in Theorem 4.1 and Corollary 4.2 hold under the sequence-form, and we empirically demonstrate strong performance under the behavioral-form as well.
7 Conclusion
This paper introduces an asymmetric perturbation technique for solving saddle-point optimization problems, addressing key challenges in learning dynamics and equilibrium computation. Unlike symmetric perturbation methods that yield only approximate equilibria, our approach guarantees last-iterate convergence to the exact equilibrium without requiring parameter adjustment. This structural insight opens promising directions for future research, including the development of learning algorithms that further exploit asymmetric perturbation to enhance stability and convergence.
Appendix A Experimental details and additional experimental results
A.1 Information on the computer resources
All experiments in this paper were conducted on macOS Sonoma 14.4.1 with Apple M2 Max and 32GB RAM.
A.2 Proximity to equilibrium under symmetric and asymmetric perturbations in Biased Matching Pennies
This section investigates the proximity of $x^{\mu}$ to the equilibrium $x^{\ast}$ in the Biased Matching Pennies (BMP) game under the symmetric/asymmetric payoff perturbation, with varying perturbation strength $\mu$. The game matrix for BMP is provided in Table 1.
Table 1: Game matrix in BMP
$y_{1}$ $y_{2}$
$x_{1}$ $1/3$ $-2/3$
$x_{2}$ $-2/3$ $1$
BMP has a unique equilibrium $x^{\ast}=y^{\ast}=(\frac{5}{8},\frac{3}{8})$, and the game value is given as $v^{\ast}=-\frac{1}{24}$.
Figure 5 exhibits the proximity of $x^{\mu}$ to $x^{\ast}$ as $\mu$ varies. Notably, under the symmetric perturbation, $x^{\mu}$ coincides with $x^{\ast}$ when $\mu$ is set to $\mu=\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}=\frac{4}{3}$. This result underscores the statement in Theorem 3.1, that $x^{\mu}$ does not coincide with $x^{\ast}$ as long as $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$.
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 5: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix $A$ is given by $[[\frac{1}{3},-\frac{2}{3}],[-\frac{2}{3},1]]$.
A.3 Additional experiments in normal-form games
In this section, we experimentally compare our AsymP-GDA with SymP-GDA, GDA, and OGDA [Daskalakis et al., 2018, Wei et al., 2021]. We conduct experiments on two normal-form games: Biased Rock-Paper-Scissors (BRPS) and Multiple Nash Equilibria (M-Ne). These games are taken from Abe et al. [2023] and Wei et al. [2021]. Tables 3 and 3 provide the game matrices for BRPS and M-Ne, respectively.
Table 2: Game matrix in BRPS
$y_{1}$ $y_{2}$ $y_{3}$
$x_{1}$ 00 $1$ $-3$
$x_{2}$ $-1$ 00 $1$
$x_{3}$ $3$ $-1$ 00
Table 3: Game matrix in M-Ne
$y_{1}$ $y_{2}$ $y_{3}$ $y_{4}$ $y_{5}$
$x_{1}$ 00 $-1$ $1$ 00 00
$x_{2}$ $1$ 00 $-1$ 00 00
$x_{3}$ $-1$ $1$ 00 00 00
$x_{4}$ $-1$ $1$ 00 $2$ $-1$
$x_{5}$ $-1$ $1$ 00 $-1$ $2$
Figure 6 illustrates the logarithm of NashConv averaged over $100$ different random seeds. For each random seed, the initial strategies $(x^{0},y^{0})$ are chosen uniformly at random within the strategy spaces $\mathcal{X}=\Delta^{m}$ and $\mathcal{Y}=\Delta^{n}$. We use a learning rate of $\eta=0.01$ for each algorithm, and a perturbation strength of $\mu=1$ for both AsymP-GDA and SymP-GDA. We observe that AsymP-GDA converges to the minimax strategy in the original game, while SymP-GDA converges to a point far from the minimax strategy.
(a) BRPS
(b) M-Ne
Figure 6: Performance of AsymP-GDA, SymP-GDA, GDA, and OGDA in normal-form games. The shaded area represents the standard errors.
Figure 7: Trajectories for SymP-GDA (top row) and AsymP-GDA (bottom row) under different perturbation strengths $\mu\in\{0.5,1.0,4.0\}$ in BRPS.
Figure 7 illustrates the trajectories of SymP-GDA (top row) and AsymP-GDA (bottom row) under varying perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS. For SymP-GDA, the trajectories do not converge directly to the equilibrium even for small values of $\mu=0.5,1.0$. Instead, they follow circuitous and elongated paths, resulting in slower convergence. Conversely, as $\mu$ increases ($\mu=2.0,4.0$), the trajectories become more direct, leading to faster convergence, but it remains farther from the equilibrium. In contrast, AsymP-GDA leads to direct convergence to the equilibrium with small perturbation strengths. For $\mu$ values up to $2.0$ the trajectories converge directly to the equilibrium. However, as $\mu$ increases beyond a threshold ($\mu=4.0$), the trajectory deviates from the equilibrium. These results provide a more detailed understanding of the trends observed in Figures 1 and 1, further illustrating the differences in convergence dynamics between symmetric and asymmetric perturbations.
A.4 Pseudo Code for Experiments in Extensive-Form Games
In this section, we present the pseudocode for AsymP-CFR+ in Algorithm 1.
Algorithm 1 AsymP-CFR+
1:  $(\boldsymbol{R}^{x}_{t},\boldsymbol{R}^{y}_{t})=0$
2:  for $t=0$ to $T$ do
3:     $\boldsymbol{x}_{t}=\left[\boldsymbol{R}^{x}_{t}\right]^{+}\big{/}\|\left[\boldsymbol{R}^{x}_{t}\right]^{+}\|_{1}$
4:     $\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu\boldsymbol{x}_{t}\right]^{+}$
5:     $\boldsymbol{y}_{t}=\left[\boldsymbol{R}^{y}_{t}\right]^{+}\big{/}\|\left[\boldsymbol{R}^{y}_{t}\right]^{+}\|_{1}$
6:     $\boldsymbol{R}^{y}_{t+1}=\left[\boldsymbol{R}^{y}_{t}+\boldsymbol{r}({y}_{t})\right]^{+}$
7:  end for
Appendix B Independently Perturbed Game
Let us consider the perturbed game where players $x$ and $y$ choose independently their perturbation strengths $\mu_{x}$ and $\mu_{y}$:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu_{x}}{2}\left\|x\right\|^{2}-\frac{\mu_{y}}{2}\left\|y\right\|^{2}\right\}.$ (11)
We establish a theoretical result similar to Theorem 3.3 for this perturbed game.
Theorem B.1.
Assume that the original game is a normal-form game with a unique interior equilibrium, and that $(x^{\ast},y^{\ast})\neq\left(\frac{\mathbf{1}_{m}}{m},\frac{\mathbf{1}_{n}}{n}\right)$, i.e., the equilibrium is not the uniform random strategy profile. Then, for any $\mu_{x}>0$ such that $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$, the minimax strategy $x^{\mu}$ in the corresponding symmetrically perturbed game (11) satisfies $x^{\mu}\neq x^{\ast}$. Furthermore, for any $\mu>0$ such that $\mu_{y}\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the maximin strategy $y^{\mu}$ in (11) satisfies $y^{\mu}\neq y^{\ast}$.
Appendix C Adaptively Asymmetric Perturbation
As noted in Remark 4.3, AsymP-GDA may require a very small value of $\mu$ to ensure convergence to an equilibrium in the original game. To address this limitation, this section introduces an enhanced variant called Adaptively AsymP-GDA (Ada-AsymP-GDA), which incorporates an adaptive anchoring strategy $\sigma\in\mathcal{X}$ [Perolat et al., 2021, Abe et al., 2024].
In Ada-AsymP-GDA, instead of using the perturbation term $\mu x^{t}$ as in the original algorithm, we apply the gradient of the squared distance between $x^{t}$ and the anchor $\sigma$, namely $\mu(x^{t}-\sigma)$. The anchoring strategy $\sigma$ is updated periodically: it is reset to the current strategy $x^{t}$ every $T_{\sigma}$ iterations.
Let $k(t)$ denote the number of times $\sigma$ has been updated up to iteration $t$, and let $\sigma^{k(t)}$ denote the anchoring strategy after $k(t)$ updates. Since $\sigma$ is updated every $T_{\sigma}$ iterations, we have $k(t)=\lfloor(t-1)/T_{\sigma}\rfloor+1$ and $\sigma^{k(t)}=x^{T_{\sigma}(k(t)-1)+1}$. In summary, Ada-AsymP-GDA updates each player’s strategy at iteration $t\in[T]$ according to:
[…]
C.1 Last-Iterate Convergence Rate
We now establish the last-iterate convergence rate of Ada-AsymP-GDA. In particular, we show that the final strategy $x^{T+1}$ converges to an equilibrium of the original game at a rate of $\mathcal{O}(1/T)$.
Theorem C.1.
Let $\mu>0$ be an arbitrary perturbation strength. If we set $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\|A\|^{2}}{\mu}\right)}\right)$ and $T_{\sigma}=cT$ for some constant $c\leq\min\left(1,\frac{\alpha^{2}}{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{1})-x^{1}\right\|^{2}}\right)$, then the strategy $x^{T+1}$ satisfies:
$\left\|\prod_{\mathcal{X}^{\ast}}(x^{T+1})-x^{T+1}\right\|^{2}$ $\leq\mathcal{O}\left(\frac{1}{T}\right).$
Notably, Ada-AsymP-GDA achieves a last-iterate convergence rate of $\mathcal{O}(1/T)$ even when using a relatively large value of $\mu$.
C.2 Empirical performance in normal-form games
In this section, we empirically evaluate the performance of Ada-AsymP-GDA on the same normal-form games presented in Appendix A.3. We compare Ada-AsymP-GDA against AsymP-GDA, SymP-GDA, and Ada-SymP-GDA, where Ada-SymP-GDA applies perturbations $\mu(x^{t}-\sigma_{x}^{k(t)})$ and $\mu(y^{t}-\sigma_{y}^{k(t)})$ to both players’ gradients.
For all algorithms, we use a learning rate of $\eta=0.01$ and a perturbation strength of $\mu=5$. The update interval for the anchoring strategy is set to $T_{\sigma}=10,000$ for both Ada-AsymP-GDA and Ada-SymP-GDA.
Figure 8 illustrates the logarithm of NashConv averaged over $100$ random seeds. Ada-AsymP-GDA exhibits substantially faster convergence to an equilibrium compared to the other methods.
(a) BRPS
(b) M-Ne
Figure 8: Performance of Ada-AsymP-GDA in normal-form games. The shaded area represents the standard errors.
C.3 Empirical performance in extensive-form games
We also extend our method to extensive-form games through a variant called Adaptive AsymP-CFR+ (Ada-AsymP-CFR+), which integrates the adaptive anchoring strategy into AsymP-CFR+. Specifically, in Ada-AsymP-CFR+, the cumulative counterfactual regret $\boldsymbol{R}_{t}^{x}$ for player $x$ is updated as:
$\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu(\boldsymbol{x}_{t}-\sigma^{k(t)})\right]^{+}.$
We compare Ada-AsymP-CFR+ with AsymP-CFR+, SymP-CFR+, and Ada-SymP-CFR+ in the same extensive-form games in Section 5. Note that Ada-SymP-CFR+ applies symmetric perturbations with the adaptive anchoring strategy. For all algorithms, we use a perturbation strength of $\mu=0.05$. The update interval for the anchoring strategy is set to $T_{\sigma}=2,500$ for both Ada-AsymP-CFR+ and Ada-SymP-CFR+.
Figure 9 illustrates the NashConv values for each game. The results indicate that Ada-AsymP-CFR+ consistently achieves lower NashConv values, demonstrating superior convergence performance.
Figure 9: Performance of Ada-AsymP-CFR+ in extensive-form games.
Appendix D Proof of Theorem 3.1
Proof of Theorem 3.1.
First, we prove that $y^{\mu}\neq y^{\ast}$ under the assumption that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}$ by contradiction. We assume that $y^{\mu}=y^{\ast}$. Since $(x^{\ast},y^{\ast})$ is in the interior of $\Delta^{m}\times\Delta^{n}$, we have:
[…]
Then, from (13), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}$ $=x^{\top}Ay^{\ast}+\frac{\mu}{2}\left\|x\right\|^{2}=v^{\ast}+\frac{\mu}{2}\left\|x\right\|^{2}.$ (14)
On the other hand,
$\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\ast}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=v^{\ast}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$ (15)
By combining (14) and (15), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}\geq\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$
Hence, from the property of the player $x$’s equilibrium strategy in the perturbed game, $x^{\mu}$ must satisfy $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$.
On the other hand, from the property of the player $y$’s equilibrium strategy $y^{\mu}$ in the perturbed game, $y^{\mu}$ is an optimal solution of the following optimization problem:
$\max_{y\in\Delta^{n}}\left\{(x^{\mu})^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$
Let us define the following Lagrangian function $L(y,\kappa,\lambda)$ as:
$L(y,\kappa,\lambda)=(x^{\mu})^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}-\sum_{i=1}^{n}\kappa_{i}g_{i}(y)-\lambda h(y),$
where $g_{i}(y)=-y_{i}$ and $h(y)=\sum_{i=1}^{n}y_{i}-1$. Then, from the KKT conditions, we get the stationarity:
$A^{\top}x^{\mu}-\mu y^{\mu}-\sum_{i=1}^{n}\kappa_{i}\nabla g_{i}(y^{\mu})-\lambda\nabla h(y^{\mu})=\mathbf{0}_{n},$ (16)
and the complementary slackness:
$\forall i\in[n],~{}\kappa_{i}g_{i}(y^{\mu})=0.$ (17)
Since $y^{\mu}=y^{\ast}$ and $y^{\ast}$ is in the interior of $\Delta^{n}$, we have $g(y^{\mu})=-y_{i}^{\mu}<0$ for all $i\in[n]$. Thus, from (17), we have $\kappa_{i}=0$ for all $i\in[n]$. Substituting this into (16), we obtain:
$A^{\top}x^{\mu}-\mu y^{\mu}-\lambda\nabla h(y^{\mu})=A^{\top}x^{\mu}-\mu y^{\mu}-\lambda\mathbf{1}_{n}=\mathbf{0}_{n}.$ (18)
Hence, we have:
$\lambda=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}.$ (19)
Putting (19) into (18) yields:
$y^{\mu}=\frac{1}{\mu}\left(A^{\top}x^{\mu}-\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}\mathbf{1}_{n}\right)=\frac{1}{\mu}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-\frac{\mathbf{1}_{n}^{\top}A^{\top}\frac{\mathbf{1}_{m}}{m}-\mu}{n}\mathbf{1}_{n}\right),$
where the second equality follows from $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$. Multiplying this by $\frac{\mathbf{1}_{m}^{\top}}{m}A$, we have:
$v^{\ast}=\frac{1}{\mu}\left(\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}\right),$ (20)
where we used the assumption that $y^{\mu}=y^{\ast}$ and (13). Here, we have:
$\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}$
$=\left\|v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=n(v^{\ast})^{2}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}+2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$\phantom{=}-n(v^{\ast})^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}-2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\mu v^{\ast}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu v^{\ast}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu v^{\ast}+\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right).$ (21)
Here, since $A^{\top}\frac{\mathbf{1}_{m}}{m}=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}\mathbf{1}_{n}+\mu y^{\mu}=\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}\mathbf{1}_{n}+\mu y^{\ast}$ from (18) and (19), we get:
$\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)$
$\phantom{=}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)$
$=\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)+\frac{\mu^{2}}{n}+\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$\phantom{=}+\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\mathbf{1}_{n}+\mu^{2}(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$=\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)+\mu^{2}\left\|y^{\ast}\right\|^{2}$
$=\mu\left(\mu\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}\right).$ (22)
By combining (20), (21), and (22), we have:
$\mu\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}=0.$
Therefore, if $y^{\mu}=y^{\ast}$, then $\mu$ must satisfy:
$\mu=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)},$
and this is equivalent to:
$\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}\Rightarrow y^{\mu}\neq y^{\ast}.$
By a similar argument, in terms of player $x$, we can conclude that:
$\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left(\left\|x^{\ast}\right\|^{2}-\frac{1}{m}\right)}\Rightarrow x^{\mu}\neq x^{\ast}.$
Appendix E Proof of Theorem 3.3
Proof of Theorem 3.3.
Let us define the function $g_{\mathrm{asym}}^{\mu}:\mathcal{X}\to\mathbb{R}$:
$g_{\mathrm{asym}}^{\mu}(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}.$
Here, we introduce the following property of the function $\max_{y\in\mathcal{Y}}x^{\top}Ay$:
Lemma E.1 (Claim 1-5 in Theorem 5 of Wei et al. [2021]).
There exists a positive constant $\alpha>0$ such that:
$\forall x\in\mathcal{X},\max_{y\in\mathcal{Y}}x^{\top}Ay-v^{\ast}\geq\alpha\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|,$
where $\alpha$ depends only on $\mathcal{X}^{\ast}$.
By using Lemma E.1, we have for any $x\in\mathcal{X}$:
$\max_{y\in\mathcal{Y}}\left(\prod_{\mathcal{X}^{\ast}}(x)\right)^{\top}Ay-\max_{y\in\mathcal{Y}}x^{\top}Ay=v^{\ast}-\max_{y\in\mathcal{Y}}x^{\top}Ay\leq-\alpha\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|.$
On the other hand, for any $x^{\ast}\in\mathcal{X}^{\ast}$ and $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$, we have:
$\frac{\mu}{2}\left\|x^{\ast}\right\|^{2}-\frac{\mu}{2}\left\|x\right\|^{2}$ $=\frac{\mu}{2}\langle x^{\ast}-x,x^{\ast}+x\rangle$
$=\frac{\mu}{2}\langle x^{\ast}-x,-x^{\ast}+x+2x^{\ast}\rangle$
$\leq\mu\langle x^{\ast}-x,x^{\ast}\rangle$
$\leq\mu\left\|x^{\ast}-x\right\|\left\|x^{\ast}\right\|.$
Summing up these inequalities, we obtain:
$g_{\mathrm{asym}}^{\mu}\left(\prod_{\mathcal{X}^{\ast}}(x)\right)-g_{\mathrm{asym}}^{\mu}(x)$ $\leq-\alpha\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|+\mu\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|\left\|\prod_{\mathcal{X}^{\ast}}(x)\right\|$
$=-\left(\alpha-\mu\left\|\prod_{\mathcal{X}^{\ast}}(x)\right\|\right)\left\|x-\prod_{\mathcal{X}^{\ast}}(x)\right\|.$
Hence, under the assumption that $\mu<\frac{\alpha}{\max_{x^{\ast}\in\mathcal{X}^{\ast}}\left\|x^{\ast}\right\|}$, we have for any $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$:
$g_{\mathrm{asym}}^{\mu}\left(\prod_{\mathcal{X}^{\ast}}(x)\right)<g_{\mathrm{asym}}^{\mu}(x).$
Thus, every $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$ is dominated by some equilibrium strategy $x^{\ast}\in\mathcal{X}^{\ast}$ with respect to the value of $g_{\mathrm{asym}}^{\mu}$. Therefore, we conclude that:
$x^{\mu}\in\mathop{\rm arg~{}min}\limits_{x\in\mathcal{X}}g_{\mathrm{asym}}^{\mu}(x)\Leftrightarrow x^{\mu}\in\mathop{\rm arg~{}min}\limits_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}\right\}\Leftrightarrow x^{\mu}\in\mathcal{X}^{\ast}.$
Appendix F Proofs for Theorem 4.1
F.1 Proof of Theorem 4.1
Proof of Theorem 4.1.
First, we have for any vectors $a,b,c$:
$\frac{1}{2}\left\|a-b\right\|^{2}-\frac{1}{2}\left\|a-c\right\|^{2}+\frac{1}{2}\left\|b-c\right\|^{2}=\langle c-b,a-b\rangle.$ (23)
From (23), we have for any $t\geq 1$:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}=\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle$ (24)
Here, we can rewrite the update rule as follows:
$x^{t+1}&=\mathop{\rm arg~{}min}\limits_{p\in\mathcal{X}}\left\{\eta\left\langle Ay^{t}+\mu x^{t},p\right\rangle+\frac{1}{2}\left\|p-x^{t}\right\|^{2}\right\},\\ y^{t+1}&=\mathop{\rm arg~{}min}\limits_{p\in\mathcal{Y}}\left\{-\eta\left\langle A^{\top}x^{t+1},p\right\rangle+\frac{1}{2}\left\|p-y^{t}\right\|^{2}\right\}.\end{aligned}$ (25)
From the first-order optimality condition for $x^{t+1}$ in (25), we have for any $t\geq 1$:
$\left\langle\eta Ay^{t}+\eta\mu x^{t}+x^{t+1}-x^{t},x^{t+1}-x^{\mu}\right\rangle\leq 0.$ (26)
Combining (24) and (26) yields:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}$
$\leq$ $\eta\left\langle Ay^{t}+\mu x^{t},x^{\mu}-x^{t+1}\right\rangle$
$=$ $\eta\left\langle Ay^{t+1}+\mu x^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1}+\mu(x^{t}-x^{t+1}),x^{\mu}-x^{t+1}\right\rangle$
$=$ $\eta\left\langle Ay^{\mu}+\mu x^{\mu},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1}+\mu(x^{t}-x^{t+1}),x^{\mu}-x^{t+1}\right\rangle$
$+\eta\left\langle Ay^{t+1}-Ay^{\mu},x^{\mu}-x^{t+1}\right\rangle-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}.$ (27)
On the other hand, from the first-order optimality condition for $x^{\mu}$ in (4), we get:
$\left\langle Ay^{\mu}+\mu x^{\mu},x^{\mu}-x^{t+1}\right\rangle\leq 0.$ (28)
By combining (27) and (28), we have for any $t\geq 1$:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}$
$\leq$ $-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\eta\left\langle Ay^{t}-Ay^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\mu\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle$
$+\eta\left\langle Ay^{t+1}-Ay^{\mu},x^{\mu}-x^{t+1}\right\rangle.$ (29)
Similar to (24), we have for any $t\geq 1$:
$\frac{1}{2}\left\|y^{\mu}-y^{t+1}\right\|^{2}-\frac{1}{2}\left\|y^{\mu}-y^{t}\right\|^{2}+\frac{1}{2}\left\|y^{t+1}-y^{t}\right\|^{2}=\left\langle y^{t}-y^{t+1},y^{\mu}-y^{t+1}\right\rangle,$ (30)
and from the first-order optimality condition for $y^{t+1}$ in (25), we have for any $t\geq 1$:
$\left\langle-\eta A^{\top}x^{t+1}+y^{t+1}-y^{t},y^{t+1}-y^{\mu}\right\rangle\leq 0.$ (31)
By combining (30) and (31), we get:
$\frac{1}{2}\left\|y^{\mu}-y^{t+1}\right\|^{2}-\frac{1}{2}\left\|y^{\mu}-y^{t}\right\|^{2}+\frac{1}{2}\left\|y^{t+1}-y^{t}\right\|^{2}$
$\leq-\eta\left\langle A^{\top}x^{t+1},y^{\mu}-y^{t+1}\right\rangle$
$=-\eta\left\langle A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle-\eta\left\langle A^{\top}x^{t+1}-A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle$
$\leq-\eta\left\langle A^{\top}x^{t+1}-A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle,$ (32)
where the last inequality stems from (5).
Summing up (29) and (32), we have for any $t\geq 1$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}+\frac{1}{2}\left\|z^{t+1}-z^{t}\right\|^{2}$
$\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\eta\mu\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1},x^{\mu}-x^{t+1}\right\rangle$
$\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|x^{t}-x^{t+1}\right\|^{2}+\eta^{2}\mu^{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|y^{t}-y^{t+1}\right\|^{2}+\eta^{2}\|A\|^{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}$
$=-\left(\eta\mu-\eta^{2}(\mu^{2}+\|A\|^{2})\right)\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|z^{t}-z^{t+1}\right\|^{2}.$
Hence, under the assumption that $\eta<\frac{\mu}{2(\mu^{2}+\|A\|^{2})}$, we have for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}\leq-\frac{\eta\mu}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{4}\left\|z^{t}-z^{t+1}\right\|^{2}.$ (33)
In terms of the path length $\left\|z^{t}-z^{t+1}\right\|$, we can we derive the following lower bound:
Lemma F.1.
For any $t\geq 1$, we have:
$\left\|z^{t}-z^{t+1}\right\|\geq\frac{\eta}{D(1+\eta\|A\|)}\left(\frac{\mu}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{\beta}{2\mu}\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right),$
where $\beta>0$ is a positive constant depending only on $\mathcal{Y}^{\mu}$.
From (33) and Lemma F.1, we obtain for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}$
$\leq-\frac{\eta\mu}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{4}\left\|z^{t}-z^{t+1}\right\|^{2}$
$\leq-\frac{\eta\mu}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{\eta^{2}}{4D^{2}(1+\eta\|A\|)^{2}}\left(\frac{\mu^{2}}{4}\left\|x^{\mu}-x^{t+1}\right\|^{4}+\frac{\beta^{2}}{4\mu^{2}}\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{4}\right)$
$\leq-\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{16D^{2}(1+\eta\|A\|)^{2}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{4}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{4}\right)$
$\leq-\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{32D^{2}(1+\eta\|A\|)^{2}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right)^{2}.$
where the second inequality follows from the fact that $(a+b)^{2}\geq a^{2}+b^{2}$ for any $a,b\geq 0$, and the fourth inequality follows from the fact that $a^{2}+b^{2}\geq\frac{1}{2}(a+b)^{2}$ for any $a,b\geq 0$. Then, under the assumption that $\eta<\frac{\mu}{2(\mu^{2}+\|A\|^{2})}$, we have for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}$
$\leq-\frac{2\eta^{2}(\mu^{2}+\|A\|^{2})\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{32D^{2}(2(\mu^{2}+\|A\|^{2})+\mu\|A\|)^{2}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right)^{2}$
$\leq-\frac{\eta^{2}\left(\mu+\|A\|\right)^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{128D^{2}\left(\mu+\|A\|\right)^{4}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right)^{2}$
$=-\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{128D^{2}\left(\mu+\|A\|\right)^{2}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right)^{2}.$
By choosing $y^{\mu}=\prod_{\mathcal{Y}^{\mu}}(y^{t})$, we get for any $t\geq 1$:
$\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}$
$\leq\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t+1}\right\|^{2}$
$\leq\left\|x^{\mu}-x^{t}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t}\right\|^{2}-\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{128D^{2}\left(\mu+\|A\|\right)^{2}}\left(\left\|x^{\mu}-x^{t+1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right)^{2}.$
To obtain the convergence rate of $\left\|x^{\mu}-x^{t}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t}\right\|^{2}$, we introduce the following lemma from Wei et al. [2021]:
Lemma F.2 (Lemma 12 of Wei et al. [2021]).
Consider a non-negative sequence $\{B_{t}\}_{t\geq 1}$ that satisfies for some $q>0$,
$B_{t+1}\leq B_{t}-qB_{t+1}^{2},~{}t\geq 1$
and
$2qB_{1}\leq 1$
Then, $B_{t}\leq\frac{\max\left(B_{1},\frac{2}{q}\right)}{t}$.
Under the assumption that $\eta\leq\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta}{\mu}\right)}$:
$2\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{128D^{2}\left(\mu+\|A\|\right)^{2}}\left(\left\|x^{\mu}-x^{1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{1})-y^{1}\right\|^{2}\right)^{2}\leq\frac{\eta^{2}D^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{64\left(\mu+\|A\|\right)^{2}}\leq 1$
Therefore, we can apply Lemma F.2 with $q=\frac{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}{128D^{2}\left(\mu+\|A\|\right)^{2}}$, and then we have:
$\left\|x^{\mu}-x^{t}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t})-y^{t}\right\|^{2}$ $\leq\frac{\max\left(\left\|x^{\mu}-x^{1}\right\|^{2}+\left\|\prod_{\mathcal{Y}^{\mu}}(y^{1})-y^{1}\right\|^{2},\frac{256D^{2}\left(\mu+\|A\|\right)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}\right)}{t}$
$\leq\frac{D^{2}\max\left(1,\frac{256\left(\mu+\|A\|\right)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}\right)}{t}$
$\leq\frac{256D^{2}\left(\mu+\|A\|\right)^{2}}{\eta^{2}t\min\left(\mu^{2},\frac{\beta^{2}}{\mu^{2}}\right)}.$
F.2 Proof of Lemma F.1
Proof of Lemma F.1.
We first derive the lower bound on the path length $\left\|z^{t}-z^{t+1}\right\|$ by the suboptimality gap:
Lemma F.3.
For any $x\in\mathcal{X}$, $y^{\mu}\in\mathcal{Y}^{\mu}$, and $t\geq 1$, we have:
$\frac{\eta}{1+\eta\|A\|}\frac{\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle-\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle}{\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{\mu}\right\|}\leq\left\|z^{t+1}-z^{t}\right\|.$
We also derive the lower bound on the suboptimality gap for any strategy profile $(x,y)\in\mathcal{X}\times\mathcal{Y}$:
Lemma F.4.
For any $x\in\mathcal{X}$, $y\in\mathcal{Y}$, and $y^{\mu}\in\mathcal{Y}^{\mu}$, we have:
$\max_{x^{\prime}\in\mathcal{X}}\left(\langle Ay+\mu x,x-x^{\prime}\rangle-\langle A^{\top}x,y-y^{\mu}\rangle\right)\geq\frac{\mu}{2}\left\|x-x^{\mu}\right\|^{2}+\frac{1}{2\mu}\left\|A(y^{\mu}-y)\right\|^{2}.$
By combining Lemmas F.3, and F.4, we have:
$\left\|z^{t+1}-z^{t}\right\|$
$\geq\frac{\eta}{1+\eta\|A\|}\max_{x\in\mathcal{X}}\frac{\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle-\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle}{\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{\mu}\right\|}$
$\geq\frac{\eta}{D(1+\eta\|A\|)}\max_{x\in\mathcal{X}}\left(\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle-\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle\right)$
$\geq\frac{\eta}{D(1+\eta\|A\|)}\left(\frac{\mu}{2}\left\|x^{t+1}-x^{\mu}\right\|^{2}+\frac{1}{2\mu}\left\|A(y^{\mu}-y^{t+1})\right\|^{2}\right).$
Here, for any maximin strategy $y^{\mu}\in\mathcal{Y}^{\mu}$ in the perturbed game, $Ay^{\mu}$ can be written as $Ay^{\mu}=b^{\ast}$ by using some vector $b^{\ast}$:
Lemma F.5.
There exists a vector $b^{\ast}$ such that $b^{\ast}=Ay^{\mu}$ for all $y^{\mu}\in\mathcal{Y}^{\mu}$. Furthermore, for any strategy $\tilde{y}\in\mathcal{Y}$, we have:
$\tilde{y}\in\mathcal{Y}^{\mu}\Leftrightarrow A\tilde{y}=b^{\ast}.$
From Lemma F.5, we can rewrite $\left\|A(y^{\mu}-y^{t+1})\right\|^{2}=\left\|Ay^{t+1}-b^{\ast}\right\|^{2}$. The following lemma demonstrates that this term can be lower bounded by the distance between $y^{t+1}$ and the set of maximin strategies $\mathcal{Y}^{\mu}$:
Lemma F.6.
There exists a positive constant $\beta>0$ such that for any $y\in\mathcal{Y}$:
$\left\|Ay-b^{\ast}\right\|^{2}\geq\beta\left\|y-\prod_{\mathcal{Y}^{\mu}}(y)\right\|^{2}.$
Therefore, we conclude that:
$\left\|z^{t+1}-z^{t}\right\|\geq\frac{\eta}{D(1+\eta\|A\|)}\left(\frac{\mu}{2}\left\|x^{t+1}-x^{\mu}\right\|^{2}+\frac{\beta}{2\mu}\left\|\prod_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}\right).$
F.3 Proof of Lemma F.3
Proof of Lemma F.3.
From the first-order optimality condition for $x^{t+1}$, we have for any $x\in\mathcal{X}$ and $t\geq 1$:
$\left\langle\eta Ay^{t}+\eta\mu x^{t}+x^{t+1}-x^{t},x-x^{t+1}\right\rangle\geq 0.$
Rearranging the terms yields:
$\left\langle x^{t+1}-x^{t},x-x^{t+1}\right\rangle$
$\geq\eta\langle Ay^{t}+\mu x^{t},x^{t+1}-x\rangle$
$=\eta\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle+\eta\langle Ay^{t}-Ay^{t+1},x^{t+1}-x\rangle+\eta\mu\langle x^{t}-x^{t+1},x^{t+1}-x\rangle.$ (34)
Similarly, from the first-order optimality condition for $y^{t+1}$, we have for any $t\geq 1$:
$\left\langle-\eta A^{\top}x^{t+1}+y^{t+1}-y^{t},y^{\mu}-y^{t+1}\right\rangle\geq 0.$
Rearranging the terms yields:
$\left\langle y^{t+1}-y^{t},y^{\mu}-y^{t+1}\right\rangle$ $\geq-\eta\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle.$ (35)
By combining (34) and (35) and Cauchy-Schwarz inequality, we have for any $t\geq 1$:
$\eta\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle-\eta\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle$
$\leq\eta\langle Ay^{t}-Ay^{t+1},x-x^{t+1}\rangle+(1-\eta\mu)\langle x^{t}-x^{t+1},x^{t+1}-x\rangle+\left\langle y^{t+1}-y^{t},y^{\mu}-y^{t+1}\right\rangle$
$\leq\eta\|A\|\left\|y^{t}-y^{t+1}\right\|\left\|x-x^{t+1}\right\|+(1-\eta\mu)\left\|x^{t}-x^{t+1}\right\|\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{t}\right\|\left\|y^{\mu}-y^{t+1}\right\|$
$\leq\left(\eta\|A\|\left\|y^{t}-y^{t+1}\right\|+(1-\eta\mu)\left\|x^{t}-x^{t+1}\right\|\right)\left\|x-x^{t+1}\right\|+\left\|y^{t+1}-y^{t}\right\|\left\|y^{\mu}-y^{t+1}\right\|.$
Therefore, we have for any $t\geq 1$:
$\eta\frac{\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle-\langle A^{\top}x^{t+1},y^{t+1}-y^{\mu}\rangle}{\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{\mu}\right\|}$
$\leq\frac{\left(\eta\|A\|\left\|y^{t}-y^{t+1}\right\|+(1-\eta\mu)\left\|x^{t}-x^{t+1}\right\|\right)\left\|x-x^{t+1}\right\|}{\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{\mu}\right\|}+\frac{\left\|y^{t+1}-y^{t}\right\|\left\|y^{\mu}-y^{t+1}\right\|}{\left\|x^{t+1}-x\right\|+\left\|y^{t+1}-y^{\mu}\right\|}$
$\leq\eta\|A\|\left\|y^{t}-y^{t+1}\right\|+(1-\eta\mu)\left\|x^{t}-x^{t+1}\right\|+\left\|y^{t+1}-y^{t}\right\|$
$\leq(1+\eta\|A\|)\left\|z^{t}-z^{t+1}\right\|.$
F.4 Proof of Lemma F.4
Proof of Lemma F.4.
From the strongly convexity of the function $\frac{1}{2}\left\|x\right\|^{2}$, we have for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $x^{\prime}\in\mathcal{X}$:
$x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}-(x^{\prime})^{\top}Ay-\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\leq\langle Ay+\mu x,x-x^{\prime}\rangle-\frac{\mu}{2}\left\|x-x^{\prime}\right\|^{2}.$
Moreover, from the linearity of the function $x^{\top}Ay$ with respect to $y\in\mathcal{Y}$, we get for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $y^{\mu}\in\mathcal{Y}^{\mu}$:
$x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}-x^{\top}Ay-\frac{\mu}{2}\left\|x\right\|^{2}=\langle A^{\top}x,y^{\mu}-y\rangle.$
Summing up the above two inequalities, we get for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\langle Ay+\mu x,x-x^{\prime}\rangle-\langle A^{\top}x,y-y^{\mu}\rangle$
$\geq x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}-(x^{\prime})^{\top}Ay-\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}+\frac{\mu}{2}\left\|x-x^{\prime}\right\|^{2}$
$\geq x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}-\left((x^{\mu})^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\right)$
$\phantom{=}+\left((x^{\mu})^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\right)-\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)$
$\geq\frac{\mu}{2}\left\|x-x^{\mu}\right\|^{2}+\left((x^{\mu})^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\right)-\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right).$
Hence, we obtain:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle-\langle A^{\top}x,y-y^{\mu}\rangle$
$\geq\frac{\mu}{2}\left\|x-x^{\mu}\right\|^{2}+\left((x^{\mu})^{\top}Ay^{\mu})+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)$
$=\frac{\mu}{2}\left\|x-x^{\mu}\right\|^{2}+\max_{\tilde{y}\in\mathcal{Y}}\min_{\tilde{x}\in\mathcal{X}}\left((\tilde{x})^{\top}A\tilde{y}+\frac{\mu}{2}\left\|\tilde{x}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right).$
Here, the suboptimality gap of the strategy for player $x$ can be lower bounded as follows:
Lemma F.7.
For any $y\in\mathcal{Y}$ and $y^{\mu}\in\mathcal{Y}^{\mu}$, we have:
$\max_{y^{\prime}\in\mathcal{Y}}\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay^{\prime}+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)\geq\frac{1}{2\mu}\left\|A(y^{\mu}-y)\right\|^{2}.$
Combining this inequality with Lemma F.7, we have:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle-\langle A^{\top}x,y-y^{\mu}\rangle\geq\frac{\mu}{2}\left\|x-x^{\mu}\right\|^{2}+\frac{1}{2\mu}\left\|A(y^{\mu}-y)\right\|^{2}.$
F.5 Proof of Lemma F.5
Proof of Lemma F.5.
Let us define :=$\mathcal{B}:=\left\{-Ay~{}|~{}y\in\mathcal{Y}\right\}$. We also define $h(x):=\frac{\mu}{2}\left\|x\right\|^{2}$ and its convex conjugate $h^{\ast}(p):=\max_{x\in\mathcal{X}}\left\{x^{\top}p-h(x)\right\}$. Since $h$ is a $\mu$-smooth function, the convex conjugate $h^{\ast}$ is $\frac{1}{\mu}$-strongly convex [Shalev-Shwartz, 2011, p. 149]. Thus, $\mathop{\rm arg~{}min}\limits_{b\in\mathcal{B}}h^{\ast}(b)$ is a singleton, and then we can write $b^{\ast}=\mathop{\rm arg~{}min}\limits_{b\in\mathcal{B}}h^{\ast}(b)$. Hence, every $\tilde{y}\in\mathcal{Y}^{\mu}$ must satisfy $A\tilde{y}=b^{\ast}$. Thus, we have $\tilde{y}\in\mathcal{Y}^{\mu}\Rightarrow A\tilde{y}=b^{\ast}$.
Next, we show that $A\tilde{y}=b^{\ast}\Rightarrow\tilde{y}\in\mathcal{Y}^{\mu}$. We have for any $\tilde{y}\in\mathcal{Y}$ satisfying $A\tilde{y}=b^{\ast}$:
$\forall y\in\mathcal{Y},~{}(x^{\mu})^{\top}A\tilde{y}=(x^{\mu})^{\top}b^{\ast}\geq(x^{\mu})^{\top}Ay,$
since $(x^{\mu})^{\top}b^{\ast}=(x^{\mu})^{\top}Ay^{\mu}\geq(x^{\mu})^{\top}Ay$. Thus, $\tilde{y}$ is the best response against $x^{\mu}$ in the perturbed game. Moreover, since $Ay^{\mu}=b^{\ast}$ for any $y^{\mu}\in\mathcal{Y}^{\mu}$ and $x^{\mu}$ is an equilibrium in the perturbed game, we have:
$\forall x\in\mathcal{Y},~{}(x^{\mu})^{\top}b^{\ast}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\leq x^{\top}b^{\ast}+\frac{\mu}{2}\left\|x\right\|^{2}.$
Hence, since $b^{\ast}=A\tilde{y}$, we have:
$\forall x\in\mathcal{X},~{}(x^{\mu})^{\top}A\tilde{y}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\leq x^{\top}A\tilde{y}+\frac{\mu}{2}\left\|x\right\|^{2},$
and then $x^{\mu}$ is also a best response against $\tilde{y}$. Consequently, it holds that $(x^{\mu},\tilde{y})$ is an equilibrium in the perturbed game, and $\tilde{y}\in\mathcal{Y}^{\mu}$. ∎
F.6 Proof of Lemma F.6
Proof of Lemma F.6.
Since the statement clearly holds for $y\in\mathcal{Y}^{\mu}$, we consider an arbitrary strategy $y\in\mathcal{Y}\setminus\mathcal{Y}^{\mu}$. First, we have for any convex set $\mathcal{S}$:
$s^{\prime}-s\in\mathcal{N}_{\mathcal{S}}(s)\Leftrightarrow s=\prod_{\mathcal{S}}(s^{\prime}).$
Thus, defining $y^{\mu}=\prod_{\mathcal{Y}^{\mu}}(y)$, we get:
$\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})=\left\{\tilde{y}-\tilde{y}^{\mu}~{}|~{}\tilde{y}\in\mathbb{R}^{n},~{}\tilde{y}^{\mu}=\prod_{\mathcal{Y}^{\mu}}(\tilde{y})\right\}.$ (36)
On the other hand, since $\mathcal{Y}$ is a polytope and also a polyhedron, we can write $\mathcal{Y}=\left\{y\in\mathbb{R}^{n}~{}|~{}\langle d_{i},y\rangle\leq e_{i}\mbox{ for }i=1,\cdots,L\right\}$ for some $\{d_{i}\}_{i=1}^{L}$ and $\{e_{i}\}_{i=1}^{L}$. Hence, from Lemma F.5, $\mathcal{Y}^{\mu}$ can be represented as:
$\mathcal{Y}^{\mu}=\left\{y\in\mathbb{R}^{n}~{}|~{}\langle d_{i},y\rangle\leq e_{i}\mbox{ for }i=1,\cdots,L,~{}\langle a_{i},y\rangle=b^{\ast}_{i}\mbox{ for }i=1,\cdots,m\right\},$
where $A=(a_{1},\cdots,a_{m})^{\top}$. Thus, $\mathcal{Y}^{\mu}$ is also a polytope. Hence, from Theorem 6.46 in Rockafellar and Wets [2009], $\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})$ can be written as:
$\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})=$$\left\{\sum_{i=1}^{L}p_{i}d_{i}+\sum_{i=1}^{m}q_{i}a_{i}-\sum_{i=1}^{m}r_{i}a_{i}~{}|~{}p_{i}\geq 0\mbox{ for }i\in I(y^{\mu}),~{}p_{i}=0\mbox{for }i\notin I(y^{\mu}),\right.$
$\left.~{}q_{i}\geq 0\mbox{ for }i\in I^{\ast}(y^{\mu}),q_{i}=0\mbox{ for }i\notin I^{\ast}(y^{\mu}),\right.$
$\left.~{}r_{i}\geq 0\mbox{ for }i\in I^{\ast}(y^{\mu}),~{}r_{i}=0\mbox{ for }i\notin I^{\ast}(y^{\mu})\right\},$
where $I(y^{\mu}):=\left\{i\in[1,L]~{}|~{}\langle d_{i},y^{\mu}\rangle=e_{i}\right\}$ and $I^{\ast}(y^{\mu}):=\left\{i\in[1,m]~{}|~{}\langle a_{i},y^{\mu}\rangle=b^{\ast}_{i}\right\}$. Since $y^{\mu}\in\mathcal{Y}^{\mu}$, it holds that $i\in I^{\ast}(y^{\mu})$ for all $i\in[1,m]$. Furthermore, without loss of generality, we assume that $\langle d_{i},y^{\mu}\rangle=e_{i}$ for $i=1,\cdots,l$ and $\langle d_{i},y^{\mu}\rangle<e_{i}$ for $i=l+1,\cdots,L$ with some $l\leq L$. Then, we obtain:
$\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})=\left\{\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}~{}|~{}p_{i}\geq 0,~{}q_{i}\geq 0,~{}r_{i}\geq 0\right\}.$ (37)
Combining (36) and (37):
$\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})$ =$=\left\{\tilde{y}-\tilde{y}^{\mu}~{}|~{}\tilde{y}\in\mathbb{R}^{n},~{}\tilde{y}^{\mu}=\prod_{\mathcal{Y}^{\mu}}(\tilde{y})\right\}$
=$=\left\{\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}~{}|~{}p_{i}\geq 0,~{}q_{i}\geq 0,~{}r_{i}\geq 0\right\}.$ (38)
Since $y^{\mu}\in\prod_{\mathcal{Y}^{\mu}}(y)$, we have $y-y^{\mu}\in\mathcal{N}_{\mathcal{Y}^{\mu}}(y^{\mu})$. Thus, we can write $y-y^{\mu}=\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}$ with $p_{i}\geq 0$, $q_{i}\geq 0$, and $r_{i}\geq 0$. Moreover, we have for any $i=1,\cdots,l$:
$\langle d_{i},y-y^{\mu}\rangle=\langle d_{i},y\rangle-e_{i}\leq 0.$
Hence, from (38), we obtain:
$y-y^{\mu}\in\mathcal{M}(y^{\mu})$ :=$:=\left\{\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}~{}|~{}p_{i}\geq 0,~{}q_{i}\geq 0,~{}r_{i}\geq 0,\right.$
$\left.~{}\left\langle d_{i},\sum_{j=1}^{l}p_{j}d_{j}+\sum_{j=1}^{m}(q_{j}-r_{j})a_{j}\right\rangle\leq 0\mbox{ for }i\in[1,l]\right\}.$
Since $\mathcal{M}(y^{\mu})$ is a cone, we also have $\frac{y-y^{\mu}}{\left\|y-y^{\mu}\right\|}\in\mathcal{M}(y^{\mu})$. Then, we have:
$\frac{y-y^{\mu}}{\left\|y-y^{\mu}\right\|}\in\mathcal{P}(y^{\mu})$ :=$:=\left\{\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}~{}|~{}p_{i}\geq 0,~{}q_{i}\geq 0,~{}r_{i}\geq 0,\right.$
$\left.~{}\left\langle d_{i},\sum_{j=1}^{l}p_{j}d_{j}+\sum_{j=1}^{m}(q_{j}-r_{j})a_{j}\right\rangle\leq 0\mbox{ for }i\in[1,l],\right.$
$\left.~{}\left\|\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}\right\|_{\infty}\leq 1\right\}.$
Since $\mathcal{P}(y^{\mu})$ is a bounded polyhedron, it is also a polytope. Now, let us take an arbitrary vertex $v$ of $\mathcal{P}(y^{\mu})$ and consider the following optimization problem:
$\min_{p_{i},q_{i},r_{i},C_{v}}C_{v},\mbox{ s.t., }v=\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i},~{}0\leq p_{i}\leq C_{v},0\leq q_{i}\leq C_{v},0\leq r_{i}\leq C_{v}.$
Since $v\in\mathcal{M}(y^{\mu})$, this optimization problem is always feasible and admits a finite solution $C_{v}<\infty$. Let us denote the optimal value of this optimization problem as $C_{v}^{\ast}$, and define $C_{y^{\mu}}=\max_{v\in\mathcal{V}(\mathcal{P}(y^{\mu}))}C_{v}^{\ast}$, where $\mathcal{V}(\mathcal{P}(y^{\mu}))$ is the set of all vertices of $\mathcal{P}(y^{\mu})$. Since any $v\in\mathcal{P}(y^{\mu})$ can be expressed as a convex combination of some points in $\mathcal{V}(\mathcal{P}(y^{\mu}))$, $v$ can be written as $\sum_{i=1}^{l}p_{i}a_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}$ with $0\leq p_{i}\leq C_{y^{\mu}}$, $0\leq q_{i}\leq C_{y^{\mu}}$, and $0\leq r_{i}\leq C_{y^{\mu}}$. Thus, since $\frac{y-y^{\mu}}{\left\|y-y^{\mu}\right\|}\in\mathcal{P}(y^{\mu})$, $\frac{y-y^{\mu}}{\left\|y-y^{\mu}\right\|}$ can be expressed as $\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}$ with $0\leq p_{i}\leq C_{y^{\mu}}$, $0\leq q_{i}\leq C_{y^{\mu}}$, and $0\leq r_{i}\leq C_{y^{\mu}}$.
From the above discussion, we can write $y-y^{\mu}$ as $\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}$ with $0\leq p_{i}\leq C_{y^{\mu}}\left\|y-y^{\mu}\right\|$, $0\leq q_{i}\leq C_{y^{\mu}}\left\|y-y^{\mu}\right\|$, and $0\leq r_{i}\leq C_{y^{\mu}}\left\|y-y^{\mu}\right\|$. Then,
$\sum_{i=1}^{l}p_{i}\langle d_{i},y-y^{\mu}\rangle+\sum_{i=1}^{m}(q_{i}-r_{i})\langle a_{i},y-y^{\mu}\rangle$
$=\left\langle\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i},y-y^{\mu}\right\rangle=\left\|y-y^{\mu}\right\|^{2}.$ (39)
Here, since $y-y^{\mu}\in\mathcal{M}(y^{\mu})$ and $y-y^{\mu}=\sum_{i=1}^{l}p_{i}d_{i}+\sum_{i=1}^{m}(q_{i}-r_{i})a_{i}$, we have:
$\sum_{i=1}^{l}p_{i}\langle d_{i},y-y^{\mu}\rangle=\sum_{i=1}^{l}p_{i}\left\langle d_{i},\sum_{j=1}^{l}p_{j}d_{j}+\sum_{j=1}^{m}(q_{j}-r_{j})a_{j}\right\rangle\leq 0.$ (40)
Moreover, since $0\leq q_{i}\leq C_{y^{\mu}}\left\|y-y^{\mu}\right\|$ and $0\leq r_{i}\leq C_{y^{\mu}}\left\|y-y^{\mu}\right\|$, we have:
$\sum_{i=1}^{m}(q_{i}-r_{i})\langle a_{i},y-y^{\mu}\rangle$ $\leq\left(\max_{i\in[1,m]}\left|\langle a_{i},y-y^{\mu}\rangle\right|\right)\sum_{i=1}^{m}|q_{i}-r_{i}|$
$\leq m\left(\max_{i\in[1,m]}\left|\langle a_{i},y-y^{\mu}\rangle\right|\right)C_{y^{\mu}}\left\|y-y^{\mu}\right\|.$ (41)
Combining (39), (40), and (41), we obtain:
$\left\|y-y^{\mu}\right\|\leq m\left(\max_{i\in[1,m]}\left|\langle a_{i},y-y^{\mu}\rangle\right|\right)C_{y^{\mu}}.$
Moreover,
$\left\|A(y-y^{\mu})\right\|^{2}=\sum_{i=1}^{m}\langle a_{i},y-y^{\mu}\rangle^{2}\geq\left(\max_{i\in[1,m]}\left|\langle a_{i},y-y^{\mu}\rangle\right|\right)^{2}.$
Thus, we finally get:
$\left\|A(y-y^{\mu})\right\|\geq\frac{\left\|y-y^{\mu}\right\|}{mC_{y^{\mu}}}.$
Note that $C_{y^{\mu}}$ only depends on the set of tight constraints $I(y^{\mu})$, and there are only finitely many different sets of tight constraints. Therefore, we conclude that there exists a positive constant $\beta>0$ such that:
$\forall y\in\mathcal{Y}\setminus\mathcal{Y}^{\mu},~{}\left\|Ay-b^{\ast}\right\|^{2}\geq\beta\left\|y-\prod_{\mathcal{Y}^{\mu}}(y)\right\|^{2}$
F.7 Proof of Lemma F.7
Proof of Lemma F.7.
Let us define $h(x):=\frac{\mu}{2}\left\|x\right\|^{2}$. Then, we have for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\max_{y^{\prime}\in\mathcal{Y}}\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay^{\prime}+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)$
$=\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)$
$=-\max_{x^{\prime}\in\mathcal{X}}\left(-(x^{\prime})^{\top}Ay^{\mu}-\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)+\max_{x^{\prime}\in\mathcal{X}}\left(-(x^{\prime})^{\top}Ay-\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)$
$=h^{\ast}(-Ay)-h^{\ast}(-Ay^{\mu}).$ (42)
Since $h$ is a $\mu$-smooth function, the convex conjugate $h^{\ast}$ is $\frac{1}{\mu}$-strongly convex [Shalev-Shwartz, 2011, p. 149]. Hence, defining $\mathcal{P}:=\left\{-Ay~{}|~{}y\in\mathcal{Y}\right\}$ and $p^{\ast}:=\mathop{\rm arg~{}min}\limits_{p\in\mathcal{P}}h^{\ast}(p)$, we have for any $p\in\mathcal{P}$:
$h^{\ast}(p)-h^{\ast}(p^{\ast})\geq\frac{1}{2\mu}\left\|p^{\ast}-p\right\|^{2}.$
Furthermore, it holds that $-Ay^{\mu}=\mathop{\rm arg~{}min}\limits_{y\in\mathcal{Y}}h^{\ast}(-Ay^{\mu})=\mathop{\rm arg~{}min}\limits_{p\in\mathcal{P}}h^{\ast}(p)\Leftrightarrow-Ay^{\mu}=p^{\ast}$. Thus, we have:
$h^{\ast}(-Ay)-h^{\ast}(-Ay^{\mu})\geq\frac{1}{2\mu}\left\|A(y^{\mu}-y)\right\|^{2}.$ (43)
By combining (42) and (43), we finally obtain for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\max_{y^{\prime}\in\mathcal{Y}}\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay^{\prime}+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\left\|x^{\prime}\right\|^{2}\right)\geq\frac{1}{2\mu}\left\|A(y^{\mu}-y)\right\|^{2}.$
Appendix G Proof of Theorem B.1
Proof of Theorem B.1.
First, we prove that $y^{\mu}\neq y^{\ast}$ under the assumption that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$ by contradiction. We assume that $y^{\mu}=y^{\ast}$. From (13), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}$ $=x^{\top}Ay^{\ast}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}=v^{\ast}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}.$ (44)
On the other hand,
$\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\ast}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=v^{\ast}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$ (45)
By combining (44) and (45), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}\geq\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$
Hence, from the property of the player $x$’s equilibrium strategy in the perturbed game, $x^{\mu}$ must satisfy $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$.
On the other hand, from the property of the player $y$’s equilibrium strategy $y^{\mu}$ in the perturbed game, $y^{\mu}$ is an optimal solution of the following optimization problem:
$\max_{y\in\Delta^{n}}\left\{(x^{\mu})^{\top}Ay-\frac{\mu_{y}}{2}\left\|y\right\|^{2}\right\}.$
Let us define the following Lagrangian function $L(y,\kappa,\lambda)$ as:
$L(y,\kappa,\lambda)=(x^{\mu})^{\top}Ay-\frac{\mu_{y}}{2}\left\|y\right\|^{2}-\sum_{i=1}^{n}\kappa_{i}g_{i}(y)-\lambda h(y),$
where $g_{i}(y)=-y_{i}$ and $h(y)=\sum_{i=1}^{n}y_{i}-1$. Then, from the KKT conditions, we get the stationarity:
$A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\sum_{i=1}^{n}\kappa_{i}\nabla g_{i}(y^{\mu})-\lambda\nabla h(y^{\mu})=\mathbf{0}_{n},$ (46)
and the complementary slackness:
$\forall i\in[n],~{}\kappa_{i}g_{i}(y^{\mu})=0.$ (47)
Since $y^{\mu}=y^{\ast}$ and $y^{\ast}$ is in the interior of $\Delta^{n}$, we have $g(y^{\mu})=-y_{i}^{\mu}<0$ for all $i\in[n]$. Thus, from (47), we have $\kappa_{i}=0$ for all $i\in[n]$. Substituting this into (46), we obtain:
$A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\lambda\nabla h(y^{\mu})=A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\lambda\mathbf{1}_{n}=\mathbf{0}_{n}.$ (48)
Hence, we have:
$\lambda=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}.$ (49)
Putting (49) into (48) yields:
$y^{\mu}=\frac{1}{\mu_{y}}\left(A^{\top}x^{\mu}-\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}\mathbf{1}_{n}\right)=\frac{1}{\mu_{y}}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-\frac{\mathbf{1}_{n}^{\top}A^{\top}\frac{\mathbf{1}_{m}}{m}-\mu_{y}}{n}\mathbf{1}_{n}\right),$
where the second equality follows from $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$. Multiplying this by $\frac{\mathbf{1}_{m}^{\top}}{m}A$, we have:
$v^{\ast}=\frac{1}{\mu_{y}}\left(\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu_{y}\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}\right),$ (50)
where we used the assumption that $y^{\mu}=y^{\ast}$ and (13). Here, we have:
$\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu_{y}\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}$
$=\left\|v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=n(v^{\ast})^{2}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}+2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$\phantom{=}-n(v^{\ast})^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}-2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\mu_{y}v^{\ast}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu_{y}v^{\ast}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu_{y}v^{\ast}+\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right).$ (51)
Here, since $A^{\top}\frac{\mathbf{1}_{m}}{m}=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}\mathbf{1}_{n}+\mu_{y}y^{\mu}=\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}\mathbf{1}_{n}+\mu_{y}y^{\ast}$ from (48) and (49), we get:
$\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)$
$\phantom{=}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)$
$=\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)+\frac{\mu_{y}^{2}}{n}+\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$\phantom{=}+\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\mathbf{1}_{n}+\mu_{y}^{2}(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$=\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)+\mu_{y}^{2}\left\|y^{\ast}\right\|^{2}$
$=\mu_{y}\left(\mu_{y}\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}\right).$ (52)
By combining (50), (51), and (52), we have:
$\mu_{y}\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}=0.$
Therefore, if $y^{\mu}=y^{\ast}$, then $\mu_{y}$ must satisfy:
$\mu_{y}=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)},$
and this is equivalent to:
$\mu_{y}\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}\Rightarrow y^{\mu}\neq y^{\ast}.$
By a similar argument, in terms of player $x$, we can conclude that:
$\mu_{x}\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left(\left\|x^{\ast}\right\|^{2}-\frac{1}{m}\right)}\Rightarrow x^{\mu}\neq x^{\ast}.$
Appendix H Proofs for Theorem C.1
H.1 Proof of Theorem C.1
Proof of Theorem C.1.
Let us define the total number of anchoring strategy updates as $K:=\frac{T}{T_{\sigma}}$. For $k\in[K]$ and the $k$-th anchoring strategy $\sigma^{k}$, we also define the stationary point $x^{\mu,k}$, which satisfies the following condition:
$x^{\mu,k}=\mathop{\rm arg~{}min}\limits_{x\in\mathcal{X}}\left\{g(x)+\frac{\mu}{2}\left\|x-\sigma^{k}\right\|^{2}\right\},$ (53)
From the first-order optimality condition for $x^{\mu,k}$, we can derive the following lemma for $k$ such that $x^{\mu,k}$:
Lemma H.1.
For any $k\in[K]$ such that $x^{\mu,k}\notin\mathcal{X}^{\ast}$, we have for any $x^{\ast}\in\mathcal{X}^{\ast}$:
$\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{4\mu^{2}}+\frac{D^{2}\mu^{2}}{\alpha^{2}}\left\|x^{\mu,k}-\sigma^{k+1}\right\|^{2}.$
Here, we derive the following upper bound on $\left\|x^{\mu,k}-\sigma^{k+1}\right\|^{2}$:
Lemma H.2.
For an arbitrary perturbation strength $\mu>0$, if we use the learning rate $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta_{k}}{\mu}\right)}\right)$, then $\sigma^{k+1}$ satisfies for any $k\in[K]$:
$\left\|x^{\mu,k}-\sigma^{k+1}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta_{k}^{2}}{\mu^{2}}\right)T_{\sigma}},$
where $\beta_{k}>0$ is a positive constant depending only on $\mathcal{Y}^{\mu,k}:=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{Y}}(x^{\mu,k})^{\top}Ay$.
Note that since $\beta_{k}\leq\|A\|^{2}$ for any $k\in[K]$, Lemma H.2 holds under the assumption that $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\|A\|^{2}}{\mu}\right)}\right)$. By combining Lemmas H.1 and H.2, we have for any $x^{\ast}\in\mathcal{X}^{\ast}$ and $k\in[K]$ such that $x^{\mu,k}\notin\mathcal{X}^{\ast}$:
$\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{4\mu^{2}}+\frac{D^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma_{k}}{T_{\sigma}},$
where $\gamma_{k}:=\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta_{k}^{2}}{\mu^{2}}\right)}$. Setting $x^{\ast}=\prod_{\mathcal{X}^{\ast}}(\sigma^{k})$ in this inequality, we get:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{k+1})-\sigma^{k+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{k})-\sigma^{k+1}\right\|^{2}$
$\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{k})-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{4\mu^{2}}+\frac{D^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma_{k}}{T_{\sigma}}.$ (54)
We now consider two cases depending on whether $x^{\mu,k}\notin\mathcal{X}^{\ast}$ for all $k\in[K]$.
Case 1:
In the first case, we suppose that $x^{\mu,k}\notin\mathcal{X}^{\ast}$ for all $k\in[K]$. In this case, we can iteratively apply (54). Thus, we obtain:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{1})-\sigma^{1}\right\|^{2}-\frac{K\alpha^{2}}{4\mu^{2}}+\frac{KD^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma}{T_{\sigma}},$
where $\gamma:=\max_{k\in[K]}\gamma_{k}$. Since $\sigma^{1}=x^{1}$, we have:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{1})-x^{1}\right\|^{2}-\frac{K\alpha^{2}}{4\mu^{2}}+\frac{KD^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma}{T_{\sigma}},$
Thus, if $K\geq\frac{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{1})-x^{1}\right\|^{2}}{\alpha^{2}}$,
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\frac{D^{2}\mu^{2}\gamma K}{\alpha^{2}T_{\sigma}}.$
Case 2:
Next, we consider the case where there exists $k\in[K]$ such that $x^{\mu,k}\in\mathcal{X}^{\ast}$. Defining $k_{\max}:=\max_{k\in[K]}\left\{k~{}|~{}x^{\mu,k}\in\mathcal{X}^{\ast}\right\}$, we have $x^{\mu,k}\notin\mathcal{X}^{\ast}$ for any $k\in\{k_{\max}+1,\cdots,K\}$. Hence, we can apply (54) from $k=K,K-1,\cdots,k_{\max}+1$.
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{k_{\max}+1})-\sigma^{k_{\max}+1}\right\|^{2}-\frac{(K-k_{\max})\alpha^{2}}{4\mu^{2}}+\frac{(K-k_{\max})D^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma}{T_{\sigma}}.$
Since $\prod_{\mathcal{X}^{\ast}}(x^{\mu,k_{\max}})=x^{\mu,k_{\max}}$ in this case, it holds that:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{k_{\max}+1})-\sigma^{k_{\max}+1}\right\|^{2}\leq\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{\mu,k_{\max}})-\sigma^{k_{\max}+1}\right\|^{2}=\frac{1}{2}\left\|x^{\mu,k_{\max}}-\sigma^{k_{\max}+1}\right\|^{2},$
Combining these two inequalities, we have:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|x^{\mu,k_{\max}}-\sigma^{k_{\max}+1}\right\|^{2}-\frac{(K-k_{\max})\alpha^{2}}{4\mu^{2}}+\frac{(K-k_{\max})D^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma}{T_{\sigma}}.$
By applying Lemma H.2 to this inequality, we have:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\frac{\gamma}{T_{\sigma}}-\frac{(K-k_{\max})\alpha^{2}}{4\mu^{2}}+\frac{(K-k_{\max})D^{2}\mu^{2}}{\alpha^{2}}\frac{\gamma}{T_{\sigma}}$
$\leq\frac{\gamma}{T_{\sigma}}+\frac{D^{2}\mu^{2}\gamma K}{\alpha^{2}T_{\sigma}}=\gamma\left(1+\frac{D^{2}\mu^{2}}{\alpha^{2}}\right)\frac{K}{T_{\sigma}}.$
Therefore, in either case, if $K\geq\frac{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{1})-\sigma^{1}\right\|^{2}}{\alpha^{2}}$, then we have:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\gamma\left(1+\frac{D^{2}\mu^{2}}{\alpha^{2}}\right)\frac{K}{T_{\sigma}}=\gamma\left(1+\frac{D^{2}\mu^{2}}{\alpha^{2}}\right)\frac{T}{T_{\sigma}^{2}}.$ (55)
Setting $T_{\sigma}=cT$ for a constant $c\leq\min\left(1,\frac{\alpha^{2}}{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{1})-\sigma^{1}\right\|^{2}}\right)$, it holds that $K=\frac{T}{T_{\sigma}}=\frac{1}{c}\geq\frac{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{1})-\sigma^{1}\right\|^{2}}{\alpha^{2}}$. Therefore, the condition for (55) is satisfied, and we obtain:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(\sigma^{K+1})-\sigma^{K+1}\right\|^{2}$ $\leq\gamma\left(1+\frac{D^{2}\mu^{2}}{\alpha^{2}}\right)\frac{1}{cT}.$
Since $\sigma^{K+1}=x^{T_{\sigma}K+1}=x^{T+1}$, we have finally:
$\frac{1}{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{T+1})-x^{T+1}\right\|^{2}$ $\leq\gamma\left(1+\frac{D^{2}\mu^{2}}{\alpha^{2}}\right)\frac{1}{cT}.$
H.2 Proof of Lemma H.1
Proof of Lemma H.1.
Since $g$ is a convex function on $\mathcal{X}$ [Boyd and Vandenberghe, 2004], we can utilize the first-order optimality condition for $x^{\mu,k}$. Specifically, there exists a subgradient $h^{k+1}\in\partial g(x^{\mu,k})$ such that:
$\forall x\in\mathcal{X},~{}\langle h^{k+1}+\mu(x^{\mu,k}-\sigma^{k}),x-x^{\mu,k}\rangle\geq 0.$
From the definition of the subgradient, we have for any $x\in\mathcal{X}$:
$\langle h^{k+1},x^{\mu,k}-x\rangle\geq g(x^{\mu,k})-g(x^{\ast}).$
Combining these inequalities, we get for any $x\in\mathcal{X}$:
$g(x^{\mu,k})-g(x)\leq\mu\langle x^{\mu,k}-\sigma^{k},x-x^{\mu,k}\rangle.$
By taking $x=x^{\ast}$ for an arbitrary $x^{\ast}\in\mathcal{X}^{\ast}$, we obtain:
$g(x^{\mu,k})-v^{\ast}=g(x^{\mu,k})-g(x^{\ast})\leq\mu\langle x^{\mu,k}-\sigma^{k},x-x^{\mu,k}\rangle.$ (56)
On the other hand, from (23), we have for any $k\geq 1$:
$\frac{1}{2}\left\|x^{\ast}-x^{\mu,k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}+\frac{1}{2}\left\|x^{\mu,k}-\sigma^{k}\right\|^{2}=\left\langle x^{\mu,k}-\sigma^{k},x^{\mu,k}-x^{\ast}\right\rangle.$ (57)
By combining (56) and (57), we get:
$\frac{1}{2}\left\|x^{\ast}-x^{\mu,k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}+\frac{1}{2}\left\|x^{\mu,k}-\sigma^{k}\right\|^{2}\leq\frac{v^{\ast}-g(x^{\mu,k})}{\mu}\leq 0.$
Thus, from Cauchy-Schwarz inequality and Young’s inequality, we have for any $\rho>0$:
$\frac{1}{2}\left\|x^{\mu,k}-\sigma^{k}\right\|^{2}$
$\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-x^{\mu,k}\right\|^{2}$
$\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}+\sigma^{k+1}-x^{\mu,k}\right\|^{2}$
$\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}-\frac{1}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}-\left\langle x^{\ast}-\sigma^{k+1},\sigma^{k+1}-x^{\mu,k}\right\rangle$
$\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}-\frac{1}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}+\frac{\rho}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}+\frac{1}{2\rho}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}.$
Hence,
$\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|x^{\mu,k}-\sigma^{k}\right\|^{2}-\frac{1}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}$
$\phantom{=}+\frac{\rho}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}+\frac{1}{2\rho}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}.$ (58)
Here, we derive the lower bound on the distance between $\sigma^{k}$ and $x^{\mu,k}$:
Lemma H.3.
For any $k\in[K]$ such that $x^{\mu,k}\notin\mathcal{X}^{\ast}$, we have:
$\frac{\alpha}{\mu}\leq\left\|x^{\mu,k}-\sigma^{k}\right\|.$
By combining (58) and Lemma H.3, we have for any $k\in[K]$ such that $x^{\mu,k}\notin\mathcal{X}^{\ast}$:
$\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{2\mu^{2}}+\frac{\rho}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}+\frac{1}{2\rho}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}$
$\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{2\mu^{2}}+\frac{\rho}{2}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}+\frac{D^{2}}{2\rho}.$
Taking $\rho=\frac{2D^{2}\mu^{2}}{\alpha^{2}}$, we finally have for any $x^{\ast}\in\mathcal{X}^{\ast}$ and $k\in[K]$ such that $x^{\mu,k}\notin\mathcal{X}^{\ast}$:
$\frac{1}{2}\left\|x^{\ast}-\sigma^{k+1}\right\|^{2}$ $\leq\frac{1}{2}\left\|x^{\ast}-\sigma^{k}\right\|^{2}-\frac{\alpha^{2}}{4\mu^{2}}+\frac{D^{2}\mu^{2}}{\alpha^{2}}\left\|\sigma^{k+1}-x^{\mu,k}\right\|^{2}.$
H.3 Proof of Lemma H.2
Proof of Lemma H.2.
By applying the same proof technique as in Theorem 4.1, under the assumption that $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\beta_{k}}{\mu}\right)}\right)$, we have for any $k\in[K]$ and $t\in\{(k-1)T_{\sigma}+1,\cdots,kT_{\sigma}\}$:
$\left\|x^{\mu,k}-x^{t+1}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta_{k}^{2}}{\mu^{2}}\right)(t-(k-1)T_{\sigma}+1)},$
where $\beta_{k}>0$ depends only on the set $\mathcal{Y}^{\mu,k}:=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{X}}(x^{\mu,k})^{\top}Ay$.
Taking $t=kT_{\sigma}$, we have $x^{t+1}=x^{kT_{\sigma}+1}=\sigma^{k+1}$ from the definition of $\sigma^{k+1}$. Thus, we have:
$\left\|x^{\mu,k}-\sigma^{k+1}\right\|^{2}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta_{k}^{2}}{\mu^{2}}\right)(T_{\sigma}+1)}\leq\frac{256D^{2}(\mu+\|A\|)^{2}}{\eta^{2}\min\left(\mu^{2},\frac{\beta_{k}^{2}}{\mu^{2}}\right)T_{\sigma}}.$
H.4 Proof of Lemma H.3
Proof of Lemma H.3.
By taking $x^{\ast}=\prod_{\mathcal{X}^{\ast}}(x^{\mu,k})$ in (56), we have:
$g(x^{\mu,k})-v^{\ast}\leq\mu\left\langle x^{\mu,k}-\sigma^{k},\prod_{\mathcal{X}^{\ast}}(x^{\mu,k})-x^{\mu,k}\right\rangle.$
By combining this inequality and Lemma E.1, we get:
$\alpha\left\|x^{\mu,k}-\prod_{\mathcal{X}^{\ast}}(x^{\mu,k})\right\|$ $\leq\mu\left\langle x^{\mu,k}-\sigma^{k},\prod_{\mathcal{X}^{\ast}}(x^{\mu,k})-x^{\mu,k}\right\rangle$
$\leq\mu\left\|x^{\mu,k}-\sigma^{k}\right\|\left\|\prod_{\mathcal{X}^{\ast}}(x^{\mu,k})-x^{\mu,k}\right\|.$
Therefore, as long as $x^{\mu,k}\notin\mathcal{X}^{\ast}$, it holds that:
$\frac{\alpha}{\mu}\leq\left\|x^{\mu,k}-\sigma^{k}\right\|.$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the main claims and contributions, including important assumptions and limitations of the proposed approach.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: This paper discusses the limitations of the proposed algorithm in Remark 4.3.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical results are presented with complete proofs, clearly stated assumptions, and proper cross-references in both the main paper and appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed algorithmic descriptions and experimental settings in Sections 4 and 5, Appendix A, and Appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The pseudocode used for the experiments is presented in Algorithm 1. We also submit the source code as supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 5, Appendix A, and Appendix C fully specify the game settings and the hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Figures 6 and 8, we report error bars over $100$ random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The computer resources used for the experiments are detailed in Appendix A.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research fully complies with the NeurIPS Code of Ethics, with all aspects of the work conducted in accordance with established ethical guidelines and best practices.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper presents work whose goal is to solve saddle-point optimization problems. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper presents theoretical and algorithmic contributions without releasing any models or datasets that could pose risks for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cited OpenSpiel [Lanctot et al., 2019], which is used in the experiments on extensive-form games.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The source code for our algorithms is provided as supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This research does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This research does not use LLMs as part of its core methodology, algorithms, or experimental validation.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.