### v2

# Asymmetric Perturbation in Solving Bilinear Saddle-Point Optimization

Abstract
This paper proposes an asymmetric perturbation technique for solving bilinear saddle-point optimization problems, commonly arising in minimax problems, game theory, and constrained optimization. Perturbing payoffs or values is known to be effective in stabilizing learning dynamics and equilibrium computation. However, it requires decreasing perturbation magnitudes to ensure convergence to an equilibrium in the underlying game, resulting in a slower rate. To overcome this, we introduce an asymmetric perturbation approach, where only one player’s payoff function is perturbed. Exploiting the near-linear structure of bilinear problems, we show that, for a sufficiently small perturbation, the equilibrium strategy of the asymmetrically perturbed game coincides with an equilibrium strategy of the original game. Building on this property, we develop a perturbation-based learning algorithm with a linear last-iterate convergence rate to an equilibrium strategy of the original game, and we further show how to construct a parameter-free procedure that retains a linear rate. Finally, we empirically demonstrate fast convergence toward equilibria in both normal-form and extensive-form games.
1 Introduction
This paper proposes an asymmetric perturbation technique for solving saddle-point optimization problems, commonly arising in minimax problems, game theory, and constrained optimization. Over the past decade, no-regret learning algorithms have been extensively studied for computing (approximate) solutions or equilibria. When each player minimizes regret, the time-averaged strategies approximate Nash equilibria in two-player zero-sum games; that is, average-iterate convergence is guaranteed. However, the actual sequence of strategies does not necessarily converge and can cycle or diverge even in simple bilinear cases (Mertikopoulos et al., 2018; Bailey and Piliouras, 2018; Cheung and Piliouras, 2019). This is problematic, especially in large-scale games with neural network policies, since averaging requires storing a separate model at every iteration.
This motivates the study of last-iterate convergence, a stronger notion than average-iterate convergence, in which the strategies themselves converge to an equilibrium. One successful approach is to use optimistic learning algorithms, which essentially incorporate a one-step optimistic prediction that the environment will behave similarly in the next step. This idea has led to several effective algorithms, including Extra-Gradient methods (EG) (Liang and Stokes, 2019; Mokhtari et al., 2020), Optimistic Gradient Descent Ascent (OGDA) (Daskalakis and Panageas, 2019; Gidel et al., 2019; Mertikopoulos and Zhou, 2019), and Optimistic Multiplicative Weights Update (OMWU) (Daskalakis and Panageas, 2019; Lei et al., 2021a). However, in large-scale settings where the gradient must be estimated from data or simulation, these algorithms can lose the last-iterate convergence property. For example, Abe et al. (2022) reports empirical non-convergence behavior under bandit feedback.
Alternatively, perturbing the payoffs with strongly convex penalties (Facchinei and Pang, 2003) has long been recognized as an effective technique for achieving last iterate convergence (Koshal et al., 2010; Tatarenko and Kamgarpour, 2019). This line of work has also shown strong performance in practical settings, including learning in large-scale games (Bakhtin et al., 2023) and fine-tuning large language models via preference optimization (Ye et al., 2024), often in place of optimistic algorithms. In prior work, the perturbation is almost always applied symmetrically, meaning both players’ payoff functions are perturbed by a strongly convex penalty. A known limitation is that, for any fixed perturbation strength $\mu>0$, the equilibrium of the symmetrically perturbed game generally remains only an approximation of an equilibrium of the original game, and the deviation scales with the strength of the perturbation (Liu et al., 2023; Abe et al., 2024). Consequently, recovering equilibria of the original game typically requires tuning $\mu$ carefully (e.g., as a function of the iteration budget) or employing a continuation-type procedure, which introduces a nontrivial trade-off between accuracy and convergence speed.
To avoid these restrictions, we develop an asymmetric perturbation approach, in which only one player’s payoff function is perturbed while the other remains unperturbed. This simple modification yields a qualitatively different outcome. For any sufficiently small perturbation strength within a broad and practical range, the equilibrium strategy of the perturbed game coincides with that of the original game (see Theorem 3.1). Intuitively, leaving player $y$ unperturbed preserves the linearity of player $x$’s original objective, so adding a strongly convex perturbation does not significantly shift the solution (see Figure 2). Consequently, solving the asymmetrically perturbed game suffices to recover an equilibrium strategy of the original game.
Furthermore, to demonstrate the effectiveness of our findings, we first incorporate the asymmetric perturbation technique into a gradient-based learning algorithm for bilinear games, and show that it converges to a saddle point at a linear (exponentially fast) rate (see Theorem 4.1 and Corollary 4.2). This rate is provably faster than an $\tilde{\mathcal{O}}(1/t)$ [^1] We use $\tilde{\mathcal{O}}$ to denote a Landau notation that disregards a polylogarithmic factor. rate for existing symmetric perturbation-based approaches under the same bilinear setting (Liu et al., 2023). Moreover, leveraging the same invariance phenomenon, we show how to construct a parameter-free variant that retains a linear last-iterate convergence rate. Second, we apply our asymmetric perturbation to gradient-based learning methods for extensive-form games using a dilated regularizer (Hoda et al., 2010), a standard choice for learning over sequence-form strategy spaces (von Stengel, 1996). While our analysis focuses on bilinear games, the structural insight behind the asymmetric perturbation may extend beyond this setting, including two-player zero-sum Markov games, and serves as a bridge to the design of new perturbation-based learning algorithms.
2 Preliminaries
Bilinear saddle-point optimization problems.
In this study, we focus on the following bilinear saddle-point problem:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay,$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ (resp. $\mathcal{Y}\subseteq\mathbb{R}^{n}$) represents the $m$-dimensional (resp. $n$-dimensional) convex strategy space for player $x$ (resp. player $y$), and $A\in\mathbb{R}^{m\times n}$ is a game matrix. We assume that $\mathcal{X}$ and $\mathcal{Y}$ are polytopes. We refer to the function $x^{\top}Ay$ as the payoff function, and write $z:=(x,y)\in\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ as the strategy profile. This formulation includes many well-studied classes of games, such as two-player normal-form games and extensive-form games with perfect recall.
Nash equilibrium.
This study aims to compute a minimax or maximin strategy in the optimization problem (1). Let $\mathcal{X}^{\ast}:=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}x^{\top}Ay$ denote the set of minimax strategies, and let $\mathcal{Y}^{\ast}:=\mathop{\rm arg~max}\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}x^{\top}Ay$ denote the set of maximin strategies. It is well-known that any strategy profile $(x^{\ast},y^{\ast})\in\mathcal{X}^{\ast}\times\mathcal{Y}^{\ast}$ is a Nash equilibrium, which satisfies the following condition:
$\forall(x,y)\in\mathcal{X}\times\mathcal{Y},~(x^{\ast})^{\top}Ay\leq(x^{\ast})^{\top}Ay^{\ast}\leq x^{\top}Ay^{\ast}.$
Based on the minimax theorem (v. Neumann, 1928), every equilibrium $(x^{\ast},y^{\ast})\in\mathcal{X}^{\ast}\times\mathcal{Y}^{\ast}$ attains the identical value, denoted as $v^{\ast}$, which can be expressed as:
$v^{\ast}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay=\max_{y\in\mathcal{Y}}\min_{x\in\mathcal{X}}x^{\top}Ay.$
We refer to $v^{\ast}$ as the game value. To quantify the proximity to equilibrium for a given strategy profile $(x,y)$, we use NashConv, which is defined as follows:
$\mathrm{NashConv}(x,y)$ $=\max_{\tilde{y}\in\mathcal{Y}}x^{\top}A\tilde{y}-\min_{\tilde{x}\in\mathcal{X}}(\tilde{x})^{\top}Ay.$
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 1: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix AA is given by [[0,1,3],[1,0,1],[3,1,0]][[0,1,-3],[-1,0,1],[3,-1,0]]. The proximity to the minimax strategy is measured by the value of $\max_{y\in\mathcal{Y}}(x^{\mu})^{\top}Ay-v^{\ast}$.
Symmetric perturbation.
Payoff perturbation is an extensively studied technique for solving games (Facchinei and Pang, 2003; Liu et al., 2023). In this approach, the payoff functions of all players are perturbed by a strongly convex function $\psi$. For example, in bilinear games, instead of solving the original game in (1), we solve the following perturbed game:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\mu\psi(x)-\mu\psi(y)\right\},$
where $\mu\in(0,\infty)$ is the perturbation strength. Since the perturbation is applied to both players’ payoff functions, we refer to this perturbed game as a symmetrically perturbed game. In this study, we specifically focus on the standard case, where the perturbation payoff function $\psi$ is given by the squared $\ell^{2}$-norm, i.e., $\psi(x)=\frac{1}{2}\|x\|^{2}$:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$ (2)
Let $x^{\mu}$ (resp. $y^{\mu}$) denote the minimax (resp. maximin) strategy in the symmetrically perturbed game (2)  [^2] The minimax strategy is uniquely determined because symmetrically perturbed games satisfy the strongly convex–strongly concave property., which can be solved at a linear rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). It is known that the solution $(x^{\mu},y^{\mu})$ is only an approximation of an equilibrium of the original game, with an error bounded by $\mathcal{O}(\mu)$ (Liu et al., 2023; Abe et al., 2024). Consequently, typical perturbation-based methods must employ a decreasing schedule for $\mu$, or use an extremely small fixed $\mu$ tuned to the number of iterations $T$ (e.g., $\mu=\mathcal{O}(1/T)$), which requires careful hyperparameter tuning (Tatarenko and Kamgarpour, 2019; Bernasconi et al., 2022; Cai et al., 2023). See Figure 1 for a biased Rock–Paper–Scissors game where the perturbed solution differs from the original equilibrium. We provide a rigorous justification for this phenomenon in Appendix B, showing that for this instance, $(x^{\mu},y^{\mu})$ fails to coincide with any equilibrium of the original game for every fixed $\mu>0$.
3 Asymmetric Payoff Perturbation
In this section, we explain our novel technique of asymmetric payoff perturbation. We demonstrate that a seemingly minor structural change—perturbing only one player’s payoff—can yield a dramatically different outcome: the solution of the perturbed game exactly matches an equilibrium strategy $x^{\ast}$ in many cases.
3.1 Asymmetric Payoff Perturbation
Instead of incorporating the perturbation into both players’ payoff functions, we consider the case where only player $x$’s payoff function is perturbed:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}\right\}.$ (3)
The procedure we are going to describe in Theorem 3.1 focuses on computing the minimax strategy $x^{\ast}$, rather than the maximin strategy $y^{\ast}$. To compute $y^{\ast}$, we simply solve the corresponding maximin problem for player $y$:
$\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$
The same reasoning applies to this perturbed maximin optimization problem. Thus, hereafter, we primarily focus on the perturbed game (3) from the perspective of player $x$.
Figure 2: The landscape of the objective function for player $x$ in asymmetrically perturbed games. The functions $g(x)$ and $g^{\mu}_{\mathrm{asym}}(x)$ are defined as $g(x):=\max\limits_{y\in\mathcal{Y}}x^{\top}Ay$ and $g^{\mu}_{\mathrm{asym}}(x):=g(x)+\frac{\mu}{2}\|x\|^{2}$, respectively.
Since the function $\max_{y\in\mathcal{Y}}x^{\top}Ay$ is convex with respect to $x$ (Boyd and Vandenberghe, 2004), the perturbed objective $\max_{y\in\mathcal{Y}}x^{\top}Ay+\frac{\mu}{2}\|x\|^{2}$ is $\mu$-strongly convex. Therefore, the minimax strategy for the perturbed game (3) is unique. We denote it by $x^{\mu}$ and denote the set of maximin strategies in (3) by $\mathcal{Y}^{\mu}$. Since both the minimax and maximin strategies constitute a Nash equilibrium of the perturbed game, the pair $(x^{\mu},y^{\mu})$ with $y^{\mu}\in\mathcal{Y}^{\mu}$ satisfies the following conditions: for all $\tilde{y}^{\mu}\in\mathcal{Y}^{\mu}$ and $x\in\mathcal{X}$,
$(x^{\mu})^{\top}A\tilde{y}^{\mu}+\frac{\mu}{2}\left\|x^{\mu}\right\|^{2}\leq x^{\top}A\tilde{y}^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2},$ (4)
and for all $y\in\mathcal{Y}$,
$(x^{\mu})^{\top}Ay^{\mu}\geq(x^{\mu})^{\top}Ay.$ (5)
3.2 Equilibrium Invariance under the Asymmetric Perturbation
In this section, we discuss the properties of the minimax strategies for asymmetrically perturbed games. Surprisingly, we can show that $x^{\mu}$ in (3) does correspond to a minimax strategy in the original game (1) for all $\mu$ smaller than a certain positive constant:
Theorem 3.1.
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\left\|x\right\|})$, where $\alpha>0$ is a constant depending only the game instance. Then, the minimax strategy $x^{\mu}$ in the corresponding asymmetrically perturbed game (3) satisfies $x^{\mu}\in\mathcal{X}^{\ast}$.
Thus, whenever $\mu$ is below a certain positive constant, $x^{\mu}$ coincides with the minimax strategy of the original game. Figure 1 illustrates this feature in a simple example (in that example, invariance holds for $\mu<2.5$). The asymmetric perturbation is closely related in spirit to Nesterov’s smoothing technique (Nesterov, 2005), which also perturbs only one side of a minimax formulation by adding a strongly convex term. A key distinction is which payoff is perturbed: smoothing is typically used by perturbing the opponent’s payoff on the maximization side, whereas we perturb the payoff of the player being optimized and leave the opponent’s payoff unchanged. This reversal is essential for Theorem 3.1, which identifies a regime of fixed $\mu$ in which the minimax strategy is invariant under the perturbation.
Remark 3.2 (Invariance without knowing the game-dependent constant).
The invariance result above is guaranteed when $\mu$ lies below a game-dependent upper bound, as characterized by $\alpha$. Nevertheless, this does not prevent us from exploiting invariance in practice for two reasons. First, we empirically observe that the invariance phenomenon appears to persist for a reasonably wide range of choices of $\mu$ in our experiments (see Figures 1 and 4). Second, in Section 4.2, we show how Theorem 3.1 can be leveraged to construct a parameter-free procedure that does not require knowledge of $\alpha$.
The key ingredient in proving Theorem 3.1 is the near-linear behavior of the objective function $g(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay$ for player $x$ in the original game. Specifically, according to Wei et al. (2021), there exists a constant $\alpha>0$ such that:
$\forall x^{\ast}\in\mathcal{X}^{\ast},~g(x)-g\left(x^{\ast}\right)\geq\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|,$
where $\Pi_{\mathcal{A}}(a)=\mathop{\rm arg~min}\limits_{a^{\prime}\in\mathcal{A}}\left\|a-a^{\prime}\right\|$ represents the projection operator onto a given closed convex set $\mathcal{A}$. This inequality implies that deviating from the minimax strategy set $\mathcal{X}^{\ast}$ results in an increase in the objective function proportionally to the distance $\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|$. In contrast, the variation (i.e., the gradient) of the perturbation payoff function $\frac{\mu}{2}\left\|x\right\|^{2}$ can always be bounded by $\mathcal{O}(\mu)$ over $\mathcal{X}$. Hence, by choosing $\mu$ sufficiently small, we can ensure that the perturbation payoff function does not significantly incentivize player $x$ to deviate from $x^{\ast}$. In Figure 2, we illustrate this fact intuitively. The addition of the strongly convex function $\frac{\mu}{2}\|x\|^{2}$ does not shift the optimum $x^{*}$ of the original $g(x)$ if $\mu$ is sufficiently small, as the kink of the lines through $g(x)$ dominates. The detailed proof of Theorem 3.1 is provided in Appendix E.
In summary, to compute a minimax strategy $x^{\ast}\in\mathcal{X}^{\ast}$ in the original game, it is sufficient to solve the asymmetrically perturbed game (3) with a small perturbation strength $\mu>0$. Note that, as mentioned above, one can also compute a maximin strategy $y^{\ast}\in\mathcal{Y}^{\ast}$ by solving the game $\max\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}$, where the payoff perturbation is applied only to player $y$.
Figure 3: Trajectories of strategies for player $x$ using AsymP-GDA, SymP-GDA, and GDA. The game matrix AA is set to A=[[0,1,3],[1,0,1],[3,1,0]]A=[[0,1,-3],[-1,0,1],[3,-1,0]], and the strategy spaces are set to $\mathcal{X}=\mathcal{Y}=\Delta^{3}$. The red point represents the minimax strategy in the original game. The trajectories originate from different initial strategies, demonstrating the learning dynamics under each method.
4 Asymmetrically Perturbed Gradient Descent Ascent
This section proposes a first-order method, Asymmetrically Perturbed Gradient Descent Ascent (AsymP-GDA), for solving asymmetrically perturbed games (3). At each iteration $t\in[T]$, AsymP-GDA updates each player’s strategy according to the following alternating updates [^3] AsymP-GDA employs alternating updates rather than simultaneous updates, as recent work has demonstrated the advantages of the former over the latter (Lee et al., 2024).:
$\begin{aligned} x^{t+1}&=\Pi_{\mathcal{X}}\left(x^{t}-\eta\left(Ay^{t}+\mu x^{t}\right)\right),\\ y^{t+1}&=\Pi_{\mathcal{Y}}\left(y^{t}+\eta A^{\top}x^{t+1}\right),\end{aligned}$ (6)
where $\eta>0$ is the learning rate. In AsymP-GDA, player $x$’s strategy $x^{t}$ is updated based on the gradient of the perturbed payoff function $x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}$, while player $y$’s strategy $y^{t}$ is updated using the gradient of the original payoff function $x^{\top}Ay$. AsymP-GDA adds only negligible per-iteration runtime or memory overhead relative to standard alternating GDA, with the only additional operation being a single vector addition.
Since the perturbed payoff function of player $x$ is strongly convex, it is anticipated that AsymP-GDA enjoys a last-iterate convergence guarantee. By combining this observation with Theorem 3.1, when $\mu$ is sufficiently small, the updated strategy $x^{t}$ should converge to a minimax strategy $x^{\ast}$ in the original game. We confirm this empirically by plotting the trajectory of $x^{t}$ updated by AsymP-GDA in a sample normal-form game, as shown in Figure 3. We also provide the trajectories of GDA and SymP-GDA; in the latter, the squared $\ell^{2}$-norm perturbs both players’ gradients. For both AsymP-GDA and SymP-GDA, the perturbation strength is set to $\mu=1$. As expected, AsymP-GDA successfully converges to the minimax strategy (red point) in the original game, whereas SymP-GDA converges to a point far from the minimax strategy, and GDA cycles around the minimax strategy. Further details and additional experiments in normal-form games can be found in Appendix A.3.
4.1 Last-Iterate Convergence Rate
In this section, we provide a last-iterate convergence result for AsymP-GDA. Let $\|A\|$ denote the largest singular value of the matrix $A$. As discussed in the previous section, when both players’ objectives are perturbed by strongly convex penalties, standard first-order methods converge to an approximate equilibrium at a linear rate (Cen et al., 2021, 2023; Pattathil et al., 2023; Sokota et al., 2023). Theorem 4.1 shows that, perhaps surprisingly, a linear convergence rate can still be achieved even when the perturbation is applied only to one player:
Theorem 4.1.
For an arbitrary perturbation strength $\mu>0$, if the learning rate satisfies $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$, then $x^{t}$ and $y^{t}$ satisfy:
$\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{t})-z^{t}\right\|^{2}\leq\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{1})-z^{1}\right\|^{2}\cdot\exp(-\Theta(t)),$
where $Z^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$.
Note that Theorem 4.1 holds for any fixed $\mu>0$. Hence, combining Theorems 3.1 and 4.1, we can conclude that if $\mu$ is sufficiently small (which does not need to depend on the number of iterations $t$ or $T$), player $x$’s strategy $x^{t}$ updated by AsymP-GDA converges to a minimax strategy in the original game (1):
Corollary 4.2.
Assume that the perturbation strength $\mu$ is set such that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\left\|x\right\|})$, and the learning rate is set so that $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$. Then, AsymP-GDA ensures the convergence of $x^{t}$ to an equilibrium $x^{\ast}$ in the original game at a rate of $\exp(-\mathcal{O}(t))$:
$\left\|x^{\ast}-x^{t}\right\|^{2}\leq\left\|\Pi_{\mathcal{Z^{\mu}}}(z^{1})-z^{1}\right\|^{2}\cdot\exp(-\Theta(t)).$
Corollary 4.2 provides a linear last-iterate convergence rate for AsymP-GDA. This rate is competitive with those of optimistic methods, such as OGDA and OMWU in certain settings (e.g., bilinear games) (Wei et al., 2021). Empirically, however, AsymP-GDA exhibits faster convergence in our experiments (see Figure 6 in Appendix A).
4.2 Parameter-free AsymP-GDA
Although Corollary 4.2 guarantees convergence to an equilibrium of the original game only when the perturbation strength $\mu$ is sufficiently small, the appropriate scale of $\mu$ may be unknown a priori. To eliminate this tuning requirement, we also introduce a parameter-free variant (Algorithm 1 in Appendix D) that solves asymmetrically perturbed games along a sequence of perturbation strengths $\mu\in(0,\mu_{\mathrm{init}}]$, starting from an arbitrarily large $\mu_{\mathrm{init}}>0$. For each perturbed game, the algorithm runs AsymP-GDA until the duality gap falls below a prescribed threshold, and then checks whether the resulting strategy profile achieves $\mathrm{NashConv}\leq\varepsilon$. If not, it halves $\mu$ and moves on to the next perturbed game.
Thanks to Theorem 3.1, it suffices to solve only finitely many perturbed games, regardless of any target accuracy $\varepsilon$. Since each perturbed game can be solved at a linear rate by AsymP-GDA, the overall procedure retains a linear rate to an equilibrium of the original game:
Theorem 4.3.
For any target accuracy $\varepsilon>0$ and any initialization $\mu_{\mathrm{init}}>0$, Algorithm 1 returns a strategy profile $(x,y)$ with $\mathrm{NashConv}(x,y)\leq\varepsilon$ after at most $\mathcal{O}(\ln(1/\varepsilon))$ iterations in total.
Whereas Liu et al. (2023) adopt a similar decreasing-$\mu$ mechanism, their analysis yields a $\tilde{\mathcal{O}}(1/t)$ rate. In contrast, our analysis in the above theorem shows a faster linear rate. This gap stems from the fact that our decreasing-$\mu$ procedure requires solving only a bounded number of perturbed games, while in Liu et al. (2023) the number of perturbed games grows as one targets higher accuracy.
Figure 4: Performance in extensive-form games. AsymP-DGDA performs two strategy updates per iteration for each player. For a fair comparison across methods with different per-iteration computational costs, we report the total number of strategy updates on the x-axis rather than iterations.
4.3 Proof Sketch of Theorem 4.1
This section outlines the proof sketch for Theorem 4.1. The complete proofs are provided in Appendix F.
(1) Monotonic decrease of the distance function.
Firstly, leveraging the strong convexity of the perturbation payoff function, $\frac{\mu}{2}\left\|x\right\|^{2}$, we can show that the distance between the current strategy profile $z^{t}=(x^{t},y^{t})$ and any equilibrium $z^{\mu}=(x^{\mu},y^{\mu})$ monotonically decreases under $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$. Specifically, we have for any $t\geq 1$:
$\left\|z^{\mu}-z^{t+1}\right\|^{2}-\left\|z^{\mu}-z^{t}\right\|^{2}$
$\!\leq\!-\!\left(\eta\mu\!-\!\eta^{2}\!\left(\mu^{2}\!+\!\|A\|^{2}\right)\right)\!\left\|x^{\mu}-x^{t+1}\right\|^{2}\!-\frac{1}{2}\left\|z^{t+1}\!-z^{t}\right\|^{2}\!.$ (7)
(2) Lower bound on the path length.
The primary technical challenge is deriving the term related to the distance between $y^{t+1}$ and the maximin strategies set $\mathcal{Y}^{\mu}$, i.e., $\left\|\Pi_{\mathcal{Y}^{\mu}}(y^{t+1})-y^{t+1}\right\|^{2}$, which leads to the last-iterate convergence rate. To this end, we derive the following lower bound on the path length $\left\|z^{t+1}-z^{t}\right\|$ by the distance $\|x^{\mu}-x^{t}\|^{2}$ and $\|y^{\mu}-y^{t}\|^{2}$ (as shown in Lemma E.1).
Let us represent the update rule of AsymP-GDA in (6) as $z^{t+1}=T_{\mu}(z^{t})$ using the operator $T_{\mu}:\mathcal{Z}\to\mathcal{Z}$. We show that $T_{\mu}$ is a piecewise affine mapping on $\mathcal{Z}$ (in Lemma F.3). Hence, there exists finitely many polyhedral sets such that $T_{\mu}$ can be represented as an affine function within each polyhedral set. For each set and the corresponding affine function, we can apply Hoffman’s bound (Hoffman, 1952), resulting in the following metric subregularity:
$\forall z\in\mathcal{Z},~\left\|T_{\mu}(z)-z\right\|\geq\Omega\left(\left\|z-\Pi_{Z^{\mu}}(z)\right\|\right),$
where $\mathcal{Z}^{\mu}:=\{x^{\mu}\}\times\mathcal{Y}^{\mu}$. By setting $z=z^{t}$ in the above inequality, we obtain the following lower bound on $\left\|z^{t+1}-z^{t}\right\|$ by the distance between the current strategy profile and the equilibrium set:
$\left\|z^{t+1}-z^{t}\right\|\geq\Omega\left(\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|\right).$ (8)
(3) Last-iterate convergence rate.
Putting (8) into (7), we have for any $t\geq 1$:
$\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{2}$
$\leq\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}-\Omega\left(\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\right),$
Therefore, by mathematical induction, we finally obtain the following upper bound on the distance $\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}$:
$\left\|\Pi_{\mathcal{Z}^{\mu}}(z^{t})-z^{t}\right\|^{2}\leq\exp\left(-\Theta(t)\right).$
Remark 4.4 (Technical challenge in proving Theorem 4.1).
The main technical challenge in proving Theorem 4.1 arises from the asymmetric nature of the perturbation, which is applied only to player $x$. Unlike the symmetric case, where strong convexity in both players’ payoff functions directly yields contraction, the asymmetric setting requires a more subtle analysis since player $y$’s payoff function remains linear. Our proof establishes a linear last-iterate convergence rate by lower-bounding the variation of $z^{t}$ by the distance from $z^{t}$ to the equilibrium set $\mathcal{Z}^{\mu}$, as shown in (8), rather than relying on strong convexity. A further complication is the presence of projections onto $\mathcal{X}$ and $\mathcal{Y}$: the update mapping $T_{\mu}$ is piecewise affine, meaning that the affine mapping governing the update can switch across iterations. As a result, the dynamics cannot be analyzed as a single linear system, in contrast to unconstrained linear-quadratic games (Zhang et al., 2022). See Appendix F for details.
5 Dilated AsymP-GDA for Extensive-Form Games
We now extend our approach to extensive-form games, which model sequential decision-making under imperfect information. To cast these games as saddle-point problems, we adopt the sequence-form representation (von Stengel, 1996). In the sequence-form representation, each player’s strategy is parameterized by realization (reach) probabilities over feasible action sequences, rather than by enumerating pure strategies (i.e., complete behavior plans). Moreover, to reduce the per-iteration computational cost of AsymP-GDA in the sequence-form representation, we use the dilated Euclidean regularizer (Hoda et al., 2010) in the following experiments.
Sequence-form representation.
We consider two-player zero-sum extensive-form games with perfect recall. Let $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$ denote the sets of information sets of players $x$ and $y$, respectively. For each information set $i\in\mathcal{I}_{x}$, let $A(i)$ be the set of actions available at $i$.
We use the sequence-form representation to express each player’s strategy. For player $x$, we write:
$x=(x_{i,a})_{i\in\mathcal{I}_{x},a\in A(i)},$
where $x_{i,a}\geq 0$ denotes the reach probability of taking action $a$ at information set $i$, counting only the contribution of player $x$’s own action probabilities. To express the constraints for the strategy space $\mathcal{X}$, for each information set $i\in\mathcal{I}_{x}$, we define $\mathrm{parent}(i)$ to be the unique parent (predecessor) pair [^4] The uniqueness of the predecessor is guaranteed by the perfect recall assumption. $(j,b)$ (an information set $j$ and an action $b$) on player $x$’s own decision sequence that immediately precedes $i$. We also write $\mathrm{parent}(i)=\varnothing$ if no such predecessor exists, namely if $i$ is a root information set for player $x$ along its own decision sequence. The strategy space $\mathcal{X}$ is then given by the following constraints:
$\forall i\in\mathcal{I}_{x},~\sum_{a\in A(i)}x_{i,a}=x_{\mathrm{parent}(i)}\mbox{ and }x_{i,a}\geq 0,$
where $x_{\mathrm{parent}(i)}$ denotes $x_{j,b}$ if there exists $(j,b)$ such that $(j,b)=\mathrm{parent}(i)$, and denotes $1$ if $\mathrm{parent}(i)=\varnothing$. We define $\mathcal{Y}$ analogously for player $y$.
In the sequence-form representation, the two-player zero-sum extensive-form game admits a bilinear saddle-point formulation: there exists a matrix $A$ such that $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}x^{\top}Ay$.
Dilated AsymP-GDA.
Due to the computational efficiency of the strategy update, we propose a variant of AsymP-GDA that employs the dilated squared Euclidean regularizer (instead of the standard squared Euclidean regularizer) for both the proximal regularizer and the perturbation term. We refer to this variant as AsymP-DGDA. For comparison, we also evaluate a symmetrically perturbed counterpart, SymP-DGDA.
For player $x$, we define the dilated squared Euclidean regularizer (Hoda et al., 2010):
$\psi_{\mathrm{dil}}(x)=\frac{1}{2}\sum_{i\in\mathcal{I}_{x}}\alpha_{i}\sum_{a\in A(i)}\frac{x_{i,a}^{2}}{x_{\mathrm{parent}(i)}},$
where $\alpha_{i}>0$ is a weight parameter, and similarly for player $y$. Using $\psi_{\mathrm{dil}}$, AsymP-DGDA updates each player’s strategy as:
$x^{t+1}$ $=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\left\{\eta\langle x,Ay^{t}+\mu\nabla\psi_{\mathrm{dil}}(x^{t})\rangle+D_{\psi_{\mathrm{dil}}}(x,x^{t})\right\},$
$y^{t+1}$ $=\mathop{\rm arg~min}\limits_{y\in\mathcal{Y}}\left\{\eta\langle y,-A^{\top}x^{t+1}\rangle+D_{\psi_{\mathrm{dil}}}(y,y^{t})\right\},$
where $D_{\psi_{\mathrm{dil}}}(a,b)=\psi_{\mathrm{dil}}(a)-\psi_{\mathrm{dil}}(b)-\langle\nabla\psi_{\mathrm{dil}}(b),a-b\rangle$. In contrast, SymP-DGDA applies perturbation symmetrically to both players. As with AsymP-GDA, AsymP-DGDA adds only negligible per-iteration runtime or memory overhead relative to standard Dilated GDA.
Empirical performance.
We compare the NashConv of the last-iterate strategy for AsymP-DGDA against SymP-DGDA and baseline algorithms, including Dilated MWU (DMWU), Dilated GDA (DGDA) [^5] DMWU corresponds to mirror descent on the sequence-form strategy space with the dilated entropy regularizer $\psi_{\mathrm{dil\mbox{-}ent}}(x)=\sum_{i\in\mathcal{I}_{x}}\alpha_{i}\sum_{a\in A(i)}x_{i,a}\ln\frac{x_{i,a}}{x_{\mathrm{parent}(i)}}$. DGDA corresponds to mirror descent with the dilated Euclidean regularizer., Dilated OMWU (Kroer et al., 2020; Lee et al., 2021), and Dilated OGDA (Farina et al., 2019; Lee et al., 2021). Our experiments focus on five different extensive-form games: Kuhn Poker, Leduc Poker, Liar’s Dice (with four-sided), and Goofspiel (with four-card and five-card variants), all of which are implemented using LiteEFG (Liu et al., 2024). The detailed hyperparameters of the algorithms, tuned for best performance, are shown in Table 4 in Appendix A.
Figure 4 shows the NashConv values for each game. As indicated by these results, AsymP-DGDA not only converges with a competitive or faster speed than any other method in all games but also directly reaches an equilibrium strategy, whereas SymP-DGDA converges near the equilibrium. These results confirm that the asymmetric perturbation leads to convergence in extensive-form games.
6 Related Literature
Saddle-point optimization problems have attracted significant attention due to their applications in machine learning, such as training generative adversarial networks (Daskalakis et al., 2018). No-regret learning algorithms have been extensively studied with the aim of achieving either average-iterate or last-iterate convergence. To attain last-iterate convergence, many recent algorithms incorporate optimism (Rakhlin and Sridharan, 2013a, b), including optimistic multiplicative weights update (Daskalakis and Panageas, 2019; Lei et al., 2021b; Wei et al., 2021), optimistic gradient descent ascent (Daskalakis et al., 2018; Mertikopoulos et al., 2019; de Montbrun and Renault, 2022), and extra-gradient methods (Golowich et al., 2020; Mokhtari et al., 2020).
As an alternative approach, payoff perturbation has gained renewed attention. In this approach, players’ payoff functions are regularized with strongly convex terms (Cen et al., 2021, 2023; Pattathil et al., 2023), which stabilizes the dynamics and leads to convergence. Some existing works have shown convergence to an approximate equilibrium under fixed perturbation (Sokota et al., 2023; Tuyls et al., 2006; Coucheney et al., 2015; Leslie and Collins, 2005; Abe et al., 2022; Hussain et al., 2023). To recover equilibria of the original game, later studies have employed a decreasing schedule or iterative regularization (Facchinei and Pang, 2003; Koshal et al., 2013; Yousefian et al., 2017; Bernasconi et al., 2022; Liu et al., 2023; Cai et al., 2023), or have updated the regularization center periodically (Perolat et al., 2021; Abe et al., 2023, 2024). In contrast to these approaches, our algorithms avoid iteration-budget-dependent tuning of the perturbation strength by exploiting the equilibrium invariance property of our asymmetric perturbation.
Extensive-form games, which model sequential decision-making under imperfect information, have been studied extensively from both theoretical and empirical perspectives. Regarding last-iterate convergence, Lee et al. (2021) establish last-iterate convergence guarantees for optimistic algorithms in the sequence-form representation (von Stengel, 1996). More recently, Liu et al. (2023) show last-iterate convergence for symmetrically perturbed algorithms with a decreasing schedule on the perturbation strength. In contrast, our asymmetric perturbation avoids such delicate scheduling of the perturbation strength, and thus may be more amenable to large-scale extensive-form settings where optimistic algorithms can be difficult to deploy reliably.
7 Conclusion
This paper introduces an asymmetric perturbation technique for solving saddle-point optimization problems, addressing key challenges in learning dynamics and equilibrium computation. Unlike symmetric perturbation, which in general yields only approximate equilibria for any fixed perturbation strength, our approach admits an invariance regime: for sufficiently small $\mu$, the minimax strategy of the perturbed game coincides with that of the original game. Building on this structural property, we propose AsymP-GDA and establish a linear last-iterate convergence rate, improving over symmetric perturbation approaches in the same bilinear setting. Moreover, we show that the same invariance phenomenon can be leveraged to design a parameter-free procedure that retains a linear rate without requiring knowledge of game-dependent constants.
Our theoretical results target bilinear two-player zero-sum games. The key insight of equilibrium invariance in Theorem 3.1 relies on the near-linear growth of the objective function. We believe that an analogous formulation can be posed beyond bilinear games, including two-player zero-sum Markov games. Moreover, extending asymmetric perturbation to behavioral-form representation in extensive-form games and analysis of its performance are promising future directions.
Appendix A Experimental Details and Additional Experimental Results
A.1 Information on the Computer Resources
All experiments in this paper were conducted on macOS Sonoma 14.4.1 with Apple M2 Max and 32GB RAM.
A.2 Proximity to Equilibrium under Symmetric and Asymmetric Perturbations in Biased Matching Pennies
This section investigates the proximity of $x^{\mu}$ to the equilibrium $x^{\ast}$ in the Biased Matching Pennies (BMP) game under the symmetric/asymmetric payoff perturbation, with varying perturbation strength $\mu$. The game matrix for BMP is provided in Table 1.
Table 1: Game matrix in BMP
$y_{1}$ $y_{2}$
$x_{1}$ $1/3$ 2/3-2/3
$x_{2}$ 2/3-2/3 11
BMP has a unique equilibrium $x^{\ast}=y^{\ast}=(\frac{5}{8},\frac{3}{8})$, and the game value is given as $v^{\ast}=-\frac{1}{24}$.
Figure 5 exhibits the proximity of $x^{\mu}$ to $x^{\ast}$ as $\mu$ varies. Notably, under the symmetric perturbation, $x^{\mu}$ coincides with $x^{\ast}$ when $\mu$ is set to $\mu=\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}=\frac{4}{3}$. This result underscores the statement in Theorem B.1, that $x^{\mu}$ does not coincide with $x^{\ast}$ as long as $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$.
(a) Symmetric Perturbation
(b) Asymmetric Perturbation
Figure 5: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix AA is given by $\frac{1}{3},-\frac{2}{3}],[-\frac{2}{3},1]]$.
A.3 Additional Experiments in Normal-Form Games
In this section, we experimentally compare our AsymP-GDA with SymP-GDA, GDA, and OGDA [Daskalakis et al., 2018, Wei et al., 2021]. We conduct experiments on two normal-form games: Biased Rock-Paper-Scissors (BRPS) and Multiple Nash Equilibria (M-Ne). These games are taken from Abe et al. [2023] and Wei et al. [2021]. Tables 3 and 3 provide the game matrices for BRPS and M-Ne, respectively.
Table 2: Game matrix in BRPS
$y_{1}$ $y_{2}$ $y_{3}$
$x_{1}$ 0 11 3-3
$x_{2}$ 1-1 0 11
$x_{3}$ 33 1-1 0
Table 3: Game matrix in M-Ne
$y_{1}$ $y_{2}$ $y_{3}$ $y_{4}$ $y_{5}$
$x_{1}$ 0 1-1 11 0 0
$x_{2}$ 11 0 1-1 0 0
$x_{3}$ 1-1 11 0 0 0
$x_{4}$ 1-1 11 0 22 1-1
$x_{5}$ 1-1 11 0 1-1 22
Figure 6 illustrates the logarithm of NashConv averaged over $100$ different random seeds. For each random seed, the initial strategies $(x^{0},y^{0})$ are chosen uniformly at random within the strategy spaces $\mathcal{X}=\Delta^{m}$ and $\mathcal{Y}=\Delta^{n}$. We use a learning rate of $\eta=0.01$ for each algorithm, and a perturbation strength of $\mu=1$ for both AsymP-GDA and SymP-GDA. We observe that AsymP-GDA converges to the minimax strategy in the original game, while SymP-GDA converges to a point far from the minimax strategy.
(a) BRPS
(b) M-Ne
Figure 6: Performance of AsymP-GDA, SymP-GDA, GDA, and OGDA in normal-form games. The shaded area represents the standard errors.
Figure 7: Trajectories for SymP-GDA (top row) and AsymP-GDA (bottom row) under different perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS.
Figure 8: Sensitivity to the perturbation strength for AsymP-GDA with $\eta=0.01$ in BRPS and M-Ne.
Figure 7 illustrates the trajectories of SymP-GDA (top row) and AsymP-GDA (bottom row) under varying perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS. For SymP-GDA, the trajectories do not converge directly to the equilibrium even for small values of $\mu=0.5,1.0$. Instead, they follow circuitous and elongated paths, resulting in slower convergence. Conversely, as $\mu$ increases ($\mu=2.0,4.0$), the trajectories become more direct, leading to faster convergence, but they remain farther from the equilibrium. In contrast, AsymP-GDA leads to direct convergence to the equilibrium with small perturbation strengths. For $\mu$ values up to $2.0$ the trajectories converge directly to the equilibrium. However, as $\mu$ increases beyond a threshold ($\mu=4.0$), the trajectory deviates from the equilibrium. These results provide a more detailed understanding of the trends observed in Figures 1 and 1, further illustrating the differences in convergence dynamics between symmetric and asymmetric perturbations.
Figure 8 illustrates the performance of AsymP-GDA on BRPS and M-Ne with $\mu\in\{0.01,0.1,1.0,2.0,4.0\}$ and $\eta=0.01$. For sufficiently small $\mu$, the limit point coincides with an equilibrium of the original game. However, decreasing $\mu$ also slows convergence. Overall, these results highlight a trade-off between accuracy and convergence speed.
A.4 Hyperparameter Settings for Extensive-Form Games
For the experiments in Section 5, we tuned the hyperparameters of each algorithm separately for each game to obtain the best performance. The resulting settings are summarized in Table 4. For the dilated regularizer, we use the unweighted version throughout, setting $\alpha_{i}=1$ for all $i$, as in Lee et al. [2021].
A.5 Comparison with CFR-Based Algorithms
This section compares the performance of AsymP-DGDA with several CFR-based algorithms, including CFR [Zinkevich et al., 2007], CFR+ [Tammelin, 2014], Discounted CFR (DCFR), and Linear CFR (LCFR) [Brown and Sandholm, 2019]. We use the same games as in Section 5 and keep the hyperparameter settings for AsymP-DGDA and SymP-DGDA unchanged. Figure 9 shows the logarithm of NashConv as a function of the number of strategy updates. For AsymP-DGDA and SymP-DGDA, we report NashConv for the last-iterate strategies. For the CFR-based algorithms, we report NashConv for the average-iterate strategies. Overall, AsymP-DGDA achieves lower NashConv than the CFR-based baselines on most games, with Leduc Poker being the exception.
Game Algorithm $\eta$ $\mu$
Kuhn Poker DMWU 0.01 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.1 0.001
AsymP-DGDA 0.1 0.01
Leduc Poker DMWU 0.05 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.05 0.0001
AsymP-DGDA 0.1 0.0001
Liars Dice (4 sides) DMWU 0.01 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.01 0.0001
AsymP-DGDA 0.1 0.001
Goofspiel (4 cards) DMWU 0.1 -
DGDA 0.1 -
DOMWU 0.01 -
DOGDA 0.01 -
SymP-DGDA 0.1 0.0001
AsymP-DGDA 0.1 0.05
Goofspiel (5 cards) DMWU 0.1 -
DGDA 0.1 -
DOMWU 0.01 -
DOGDA 0.01 -
SymP-DGDA 0.1 0.0001
AsymP-DGDA 0.1 0.001
Table 4: Hyperparameters
Figure 9: Performance in extensive-form games. AsymP-DGDA performs two strategy updates per iteration for each player. For a fair comparison across methods with different per-iteration computational costs, we report the total number of strategy updates on the x-axis rather than iterations.
Appendix B Impossibility Results for Symmetric Perturbation
As we stated in Section 2, existing works [Liu et al., 2023, Abe et al., 2024] have shown that the distance between the solution of the symmetrically perturbed game $(x^{\mu},y^{\mu})$ and the solution in the original game $(x^{\ast},y^{\ast})$ is upper bounded by $\mathcal{O}(\mu)$. However, they do not guarantee that the two solutions coincide, even for a small $\mu>0$. In contrast, the following theorem provides the first formal impossibility result, proving that $(x^{\mu},y^{\mu})$ almost never coincides with $(x^{\ast},y^{\ast})$.
Theorem B.1.
Consider a normal-form game with a unique interior equilibrium. Assume that at this equilibrium, neither player chooses their actions uniformly at random, i.e., $(x^{\ast},y^{\ast})\neq\left(\frac{1}{m}\mathbf{1}_{m},\frac{1}{n}\mathbf{1}_{n}\right)$. Then, for any $\mu>0$ such that $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$, the minimax strategy $x^{\mu}$ in (2) satisfies $x^{\mu}\neq x^{\ast}$. Furthermore, for any $\mu>0$ such that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the maximin strategy $y^{\mu}$ in (2) satisfies $y^{\mu}\neq y^{\ast}$.
The proof is provided in Appendix H. Additionally, we extend our analysis to the case where both players have different perturbation strengths, i.e., $\mu_{x}>0$ and $\mu_{y}>0$, as shown in Appendix C.
Discussion on Theorem B.1.
The term $\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$ can be interpreted as a measure of the difference between the equilibrium $(x^{\ast},y^{\ast})$ and the uniform random strategy profile $(\frac{1}{m}\mathbf{1}_{m},\frac{1}{n}\mathbf{1}_{n})$. Specifically, the numerator $\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}$ represents the difference in the payoffs, while the denominator $\left\|x^{\ast}\right\|^{2}-\frac{1}{m}$ represents the difference in the squared $\ell^{2}$-norms, respectively. A promising direction for future research is to theoretically demonstrate that, when $\mu=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the corresponding equilibrium coincides exactly with the equilibrium in the original game, i.e., $(x^{\mu},y^{\mu})=(x^{*},y^{*})$. We have experimentally confirmed this, and the results are presented in the Appendix A.2.
When the game is symmetric.
Next, let us consider the case when $A^{\top}=-A$, as in Rock-Paper-Scissors. In this scenario, the equilibrium strategies $x^{\mu}$ and $y^{\mu}$ are not identical to the minimax or maximin strategies of the original game, regardless of the choice of $\mu>0$.
Corollary B.2.
Assume that $A^{\top}=-A$. Under the same setup as Theorem B.1, the equilibrium $(x^{\mu},y^{\mu})$ in (2) always satisfies $x^{\mu}\neq x^{\ast}$ and $y^{\mu}\neq y^{\ast}$ for any $\mu>0$.
This is because it always holds that $v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)=0$ when $A^{\top}=-A$. Figure 1 shows the proximity of $x^{\mu}$ to $x^{\ast}$ with varying perturbation strength $\mu$ in a simple biased Rock-Paper-Scissors game. We observe that as long as $\mu>0$, $x^{\mu}$ remains distant from $x^{\ast}$. This observation supports the theoretical results in Theorem B.1 and Corollary B.2.
Appendix C Independently Perturbed Game
Let us consider the perturbed game where players $x$ and $y$ choose independently their perturbation strengths $\mu_{x}$ and $\mu_{y}$:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu_{x}}{2}\left\|x\right\|^{2}-\frac{\mu_{y}}{2}\left\|y\right\|^{2}\right\}.$ (9)
We establish a theoretical result similar to Theorem 3.1 for this perturbed game.
Theorem C.1.
Assume that the original game is a normal-form game with a unique interior equilibrium, and that $(x^{\ast},y^{\ast})\neq\left(\frac{\mathbf{1}_{m}}{m},\frac{\mathbf{1}_{n}}{n}\right)$, i.e., the equilibrium is not the uniform random strategy profile. Then, for any $\mu_{x}>0$ such that $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$, the minimax strategy $x^{\mu}$ in the corresponding symmetrically perturbed game (9) satisfies $x^{\mu}\neq x^{\ast}$. Furthermore, for any $\mu>0$ such that $\mu_{y}\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$, the maximin strategy $y^{\mu}$ in (9) satisfies $y^{\mu}\neq y^{\ast}$.
[…]
2 $(\hat{x}_{x}^{\mu_{0}},\hat{y}_{x}^{\mu_{0}})\leftarrow(x_{\mathrm{init}},y_{\mathrm{init}}),~(\hat{x}_{y}^{\mu_{0}},\hat{y}_{y}^{\mu_{0}})\leftarrow(x_{\mathrm{init}},y_{\mathrm{init}})$
3 for $k=1,2,\cdots$ do
4 $\mu_{k}\leftarrow\frac{\mu_{\mathrm{init}}}{2^{k-1}}$
5 $\eta_{k}\leftarrow\min\left(\eta_{k-1},\frac{\mu}{\mu^{2}+mn}\right)$
6 $\delta_{k}\leftarrow\frac{\mu_{k}}{2\|A\|^{2}\max_{z\in\mathcal{Z}}\|z\|^{2}}\varepsilon^{2}$
7 $(x_{x}^{1},y_{x}^{1})\leftarrow(\hat{x}_{x}^{\mu_{k-1}},\hat{y}_{x}^{\mu_{k-1}})$
8 $(x_{y}^{1},y_{y}^{1})\leftarrow(\hat{x}_{y}^{\mu_{k-1}},\hat{y}_{y}^{\mu_{k-1}})$
9 $(\hat{x}_{x}^{\mu_{k}},\hat{y}_{x}^{\mu_{k}})\leftarrow\textsc{AsymP-GDA}_{x}(x_{x}^{1},y_{x}^{1},\eta_{k},\mu_{k},\delta_{k})$
10 $(\hat{x}_{y}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}})\leftarrow\textsc{AsymP-GDA}_{y}(x_{y}^{1},y_{y}^{1},\eta_{k},\mu_{k},\delta_{k})$
11 if $\max_{x\in\mathcal{X}}\langle A\hat{y}_{y}^{\mu_{k}},\hat{x}_{x}^{\mu_{k}}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}\hat{x}_{x}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}}-y\rangle\leq\varepsilon$ then
12    return $\hat{x}_{x}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}}$
13   end if
14
15 end for
-3mm
1mm
16 subroutine $\textsc{AsymP-GDA}_{x}(x^{1},y^{1},\eta,\mu,\delta)$
17 for $t=1,2,\cdots$ do
18    $x^{t+1}=\Pi_{\mathcal{X}}\left(x^{t}-\eta(Ay^{t}+\mu x^{t})\right)$
19    $y^{t+1}=\Pi_{\mathcal{Y}}\left(y^{t}+\eta A^{\top}x^{t+1}\right),$~
20    if $\max_{x\in\mathcal{X}}\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}x^{t+1},y^{t+1}-y\rangle\leq\delta$ then
21       return $(x^{t+1},y^{t+1})$
22
23
24   end for
25
26
-3mm
1mm
27 subroutine $\textsc{AsymP-GDA}_{y}(x^{1},y^{1},\eta,\mu,\delta)$
28 for $t=1,2,\cdots$ do
29    $y^{t+1}=\Pi_{\mathcal{Y}}\left(y^{t}+\eta(A^{\top}x^{t}-\mu y^{t})\right),$~
30    $x^{t+1}=\Pi_{\mathcal{X}}\left(x^{t}-\eta Ay^{t+1}\right)$
31    if $\max_{x\in\mathcal{X}}\langle Ay^{t+1},x^{t+1}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}x^{t+1}+\mu y^{t+1},y^{t+1}-y\rangle\leq\delta$ then
32       return $(x^{t+1},y^{t+1})$
33
34
35   end for
36
37
Algorithm 1 Adaptive-$\mu$ AsymP-GDA
Appendix D Parameter-Free AsymP-GDA
As noted in Remark 3.2, AsymP-GDA may require a very small perturbation strength $\mu$ to guarantee convergence to an equilibrium of the original game. To address this limitation, this section introduces a parameter-free variant (Algorithm 1) that adaptively shrinks $\mu$.
Algorithm 1 takes a target NashConv value $\varepsilon$ as an input, and runs AsymP-GDA over multiple episodes. In each episode, the algorithm runs AsymP-GDA until the duality gap of the current strategies falls below $\Theta(\varepsilon^{2})$. It then checks whether the resulting strategy profile has a NashConv value at most $\varepsilon$. If the condition is satisfied, the algorithm terminates and outputs the strategies. Otherwise, it halves the perturbation strength $\mu$ and proceeds to the next episode.
D.1 Last-Iterate Convergence Rate
We establish the last-iterate convergence rate for Algorithm 1. Without loss of generality, we assume $\max_{x\in\mathcal{X}}\|A^{\top}x\|_{\infty}\leq 1$, $\max_{y\in\mathcal{Y}}\|Ay\|_{\infty}\leq 1$, and $\max_{(i,j)\in[m]\times[n]}|A_{ij}|\leq 1$ since these conditions can be satisfied by re-scaling $A$. Under this assumption, we show that the iteration complexity for the algorithm with an arbitrarily large initial perturbation strength is $\mathcal{O}(\ln(1/\varepsilon))$, in Theorem 4.3,
Theorem 4.3 implies that Algorithm 1 attains a linear last-iterate convergence rate, i.e., $\mathcal{O}(\exp(-t))$, even when the perturbation strength $\mu$ is initialized arbitrarily large. While Liu et al. [2023] have adopted a similar shrinking mechanism, their analysis yields a $\tilde{\mathcal{O}}(1/t)$ rate, which is slower than the linear rate established here.
Appendix E Proof of Theorem 3.1
Proof of Theorem 3.1.
Let us define the function $g_{\mathrm{asym}}^{\mu}:\mathcal{X}\to\mathbb{R}$:
$g_{\mathrm{asym}}^{\mu}(x):=\max_{y\in\mathcal{Y}}x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}.$
Here, we introduce the following property of the function $\max_{y\in\mathcal{Y}}x^{\top}Ay$:
Lemma E.1 (Claim 1-5 in Theorem 5 of Wei et al. [2021]).
There exists a positive constant $\alpha>0$ such that:
$\forall x\in\mathcal{X},$~ $\max_{y\in\mathcal{Y}}x^{\top}Ay-v^{\ast}\geq\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|,$
$\forall y\in\mathcal{Y},$~ $v^{\ast}-\min_{x\in\mathcal{X}}x^{\top}Ay\geq\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|,$
where $\alpha$ depends only on $\mathcal{X}^{\ast}$ and $\mathcal{Y}^{\ast}$.
By using Lemma E.1, we have for any $x\in\mathcal{X}$:
$\max_{y\in\mathcal{Y}}\left(\Pi_{\mathcal{X}^{\ast}}(x)\right)^{\top}Ay-\max_{y\in\mathcal{Y}}x^{\top}Ay=v^{\ast}-\max_{y\in\mathcal{Y}}x^{\top}Ay\leq-\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|.$
On the other hand, for any $x^{\ast}\in\mathcal{X}^{\ast}$ and $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$, we have:
$\frac{\mu}{2}\left\|x^{\ast}\right\|^{2}-\frac{\mu}{2}\left\|x\right\|^{2}$ $=\frac{\mu}{2}\langle x^{\ast}-x,x^{\ast}+x\rangle$
$=\frac{\mu}{2}\langle x^{\ast}-x,-x^{\ast}+x+2x^{\ast}\rangle$
$\leq\mu\langle x^{\ast}-x,x^{\ast}\rangle$
$\leq\mu\left\|x^{\ast}-x\right\|\left\|x^{\ast}\right\|.$
Summing up these inequalities, we obtain:
$g_{\mathrm{asym}}^{\mu}\left(\Pi_{\mathcal{X}^{\ast}}(x)\right)-g_{\mathrm{asym}}^{\mu}(x)$ $\leq-\alpha\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|+\mu\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|\left\|\Pi_{\mathcal{X}^{\ast}}(x)\right\|$
$=-\left(\alpha-\mu\left\|\Pi_{\mathcal{X}^{\ast}}(x)\right\|\right)\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|$
$\leq-\left(\alpha-\mu\max_{x\in\mathcal{X}}\left\|x\right\|\right)\left\|x-\Pi_{\mathcal{X}^{\ast}}(x)\right\|.$
Hence, under the assumption that $\mu<\frac{\alpha}{\max_{x\in\mathcal{X}}\left\|x\right\|}$, we have for any $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$:
$g_{\mathrm{asym}}^{\mu}\left(\Pi_{\mathcal{X}^{\ast}}(x)\right)<g_{\mathrm{asym}}^{\mu}(x).$
Thus, every $x\in\mathcal{X}\setminus\mathcal{X}^{\ast}$ is dominated by some equilibrium strategy $x^{\ast}\in\mathcal{X}^{\ast}$ with respect to the value of $g_{\mathrm{asym}}^{\mu}$. Therefore, we conclude that:
$x^{\mu}\in\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}g_{\mathrm{asym}}^{\mu}(x)\Leftrightarrow x^{\mu}\in\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\left\|x\right\|^{2}\right\}\Leftrightarrow x^{\mu}\in\mathcal{X}^{\ast}.$
Appendix F Proofs for Theorem 4.1
F.1 Proof of Theorem 4.1
Proof of Theorem 4.1.
First, we have for any vectors $a,b,c$:
$\frac{1}{2}\left\|a-b\right\|^{2}-\frac{1}{2}\left\|a-c\right\|^{2}+\frac{1}{2}\left\|b-c\right\|^{2}=\langle c-b,a-b\rangle.$ (10)
From (10), we have for any $t\geq 1$:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}=\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle$ (11)
Here, from the first-order optimality condition for $x^{t+1}$ in (6), we have for any $t\geq 1$:
$\left\langle\eta Ay^{t}+\eta\mu x^{t}+x^{t+1}-x^{t},x^{t+1}-x^{\mu}\right\rangle\leq 0.$ (12)
Combining (11) and (12) yields:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}$
$\leq$ $\eta\left\langle Ay^{t}+\mu x^{t},x^{\mu}-x^{t+1}\right\rangle$
$=$ $\eta\left\langle Ay^{t+1}+\mu x^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1}+\mu(x^{t}-x^{t+1}),x^{\mu}-x^{t+1}\right\rangle$
$=$ $\eta\left\langle Ay^{\mu}+\mu x^{\mu},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1}+\mu(x^{t}-x^{t+1}),x^{\mu}-x^{t+1}\right\rangle$
$+\eta\left\langle Ay^{t+1}-Ay^{\mu},x^{\mu}-x^{t+1}\right\rangle-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}.$ (13)
On the other hand, from the first-order optimality condition for $x^{\mu}$ in (4), we get:
$\left\langle Ay^{\mu}+\mu x^{\mu},x^{\mu}-x^{t+1}\right\rangle\leq 0.$ (14)
By combining (13) and (14), we have for any $t\geq 1$:
$\frac{1}{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}-\frac{1}{2}\left\|x^{\mu}-x^{t}\right\|^{2}+\frac{1}{2}\left\|x^{t+1}-x^{t}\right\|^{2}$
$\leq$ $-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\eta\left\langle Ay^{t}-Ay^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\mu\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle$
$+\eta\left\langle Ay^{t+1}-Ay^{\mu},x^{\mu}-x^{t+1}\right\rangle.$ (15)
Similar to (11), we have for any $t\geq 1$:
$\frac{1}{2}\left\|y^{\mu}-y^{t+1}\right\|^{2}-\frac{1}{2}\left\|y^{\mu}-y^{t}\right\|^{2}+\frac{1}{2}\left\|y^{t+1}-y^{t}\right\|^{2}=\left\langle y^{t}-y^{t+1},y^{\mu}-y^{t+1}\right\rangle,$ (16)
and from the first-order optimality condition for $y^{t+1}$ in (6), we have for any $t\geq 1$:
$\left\langle-\eta A^{\top}x^{t+1}+y^{t+1}-y^{t},y^{t+1}-y^{\mu}\right\rangle\leq 0.$ (17)
By combining (16) and (17), we get:
$\frac{1}{2}\left\|y^{\mu}-y^{t+1}\right\|^{2}-\frac{1}{2}\left\|y^{\mu}-y^{t}\right\|^{2}+\frac{1}{2}\left\|y^{t+1}-y^{t}\right\|^{2}$
$\leq-\eta\left\langle A^{\top}x^{t+1},y^{\mu}-y^{t+1}\right\rangle$
$=-\eta\left\langle A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle-\eta\left\langle A^{\top}x^{t+1}-A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle$
$\leq-\eta\left\langle A^{\top}x^{t+1}-A^{\top}x^{\mu},y^{\mu}-y^{t+1}\right\rangle,$ (18)
where the last inequality stems from (5).
Summing up (15) and (18), we have for any $t\geq 1$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}+\frac{1}{2}\left\|z^{t+1}-z^{t}\right\|^{2}$
$\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\eta\mu\left\langle x^{t}-x^{t+1},x^{\mu}-x^{t+1}\right\rangle+\eta\left\langle Ay^{t}-Ay^{t+1},x^{\mu}-x^{t+1}\right\rangle$
$\leq-\eta\mu\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|x^{t}-x^{t+1}\right\|^{2}+\eta^{2}\mu^{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|y^{t}-y^{t+1}\right\|^{2}+\eta^{2}\|A\|^{2}\left\|x^{\mu}-x^{t+1}\right\|^{2}$
$=-\left(\eta\mu-\eta^{2}(\mu^{2}+\|A\|^{2})\right)\left\|x^{\mu}-x^{t+1}\right\|^{2}+\frac{1}{4}\left\|z^{t}-z^{t+1}\right\|^{2}.$
Hence, under the assumption that $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$, we have for any $y^{\mu}\in\mathcal{Y}^{\mu}$:
$\frac{1}{2}\left\|z^{\mu}-z^{t+1}\right\|^{2}-\frac{1}{2}\left\|z^{\mu}-z^{t}\right\|^{2}\leq-\frac{1}{4}\left\|z^{t}-z^{t+1}\right\|^{2}.$
Defining $z_{t}^{\mu}=\mathop{\rm arg~min}\limits_{z\in Z^{\mu}}\|z-z^{t}\|$, we obtain:
$\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{t+1}\right\|^{2}\leq\left\|z_{t}^{\mu}-z^{t+1}\right\|^{2}$ $\leq\left\|z^{\mu,t}-z^{t}\right\|^{2}-\frac{1}{2}\left\|z^{t}-z^{t+1}\right\|^{2}$
$=\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{t}\right\|^{2}-\frac{1}{2}\left\|z^{t}-z^{t+1}\right\|^{2}.$ (19)
Here, let us define the operator $T_{\mu}:\mathcal{Z}\to\mathcal{Z}$ as follows:
$T_{\mu}(x,y):=\left(\Pi_{\mathcal{X}}\left(x-\eta\left(Ay+\mu x\right)\right),~\Pi_{\mathcal{Y}}\left(y+\eta A^{\top}\Pi_{\mathcal{X}}\left(x-\eta\left(Ay+\mu x\right)\right)\right)\right).$
Then, the update rule of AsymP-GDA in (6) can be written as $z^{t+1}=T_{\mu}(z^{t})$. Regarding $T_{\mu}$, we can show the following property:
Lemma F.1.
There exists a positive constant $\beta>0$ such that: for any $z\in\mathcal{Z}$,
$\min_{z^{\mu}\in Z^{\mu}}\left\|z-z^{\mu}\right\|\leq\beta\|z-T_{\mu}(z)\|.$
Setting $z=z^{t}$ in Lemma F.1, we obtain:
$\|z^{t+1}-z^{t}\|$ $=\|z^{t}-T_{\mu}(z^{t})\|$
$\geq\frac{1}{\beta}\min_{z^{\mu}\in Z^{\mu}}\left\|z^{t}-z^{\mu}\right\|$
$\geq\frac{1}{\beta}\min_{z^{\mu}\in Z^{\mu}}\left\|z^{t+1}-z^{\mu}\right\|,$
where the last inequality follows from (19).
Plugging the above inequality into (19), we have:
$\left(1+\frac{1}{2\beta^{2}}\right)\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{t+1}\right\|^{2}\leq\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{t}\right\|^{2}.$
Therefore, when $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$, we have for any $t\geq 1$:
$\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{t+1}\right\|^{2}\leq\left(\frac{1}{1+1/(2\beta^{2})}\right)^{t}\min_{z^{\mu}\in Z^{\mu}}\left\|z^{\mu}-z^{1}\right\|^{2}.$
F.2 Proof of Lemma F.1
Proof of Lemma F.1.
First, we introduce the definition of piecewise affine mappings from Rockafellar and Wets [2009].
Definition F.2 (Definition 2.47 in Rockafellar and Wets [2009]).
A mapping $F:\mathcal{S}\to\mathbb{R}^{m}$ for a set $\mathcal{S}\subseteq\mathbb{R}^{n}$ is piecewise affine on $\mathcal{S}$ if $\mathcal{S}$ can be represented as the union of finitely many polyhedral sets, relative to each of which $F(x)$ is given by an expression of the form $Ax+a$ for some matrix $A\in\mathbb{R}^{m\times n}$ and vector $a\in\mathbb{R}^{m}$.
Based on the above definition, we show that the piecewise affineness of $T_{\mu}$:
Lemma F.3.
The operator $T_{\mu}$ is piecewise affine on $\mathcal{Z}$.
By Lemma F.3, there exists finitely many polyhedral sets $\{R_{k}\}_{k=1}^{K}$ such that $\bigcup_{k=1}^{K}R_{k}=\mathcal{Z}$ and $T_{\mu}(z)=M_{k}z+q_{k}$ for some matrix $M_{k}$ and vector $q_{k}$ for any $z\in R_{k}$. Since $R_{k}$ is polyhedral, $R_{k}$ can be represented as $R_{k}=\{z~|~G_{k}z\leq h_{k}\}$ for some matrix $G_{k}$ and vector $h_{k}$. Then, we can represent the set of fixed points within $R_{k}$ with respect to $T_{\mu}$ as:
$S_{k}:=\{z\in R_{k}~|~T_{\mu}(z)=z\}=\{z~|~G_{k}z\leq h_{k}\mbox{ and }(I-M_{k})z=q_{k}\}$
From Hoffman’s bound [Hoffman, 1952], for each $S_{k}$ satisfying $S_{k}\neq\emptyset$, there exists a positive constant $c_{k}>0$ such that for any $z$:
$\min_{z^{\prime}\in S_{k}}\left\|z-z^{\prime}\right\|\leq c_{k}\left(\left\|(G_{k}z-h_{k})_{+}\right\|+\left\|(I-M_{k})z-q_{k}\right\|\right),$
where $(a)_{+}$ denotes the elementwise positive part of a given vector $a$, i.e., $((a)_{+})_{i}:=\max(a_{i},0)$ for all $i$. If $z\in R_{k}$, then $\left\|(G_{k}z-h_{k})_{+}\right\|=0$. Hence, we have for any $z\in R_{k}$:
$\min_{z^{\prime}\in S_{k}}\left\|z-z^{\prime}\right\|\leq c_{k}\left\|(I-M_{k})z-q_{k}\right\|=c_{k}\left\|z-T_{\mu}(z)\right\|.$
Thus, defining $\mathrm{Fix}(T_{\mu}):=\{z\in\mathcal{Z}~|~T_{\mu}(z)=z\}$, we get for any $z\in R_{k}$ satisfying $S_{k}\neq\emptyset$:
$\min_{z^{\prime}\in\mathrm{Fix}(T_{\mu})}\left\|z-z^{\prime}\right\|\leq\min_{z^{\prime}\in S_{k}}\left\|z-z^{\prime}\right\|\leq c_{k}\left\|z-T_{\mu}(z)\right\|,$ (20)
since $\mathrm{Fix}(T_{\mu})=\bigcup_{k=1}^{K}S_{k}$.
On the other hand, since $z-T_{\mu}(z)$ is continuous and $R_{k}\cap\mathcal{Z}$ is compact, $\|z-T_{\mu}(z)\|$ has a minimum over $R_{k}\cap\mathcal{Z}$. Thus, if $S_{k}=\emptyset$, we can define the following positive constant:
$\delta_{k}:=\min_{z\in R_{k}\cap\mathcal{Z}}\left\|z-T_{\mu}(z)\right\|>0.$
Hence, letting $D:=\max_{z,z^{\prime}\in\mathcal{Z}}\left\|z-z^{\prime}\right\|$ denote the diameter of $\mathcal{Z}$, we obtain for $z\in R_{k}$ satisfying $S_{k}=\emptyset$:
$\min_{z^{\prime}\in\mathrm{Fix}(T_{\mu})}\left\|z-z^{\prime}\right\|\leq D=\frac{D}{\|z-T_{\mu}(z)\|}\|z-T_{\mu}(z)\|\leq\frac{D}{\delta_{k}}\|z-T_{\mu}(z)\|.$ (21)
By combining (20) and (21), it holds that there exists a positive constant $\beta>0$ such that:
$\forall z\in\mathcal{Z},~\min_{z^{\prime}\in\mathrm{Fix}(T_{\mu})}\left\|z-z^{\prime}\right\|\leq\beta\|z-T_{\mu}(z)\|.$ (22)
Lemma F.4.
For any $\mu$, it holds that $\mathrm{Fix}(T_{\mu})=Z^{\mu}$.
By combining (22) and Lemma F.4, we have:
$\forall z\in\mathcal{Z},~\min_{z^{\prime}\in Z^{\mu}}\left\|z-z^{\prime}\right\|\leq\beta\|z-T_{\mu}(z)\|.$
F.3 Proof of Lemma F.3
Proof of Lemma F.3.
We introduce the following lemma:
Lemma F.5 (Example 12.31 in Rockafellar and Wets [2009]).
For a given nonempty, closed, convex set $\mathcal{S}$, the following properties are equivalent:
  • the set $\mathcal{S}$ is polyhedral,
  • the projection mapping $\Pi_{\mathcal{S}}$ is piecewise affine.
By using Lemma F.5, $\Pi_{\mathcal{X}}$ and $\Pi_{\mathcal{Y}}$ are piecewise affine. Since the composition of piecewise affine mappings is also piecewise affine [Gorokhovik, 2011], the operator $T_{\mu}$ is a piecewise affine map. ∎
F.4 Proof of Lemma F.4
Proof of Lemma F.4.
We first show that $z\in\mathrm{Fix}(T_{\mu})\Rightarrow z\in Z^{\mu}$. If $z=(x,y)\in\mathrm{Fix}(T_{\mu})$, we have:
$\begin{aligned} x&=\Pi_{\mathcal{X}}\left(x-\eta\left(Ay+\mu x\right)\right),\\ y&=\Pi_{\mathcal{Y}}\left(y+\eta A^{\top}\Pi_{\mathcal{X}}\left(x-\eta\left(Ay+\mu x\right)\right)\right)=\Pi_{\mathcal{Y}}\left(y+\eta A^{\top}x\right).\end{aligned}$ (23)
From the first-order optimality condition for the above properties, we obtain:
$\langle Ay+\mu x,x-x^{\prime}\rangle\leq 0,~\forall x^{\prime}\in\mathcal{X},$
and
$\langle-A^{\top}x,y-y^{\prime}\rangle\leq 0,~\forall y^{\prime}\in\mathcal{Y}.$
These are equivalent to the first-order optimality condition for (4) and (5). Thus, if $z\in\mathrm{Fix}(T_{\mu})$, then $z\in Z^{\mu}$.
We next show that $z\in Z^{\mu}\Rightarrow z\in\mathrm{Fix}(T_{\mu})$. From the first-order optimality condition for (4) and (5), if $z\in Z^{\mu}$, then we have:
$\langle Ay+\mu x,x-x^{\prime}\rangle\leq 0,~\forall x^{\prime}\in\mathcal{X},$
and
$\langle-A^{\top}x,y-y^{\prime}\rangle\leq 0,~\forall y^{\prime}\in\mathcal{Y}.$
This is equivalent to (23). Hence, if $z\in Z^{\mu}$, then $z\in\mathrm{Fix}(T_{\mu})$.
Therefore, we have $\mathrm{Fix}(T_{\mu})=Z^{\mu}$. ∎
Appendix G Proofs for Theorem 4.3
G.1 Proof of Theorem 4.3
Proof of Theorem 4.3.
First, we show that the number of perturbed games that must be solved is constant and does not depend on the target accuracy $\varepsilon$:
Lemma G.1.
For any target accuracy $\varepsilon>0$, Algorithm 1 with an initial perturbation strength $\mu_{\mathrm{init}}>0$ terminates after at most $K:=\max\left(1,\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{x\in\mathcal{X}}\|x\|/\alpha)}{\ln 2}\rceil+2\right)$ episodes, and outputs a strategy profile $(x,y)$ satisfying $\mathrm{NashConv}(x,y)\leq\varepsilon$.
By Lemma G.1, it suffices to upper bound the iteration complexity for each episode. We thus derive the following iteration complexity for solving asymmetrically perturbed games:
Lemma G.2.
Let us consider an arbitrary $\mu>0$, and assume that $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$. There exists a positive constant $\beta_{\mu}$ depending on $A,\mathcal{X},\mathcal{Y},\eta$, and $\mu$ such that: for any $\delta>0$,
  • $\textsc{AsymP-GDA}_{x}(x^{1},y^{1},\eta,\mu,\delta)$ in Algorithm 1 outputs the strategy profile $(x,y)$ satisfying
    $\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle\leq\delta$
    within at most
    $\frac{2\left(\ln D+\ln\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln(1/\delta)\right)}{\ln\left(1+1/(2\beta_{\mu}^{2})\right)}$
    inner-iterations.
  • $\textsc{AsymP-GDA}_{y}(x^{1},y^{1},\eta,\mu,\delta)$ in Algorithm 1 outputs the strategy profile $(x,y)$ satisfying
    $\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\delta$
    within at most
    $\frac{2\left(\ln D+\ln\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln(1/\delta)\right)}{\ln\left(1+1/(2\beta_{\mu}^{2})\right)}$
    inner-iterations.
By Lemma G.2, for any $k\geq 1$, we obtain the strategy profiles $(\hat{x}_{x}^{\mu_{k}},\hat{y}_{x}^{\mu_{k}})$ and $(\hat{x}_{y}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}})$ satisfying:
$\max_{x^{\prime}\in\mathcal{X}}\langle A\hat{y}_{x}^{\mu_{k}}+\mu\hat{x}_{x}^{\mu_{k}},\hat{x}_{x}^{\mu_{k}}-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}\hat{x}_{x}^{\mu_{k}},\hat{y}_{x}^{\mu_{k}}-y^{\prime}\rangle\leq\frac{\mu_{k}\varepsilon^{2}}{2m},$
$\max_{x^{\prime}\in\mathcal{X}}\langle A\hat{y}_{y}^{\mu_{k}},\hat{x}_{y}^{\mu_{k}}-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}\hat{x}_{y}^{\mu_{k}}+\mu\hat{y}_{y}^{\mu_{k}},y-y^{\prime}\rangle\leq\frac{\mu_{k}\varepsilon^{2}}{2n},$
after at most $\frac{4\left(\ln\left(3\mu_{k}\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln\left(m+n\right)+\ln(2D/\mu_{k})+2\ln(1/\varepsilon)\right)}{\ln\left(1+1/(2\beta_{\mu_{k}}^{2})\right)}$ strategy updates. Therefore, the total iteration complexity is bounded by:
$\sum_{k=1}^{K}\frac{4\left(\ln\left(3\mu_{k}\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln\left(m+n\right)+\ln(2D/\mu_{k})+2\ln(1/\varepsilon)\right)}{\ln\left(1+1/(2\beta_{\mu_{k}}^{2})\right)}$
$\leq 4K\frac{\ln\left(3\mu_{\mathrm{init}}\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln\left(m+n\right)+\ln(4D\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|/\alpha)+2\ln(1/\varepsilon)}{\ln\left(1+1/(2\max_{k\in[K]}\beta_{\mu_{k}}^{2})\right)}.$
Since $K=\max\left(1,\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+2\right)$ is a constant that does not depend on $\varepsilon$, we conclude that we can obtain $\varepsilon$-equilibrium in the original game within $\mathcal{O}(\ln(1/\varepsilon))$ strategy updates. ∎
G.2 Proof of Lemma G.1
Proof of Lemma G.1.
We first show that, for sufficiently large $k$, the equilibrium strategy of the perturbed game exactly matches an equilibrium strategy in the original game:
Lemma G.3.
Let us define $\mu_{k}:=\frac{\mu_{\mathrm{init}}}{2^{k-1}}$, $x^{\mu_{k}}:=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu_{k}}{2}\|x\|^{2}\right\}$, and $y^{\mu_{k}}:=\mathop{\rm arg~max}\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu_{k}}{2}\|y\|^{2}\right\}$. Then, for any $k>\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+1$, we have $x^{\mu_{k}}\in\mathcal{X}^{\ast}$ and $y^{\mu_{k}}\in\mathcal{Y}^{\ast}$.
We also prove that, in the regime where $x^{\mu}\in\mathcal{X}^{\ast}$ and $y^{\mu}\in\mathcal{Y}^{\ast}$, the NashConv can be upper bounded by the suboptimality gap in the perturbed game:
Lemma G.4.
Assume that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\|x\|+\max_{y\in\mathcal{Y}}\|y\|})$. If $(x,y)\in\mathcal{X}\times\mathcal{Y}$ satisfies for an arbitrary non-negative value $\varepsilon\geq 0$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle\leq\frac{\mu}{2m}\varepsilon^{2},$
then:
$\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}-v^{\ast}\leq\varepsilon.$
Furthermore, if $(x,y)$ satisfies:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\frac{\mu}{2n}\varepsilon^{2},$
then:
$v^{\ast}-\min_{x^{\prime}\in\mathcal{X}}(x^{\prime})^{\top}Ay\leq\varepsilon.$
Let us define $K:=\max\left(1,\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+2\right)$. By combining Lemmas G.3 and G.4, whenever $\textsc{AsymP-GDA}_{x}$ and $\textsc{AsymP-GDA}_{y}$ output $\frac{\mu_{k}\varepsilon^{2}}{2\|A\|^{2}\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|^{2}}$-equilibrium in perturbed games for any $k\geq 1$, Algorithm 1 terminates after at most $K$ outer-iterations for achieving the strategy profile $(x,y)$ satisfying $\mathrm{NashConv}(x,y)\leq\varepsilon$. ∎
G.3 Proof of Lemma G.2
Proof of Lemma G.2.
Let us denote the equilibrium for player x’s perturbed game as $x_{x}^{\mu}:=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\max\limits_{y\in\mathcal{Y}}\left\{x^{\top}Ay+\frac{\mu}{2}\|x\|^{2}\right\}$ and $\mathcal{Y}_{x}^{\mu}=\mathop{\rm arg~max}\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay+\frac{\mu}{2}\|x\|^{2}\right\}$. Also, let us denote the equilibrium for player y’s perturbed game as $y_{y}^{\mu}:=\mathop{\rm arg~max}\limits_{y\in\mathcal{Y}}\min\limits_{x\in\mathcal{X}}\left\{x^{\top}Ay-\frac{\mu}{2}\|y\|^{2}\right\}$ and $\mathcal{X}_{y}^{\mu}:=\mathop{\rm arg~min}\limits_{x\in\mathcal{X}}\min\limits_{y\in\mathcal{Y}}\left\{x^{\top}Ay-\frac{\mu}{2}\|y\|^{2}\right\}$,.
By applying the same proof technique as in Theorem 4.1 to a given $\mu$, under the assumption that $\eta\leq\frac{\mu}{\mu^{2}+\|A\|^{2}}$, we have for any $t\geq 1$:
$\left\|\Pi_{\{x_{x}^{\mu}\}\times\mathcal{Y}_{x}^{\mu}}(z^{t+1})-z^{t+1}\right\|^{2}\leq\left(\frac{1}{1+1/(2\beta_{\mu}^{2})}\right)^{t}\left\|\Pi_{\{x_{x}^{\mu}\}\times\mathcal{Y}_{x}^{\mu}}(z^{1})-z^{1}\right\|^{2},$ (24)
where $\beta_{\mu}>0$ is a positive constant depending on $A,\mathcal{X},\mathcal{Y},\eta$, and $\mu$. Then, we can bound the gap function by the distance from the equilibrium set as follows:
Lemma G.5.
We have for any $z\in\mathcal{Z}$ and $z_{x}^{\mu}\in\{x_{x}^{\mu}\}\times\mathcal{Y}_{x}^{\mu}$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z-z_{x}^{\mu}\|.$
Moreover, we have for any $z\in\mathcal{Z}$ and $z_{x}^{\mu}\in\mathcal{X}_{y}^{\mu}\times\{y_{y}^{\mu}\}$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z-z_{y}^{\mu}\|.$
By combining Lemma G.5 and (24), we have for any $t\geq 1$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x^{t+1},y^{t+1}-y^{\prime}\rangle$
$\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z^{t+1}-\Pi_{\{x_{x}^{\mu}\}\times\mathcal{Y}_{x}^{\mu}}(z^{t+1})\|$
$\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\left(\frac{1}{1+1/(2\beta_{\mu}^{2})}\right)^{t/2}\left\|z^{1}-\Pi_{\{x_{x}^{\mu}\}\times\mathcal{Y}_{x}^{\mu}}(z^{1})\right\|$
$\leq D\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\left(\frac{1}{1+1/(2\beta_{\mu}^{2})}\right)^{t/2}.$
Thus, $\textsc{AsymP-GDA}_{x}$ returns the strategies $(x,y)$ satisfying $\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle\leq\delta$ within at most $\frac{2\left(\ln D+\ln\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln(1/\delta)\right)}{\ln\left(1+1/(2\beta_{\mu}^{2})\right)}$ iterations.
By a similar argument, we can show that $\textsc{AsymP-GDA}_{y}$ returns the strategies $(x,y)$ satisfying $\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\delta$ within at most $\frac{2\left(\ln D+\ln\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)+\ln(1/\delta)\right)}{\ln\left(1+1/(2\beta_{\mu}^{2})\right)}$ iterations. ∎
G.4 Proof of Lemma G.3
Proof of Lemma G.3.
When $k>\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+1$, we have $\mu_{k}=\frac{\mu_{\mathrm{init}}}{2^{k-1}}<\frac{\alpha}{\max_{z\in\mathcal{Z}}\|z\|}\leq\frac{\alpha}{\max_{x\in\mathcal{X}}\|x\|}$. Hence, the assumption that $\mu_{k}<\frac{\alpha}{\max_{x\in\mathcal{X}}\|x\|}$ in Theorem 3.1 satisfies when $k>\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{x\in\mathcal{X}}\|x\|/\alpha)}{\ln 2}\rceil+1$. Therefore, we have $x^{\mu_{k}}\in\mathcal{X}^{\ast}$ for any $k>\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+1$. By a similar argument, we have $y^{\mu_{k}}\in\mathcal{Y}^{\ast}$ for any $k>\lceil\frac{\ln(\mu_{\mathrm{init}}\max_{z\in\mathcal{Z}}\|z\|/\alpha)}{\ln 2}\rceil+1$. ∎
G.5 Proof of Lemma G.4
Proof of Lemma G.4.
We have for any $x\in\mathcal{X}$ and $x^{\ast}\in\mathcal{X}^{\ast}$:
$\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}-v^{\ast}$ $=\max_{y^{\prime}\in\mathcal{Y}}\left((x-x^{\ast})^{\top}Ay^{\prime}+(x^{\ast})^{\top}Ay^{\prime}\right)-v^{\ast}$
$\leq\max_{y^{\prime}\in\mathcal{Y}}(x-x^{\ast})^{\top}Ay^{\prime}+\max_{y^{\prime}\in\mathcal{Y}}(x^{\ast})^{\top}Ay^{\prime}-v^{\ast}$
$=\max_{y^{\prime}\in\mathcal{Y}}(x-x^{\ast})^{\top}Ay^{\prime}$
$\leq\sqrt{m}\|x-x^{\ast}\|.$
From Theorem 3.1, under the assumption that $\mu\in(0,\frac{\alpha}{\max_{x\in\mathcal{X}}\|x\|})$, it holds that $x^{\mu}\in\mathcal{X}^{\ast}$. Hence, we get for any $x\in\mathcal{X}$:
$\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}-v^{\ast}$ $\leq\sqrt{m}\|x-x^{\mu}\|.$ (25)
On the other hand, since the function $g_{\mathrm{asym}}^{\mu}(x)=\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}+\frac{\mu}{2}\|x\|^{2}$ is $\mu$-strongly convex, we have for any subgradient $s\in\partial g_{\mathrm{asym}}^{\mu}(x^{\prime})$:
$\forall x\in\mathcal{X},~\frac{\mu}{2}\|x-x^{\prime}\|^{2}$ $\leq\langle s,x^{\prime}-x\rangle+g_{\mathrm{asym}}^{\mu}(x)-g_{\mathrm{asym}}^{\mu}(x^{\prime}).$
From the first-order optimality condition for $x^{\mu}$, there must be a subgradient $s^{\ast}\in\partial g_{\mathrm{asym}}^{\mu}(x^{\mu})$ such that:
$\forall x\in\mathcal{X},~\langle s^{\ast},x^{\mu}-x\rangle\leq 0.$
Thus, setting $x^{\prime}=x^{\mu}$, we have:
$\forall x\in\mathcal{X},~\frac{\mu}{2}\|x-x^{\mu}\|^{2}$ $\leq g_{\mathrm{asym}}^{\mu}(x)-g_{\mathrm{asym}}^{\mu}(x^{\mu}).$
The right-hand side in the above inequality can be upper bounded as follows: for any $y\in\mathcal{Y}$,
$g_{\mathrm{asym}}^{\mu}(x)-g_{\mathrm{asym}}^{\mu}(x^{\mu})$
$=\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}+\frac{\mu}{2}\|x\|^{2}-\min_{x^{\prime}\in\mathcal{X}}\max_{y^{\prime}\in\mathcal{Y}}\left((x^{\prime})^{\top}Ay^{\prime}+\frac{\mu}{2}\|x^{\prime}\|^{2}\right)$
$=\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}+\frac{\mu}{2}\|x\|^{2}-\max_{y^{\prime}\in\mathcal{Y}}\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay^{\prime}+\frac{\mu}{2}\|x^{\prime}\|^{2}\right)$
$\leq\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}+\frac{\mu}{2}\|x\|^{2}-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\|x^{\prime}\|^{2}\right)$
$=\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+x^{\top}Ay+\frac{\mu}{2}\|x\|^{2}-\min_{x^{\prime}\in\mathcal{X}}\left((x^{\prime})^{\top}Ay+\frac{\mu}{2}\|x^{\prime}\|^{2}\right)$
$\leq\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+\langle Ay+\mu x,x-x^{\dagger}\rangle$
$\leq\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle.$
Hence, we obtain:
$\forall(x,y)\in\mathcal{X}\times\mathcal{Y},~\frac{\mu}{2}\|x-x^{\mu}\|^{2}$ $\leq\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle.$ (26)
By combining (25) and (26), we finally obtain:
$\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}-v^{\ast}$ $\leq\sqrt{\frac{2m}{\mu}}\sqrt{\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle}.$
Therefore,
$\max_{y^{\prime}\in\mathcal{Y}}\langle A^{\top}x,y^{\prime}-y\rangle+\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle\leq\frac{\mu}{2m}\varepsilon^{2}\Rightarrow\max_{y^{\prime}\in\mathcal{Y}}x^{\top}Ay^{\prime}-v^{\ast}\leq\varepsilon.$
By a similar argument, under the assumption that $\mu\in(0,\frac{\alpha}{\max_{y\in\mathcal{Y}}\|y\|})$, we have:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\frac{\mu}{2n}\varepsilon^{2}\Rightarrow v^{\ast}-\min_{x^{\prime}\in\mathcal{X}}(x^{\prime})^{\top}Ay\leq\varepsilon.$
G.6 Proof of Lemma G.5
Proof of Lemma G.5.
Let us define:
$F_{\mu}^{x}(x,y,x^{\prime},y^{\prime}):=\langle Ay+\mu x,x-x^{\prime}\rangle+\langle-A^{\top}x,y-y^{\prime}\rangle=-\langle Ay,x^{\prime}\rangle+\mu\langle x,x-x^{\prime}\rangle+\langle A^{\top}x,y^{\prime}\rangle.$
Then, we have:
$F_{\mu}^{x}(x,y,x^{\prime},y^{\prime})-F_{\mu}^{x}(\hat{x},\hat{y},x^{\prime},y^{\prime})$
$=-\langle Ay,x^{\prime}\rangle+\mu\langle x,x-x^{\prime}\rangle+\langle A^{\top}x,y^{\prime}\rangle+\langle A\hat{y},x^{\prime}\rangle-\mu\langle\hat{x},\hat{x}-x^{\prime}\rangle-\langle A^{\top}\hat{x},y^{\prime}\rangle$
$=-\langle A(y-\hat{y}),x^{\prime}\rangle+\mu\langle x,x-x^{\prime}\rangle-\mu\langle\hat{x},\hat{x}-x^{\prime}\rangle+\langle A^{\top}(x-\hat{x}),y^{\prime}\rangle$
$=-\langle A(y-\hat{y}),x^{\prime}\rangle+\mu\langle x-\hat{x},x-x^{\prime}\rangle+\mu\langle\hat{x},x-\hat{x}\rangle+\langle A^{\top}(x-\hat{x}),y^{\prime}\rangle$
$\leq\|A^{\top}x^{\prime}\|\|y-\hat{y}\|+\mu\|x-x^{\prime}\|\|x-\hat{x}\|+\mu\|\hat{x}\|\|x-\hat{x}\|+\|Ay^{\prime}\|\|x-\hat{x}\|$.
Hence, we get for any $z,\hat{z}\in\mathcal{Z}$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle-\max_{x^{\prime}\in\mathcal{X}}\langle A\hat{y}+\mu\hat{x},\hat{x}-x^{\prime}\rangle-\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}\hat{x},\hat{y}-y^{\prime}\rangle$
$=\max_{(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}}F_{\mu}^{x}(x,y,x^{\prime},y^{\prime})-\max_{(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}}F_{\mu}^{x}(\hat{x},\hat{y},x^{\prime},y^{\prime})$
$\leq\max_{(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}}\left(F_{\mu}^{x}(x,y,x^{\prime},y^{\prime})-F_{\mu}^{x}(\hat{x},\hat{y},x^{\prime},y^{\prime})\right)$
[…]
$\leq\max_{x^{\prime}\in\mathcal{X}}\|A^{\top}x^{\prime}\|\|y-\hat{y}\|+3\mu\max_{x^{\prime}\in\mathcal{X}}\|x^{\prime}\|\|x-\hat{x}\|+\max_{y^{\prime}\in\mathcal{Y}}\|Ay^{\prime}\|\|x-\hat{x}\|$
$\leq\left(3\mu\max_{x^{\prime}\in\mathcal{X}}\|x^{\prime}\|+\max_{x^{\prime}\in\mathcal{X}}\|A^{\top}x^{\prime}\|+\max_{y^{\prime}\in\mathcal{Y}}\|Ay^{\prime}\|\right)\|z-\hat{z}\|$
$\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z-\hat{z}\|$.
Here, from the first-order optimality condition for $x^{\mu}$ and $y^{\mu}\in\mathcal{Y}^{\mu}$, we have:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay^{\mu}+\mu x^{\mu},x^{\mu}-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x^{\mu},y^{\mu}-y^{\prime}\rangle\leq 0.$
By combining these inequalities, we obtain for any $z\in\mathcal{Z}$ and $z^{\mu}\in\mathcal{Z}^{\mu}$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay+\mu x,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x,y-y^{\prime}\rangle\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z-z^{\mu}\|.$
By a similar argument, we have for any $z\in\mathcal{Z}$ and $z^{\mu}\in\mathcal{Z}^{\mu}$:
$\max_{x^{\prime}\in\mathcal{X}}\langle Ay,x-x^{\prime}\rangle+\max_{y^{\prime}\in\mathcal{Y}}\langle-A^{\top}x+\mu y,y-y^{\prime}\rangle\leq\left(3\mu\max_{z^{\prime}\in\mathcal{Z}}\|z^{\prime}\|+\sqrt{m}+\sqrt{n}\right)\|z-z^{\mu}\|.$
Appendix H Proof of Theorem B.1
Proof of Theorem B.1.
First, we prove that $y^{\mu}\neq y^{\ast}$ under the assumption that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}$ by contradiction. We assume that $y^{\mu}=y^{\ast}$. Since $(x^{\ast},y^{\ast})$ is in the interior of $\Delta^{m}\times\Delta^{n}$, we have:
$\begin{aligned} (A^{\top}x^{\ast})_{i}=v^{\ast},~\forall i\in[m]\\ (Ay^{\ast})_{i}=v^{\ast},~\forall i\in[n].\end{aligned}$ (27)
Then, from (27), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}$ $=x^{\top}Ay^{\ast}+\frac{\mu}{2}\left\|x\right\|^{2}=v^{\ast}+\frac{\mu}{2}\left\|x\right\|^{2}.$ (28)
On the other hand,
$\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\ast}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=v^{\ast}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$ (29)
By combining (28) and (29), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu}{2}\left\|x\right\|^{2}\geq\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$
Hence, from the property of the player $x$’s equilibrium strategy in the perturbed game, $x^{\mu}$ must satisfy $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$.
On the other hand, from the property of the player $y$’s equilibrium strategy $y^{\mu}$ in the perturbed game, $y^{\mu}$ is an optimal solution of the following optimization problem:
$\max_{y\in\Delta^{n}}\left\{(x^{\mu})^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}\right\}.$
Let us define the following Lagrangian function $L(y,\kappa,\lambda)$ as:
$L(y,\kappa,\lambda)=(x^{\mu})^{\top}Ay-\frac{\mu}{2}\left\|y\right\|^{2}-\sum_{i=1}^{n}\kappa_{i}g_{i}(y)-\lambda h(y),$
where $g_{i}(y)=-y_{i}$ and $h(y)=\sum_{i=1}^{n}y_{i}-1$. Then, from the KKT conditions, we get the stationarity:
$A^{\top}x^{\mu}-\mu y^{\mu}-\sum_{i=1}^{n}\kappa_{i}\nabla g_{i}(y^{\mu})-\lambda\nabla h(y^{\mu})=\mathbf{0}_{n},$ (30)
and the complementary slackness:
$\forall i\in[n],~\kappa_{i}g_{i}(y^{\mu})=0.$ (31)
Since $y^{\mu}=y^{\ast}$ and $y^{\ast}$ is in the interior of $\Delta^{n}$, we have $g(y^{\mu})=-y_{i}^{\mu}<0$ for all $i\in[n]$. Thus, from (31), we have $\kappa_{i}=0$ for all $i\in[n]$. Substituting this into (30), we obtain:
$A^{\top}x^{\mu}-\mu y^{\mu}-\lambda\nabla h(y^{\mu})=A^{\top}x^{\mu}-\mu y^{\mu}-\lambda\mathbf{1}_{n}=\mathbf{0}_{n}.$ (32)
Hence, we have:
$\lambda=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}.$ (33)
Putting (33) into (32) yields:
$y^{\mu}=\frac{1}{\mu}\left(A^{\top}x^{\mu}-\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}\mathbf{1}_{n}\right)=\frac{1}{\mu}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-\frac{\mathbf{1}_{n}^{\top}A^{\top}\frac{\mathbf{1}_{m}}{m}-\mu}{n}\mathbf{1}_{n}\right),$
where the second equality follows from $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$. Multiplying this by $\frac{\mathbf{1}_{m}^{\top}}{m}A$, we have:
$v^{\ast}=\frac{1}{\mu}\left(\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}\right),$ (34)
where we used the assumption that $y^{\mu}=y^{\ast}$ and (27). Here, we have:
$\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}$
$=\left\|v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=n(v^{\ast})^{2}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}+2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$\phantom{=}-n(v^{\ast})^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}-2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\mu v^{\ast}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu v^{\ast}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu v^{\ast}+\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right).$ (35)
Here, since $A^{\top}\frac{\mathbf{1}_{m}}{m}=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu}{n}\mathbf{1}_{n}+\mu y^{\mu}=\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}\mathbf{1}_{n}+\mu y^{\ast}$ from (32) and (33), we get:
$\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)$
$\phantom{=}+\frac{\mu}{n}\mathbf{1}_{n}^{\top}\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu y^{\ast}\right)$
$=\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)+\frac{\mu^{2}}{n}+\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)\mathbf{1}_{n}^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$\phantom{=}+\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\mathbf{1}_{n}+\mu^{2}(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$=\mu\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu}{n}-v^{\ast}\right)+\mu^{2}\left\|y^{\ast}\right\|^{2}$
$=\mu\left(\mu\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}\right).$ (36)
By combining (34), (35), and (36), we have:
$\mu\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}=0.$
Therefore, if $y^{\mu}=y^{\ast}$, then $\mu$ must satisfy:
$\mu=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)},$
and this is equivalent to:
$\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}\Rightarrow y^{\mu}\neq y^{\ast}.$
By a similar argument, in terms of player $x$, we can conclude that:
$\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left(\left\|x^{\ast}\right\|^{2}-\frac{1}{m}\right)}\Rightarrow x^{\mu}\neq x^{\ast}.$
Appendix I Proof of Theorem C.1
Proof of Theorem C.1.
First, we prove that $y^{\mu}\neq y^{\ast}$ under the assumption that $\mu\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left\|y^{\ast}\right\|^{2}-\frac{1}{n}}$ by contradiction. We assume that $y^{\mu}=y^{\ast}$. From (27), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}$ $=x^{\top}Ay^{\ast}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}=v^{\ast}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}.$ (37)
On the other hand,
$\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\ast}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}=v^{\ast}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$ (38)
By combining (37) and (38), we have for any $x\in\Delta^{m}$:
$x^{\top}Ay^{\mu}+\frac{\mu_{x}}{2}\left\|x\right\|^{2}\geq\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A^{\top}y^{\mu}+\frac{\mu_{x}}{2}\left\|\frac{\mathbf{1}_{m}}{m}\right\|^{2}.$
Hence, from the property of the player $x$’s equilibrium strategy in the perturbed game, $x^{\mu}$ must satisfy $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$.
On the other hand, from the property of the player $y$’s equilibrium strategy $y^{\mu}$ in the perturbed game, $y^{\mu}$ is an optimal solution of the following optimization problem:
$\max_{y\in\Delta^{n}}\left\{(x^{\mu})^{\top}Ay-\frac{\mu_{y}}{2}\left\|y\right\|^{2}\right\}.$
Let us define the following Lagrangian function $L(y,\kappa,\lambda)$ as:
$L(y,\kappa,\lambda)=(x^{\mu})^{\top}Ay-\frac{\mu_{y}}{2}\left\|y\right\|^{2}-\sum_{i=1}^{n}\kappa_{i}g_{i}(y)-\lambda h(y),$
where $g_{i}(y)=-y_{i}$ and $h(y)=\sum_{i=1}^{n}y_{i}-1$. Then, from the KKT conditions, we get the stationarity:
$A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\sum_{i=1}^{n}\kappa_{i}\nabla g_{i}(y^{\mu})-\lambda\nabla h(y^{\mu})=\mathbf{0}_{n},$ (39)
and the complementary slackness:
$\forall i\in[n],~\kappa_{i}g_{i}(y^{\mu})=0.$ (40)
Since $y^{\mu}=y^{\ast}$ and $y^{\ast}$ is in the interior of $\Delta^{n}$, we have $g(y^{\mu})=-y_{i}^{\mu}<0$ for all $i\in[n]$. Thus, from (40), we have $\kappa_{i}=0$ for all $i\in[n]$. Substituting this into (39), we obtain:
$A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\lambda\nabla h(y^{\mu})=A^{\top}x^{\mu}-\mu_{y}y^{\mu}-\lambda\mathbf{1}_{n}=\mathbf{0}_{n}.$ (41)
Hence, we have:
$\lambda=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}.$ (42)
Putting (42) into (41) yields:
$y^{\mu}=\frac{1}{\mu_{y}}\left(A^{\top}x^{\mu}-\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}\mathbf{1}_{n}\right)=\frac{1}{\mu_{y}}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-\frac{\mathbf{1}_{n}^{\top}A^{\top}\frac{\mathbf{1}_{m}}{m}-\mu_{y}}{n}\mathbf{1}_{n}\right),$
where the second equality follows from $x^{\mu}=\frac{\mathbf{1}_{m}}{m}$. Multiplying this by $\frac{\mathbf{1}_{m}^{\top}}{m}A$, we have:
$v^{\ast}=\frac{1}{\mu_{y}}\left(\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu_{y}\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}\right),$ (43)
where we used the assumption that $y^{\mu}=y^{\ast}$ and (27). Here, we have:
$\frac{1}{m^{2}}\left\|A^{\top}\mathbf{1}_{m}\right\|^{2}-\frac{1}{m^{2}n}(\mathbf{1}_{n}^{\top}A^{\top}\mathbf{1}_{m})^{2}+\mu_{y}\frac{1}{mn}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}$
$=\left\|v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(v^{\ast}\mathbf{1}_{n}+A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=n(v^{\ast})^{2}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}+2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$\phantom{=}-n(v^{\ast})^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}-2v^{\ast}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\mu_{y}v^{\ast}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu_{y}v^{\ast}+\left\|A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right\|^{2}-\frac{1}{n}\left(\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)\right)^{2}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\mu_{y}v^{\ast}+\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right).$ (44)
Here, since $A^{\top}\frac{\mathbf{1}_{m}}{m}=\frac{\mathbf{1}_{n}^{\top}A^{\top}x^{\mu}-\mu_{y}}{n}\mathbf{1}_{n}+\mu_{y}y^{\mu}=\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}\mathbf{1}_{n}+\mu_{y}y^{\ast}$ from (41) and (42), we get:
$\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(A^{\top}\frac{\mathbf{1}_{m}}{m}-v^{\ast}\mathbf{1}_{n}\right)$
$=\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)$
$\phantom{=}+\frac{\mu_{y}}{n}\mathbf{1}_{n}^{\top}\left(\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}+\mu_{y}y^{\ast}\right)$
$=\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)+\frac{\mu_{y}^{2}}{n}+\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)\mathbf{1}_{n}^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$\phantom{=}+\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)\mathbf{1}_{n}+\mu_{y}^{2}(y^{\ast})^{\top}\left(\mathbb{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\right)y^{\ast}$
$=\mu_{y}\left(\frac{\frac{1}{m}\mathbf{1}_{m}^{\top}A\mathbf{1}_{n}-\mu_{y}}{n}-v^{\ast}\right)+\mu_{y}^{2}\left\|y^{\ast}\right\|^{2}$
$=\mu_{y}\left(\mu_{y}\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}\right).$ (45)
By combining (43), (44), and (45), we have:
$\mu_{y}\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)+\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}=0.$
Therefore, if $y^{\mu}=y^{\ast}$, then $\mu_{y}$ must satisfy:
$\mu_{y}=\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)},$
and this is equivalent to:
$\mu_{y}\neq\frac{v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)}{\left(\left\|y^{\ast}\right\|^{2}-\frac{1}{n}\right)}\Rightarrow y^{\mu}\neq y^{\ast}.$
By a similar argument, in terms of player $x$, we can conclude that:
$\mu_{x}\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left(\left\|x^{\ast}\right\|^{2}-\frac{1}{m}\right)}\Rightarrow x^{\mu}\neq x^{\ast}.$